### v1

# Graph Limits via Quotients

Abstract
We present a new notion of limits of weighted directed graphs of growing size based on convergence of their random quotients. These limits are specified in terms of random exchangeable measures on the unit square. We call our limits grapheurs and show that these are dual to graphons in a precise sense. Grapheurs are well-suited to modeling hubs and connections between them in large graphs; previous notions of graph limits based on subgraph densities fail to adequately model such global structures as subgraphs are inherently local. Using our framework, we present an edge-based sampling approach for testing properties pertaining to hubs in large graphs. This method relies on an edge-based analog of the Szemerédi regularity lemma, whereby we show that sampling a small number of edges from a large graph approximately preserves its quotients. Finally, we observe that the random quotients of a graph are related to each other by equipartitions, and we conclude with a characterization of such random graph models.
Keywords: De Finetti, graphons, hubs, property testing, Szemerédi regularity
1 Introduction
Analyzing the structure of large graphs is a challenge that arises in numerous problem domains such as bioinformatics, network tomography, operations research, and the social sciences [newman_networks, network_tomography, social_network, JACKSON2002265, bassett2017network]. Graphons and other related notions of graph limits in the literature employ subgraph densities to summarize and compare large graphs. However, subgraph-based measures are inherently local and they fail to suitably account for global structures such as hubs and connections between them, which are commonly encountered in real-world networks [OKELLY1998171, hub_and_spoke, power2013evidence, hubs_in_drugs, van2013network]. For example, most subgraphs of a star, the prototypical example of a hub, do not contain any edges.
A hub is a vertex in a graph [^1] We consider directed graphs with possible self-loops and nonnegative edge weights. with a significant fraction of the total edge weight incident to it. Our central thesis is that hubs in large graphs can be detected by their quotients. Formally, a quotient of size $k$ of a graph on $n$ vertices is defined by an ordered partition $f_{k,n}:[n]\rightarrow[k]$, which partitions the vertex set $[n]$ into $k$ parts that are given by the fibers $f^{-1}_{k,n}(j)$ for $j\in[k]$. In particular, for a graph on $n$ vertices with adjacency matrix $G\in\mathbb{R}^{n\times n}_{+}$, its quotient by the ordered partition $f_{k,n}$ is given by the graph $\rho(f_{k,n})G\in\mathbb{R}^{k\times k}_{+}$ on $k$ vertices with edge weights:
$\rho(f_{k,n})G]_{i,j}=\sum_{\begin{subarray}{c}v\in f_{k,n}^{-1}(i)\\ u\in f_{k,n}^{-1}(j)\end{subarray}}G_{v,u}.$ (1)
Here we view $\rho(f_{k,n}):\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{k\times k}$ as a ‘linear representation’ of $f_{k,n}$. In words, the vertices of the quotient $\rho(f_{k,n})G$ correspond to the parts in the partition defined by $f_{k,n}$, and the edge weight between parts $i,j\in[k]$ is given by sums of all edge weights in $G$ between vertices in parts $f_{k,n}^{-1}(i)$ and $f_{k,n}^{-1}(j)$. Thus, a quotient of $G$ represents the edge weights between each pair of parts in a vertex partition.
As an illustration, we give in Figures 1-2 examples of small quotients of two networks. The first, depicted in Figure 1, is a traffic network representing the fraction of trips starting and ending at each zone of the Gold Coast, Australia [traffic_networks]. The second, depicted in Figure 2, is a brain connectome, the network of connections between neurons of an adult male C. elegans roundworm [cook2019whole, worms_brain]. The structure of each quotient indicates some of the features of the underlying large network. As we will see in Section 4.3, the presence of hubs in the original network leads to nonuniform distribution of edge weight in their random quotients.
Figure 1: Traffic network for the Gold Coast, Australia, and an associated random quotient. The city is divided into 1068 zones, and the weight of an edge is the fraction of trips from a dataset with given start and end zones [traffic_networks]. Here we plot the first 300 of these zones and their interconnections, displaying the width of an edge proportionally to its weight. We color the vertices according to their image in the quotient, and color and edge connecting two vertices with the same image by the color of that image. The nonuniformity in the edge weights of the quotient indicates the presence of hubs that we can see in the plot of the entire network.
Figure 2: Network of chemical connections among the 575 neurons of adult male C. elegans (a species of roundworm), and its random quotient. Edges are directed from pre-synaptic to post-synaptic cells, and their weight is the total number of EM serial sections of connectivity [cook2019whole, worms_brain]. The nonuniform distribution of self-loops and the unequal in-degrees and out-degrees among the four vertices in the quotient again indicate the presence of hubs.
In this paper, we define a new notion of convergence for sequences of growing-sized graphs based on convergence of their random quotients. In analogy with graphons [lovasz2012large, §1], we define suitable limit objects in terms of random measures for any quotient-convergent sequence of graphs. Finally, we show that finite (random) graphs sampled from our limit objects are quotient-convergent to the limit almost surely. The key concepts in our paper are dual to those in the literature on graphons, and we emphasize this duality throughout our development. In particular, we shall see that random quotients and random subgraphs, in terms of which graphon convergence is defined, are random linear maps which are adjoints of each other. Further, a dual analog of the cut metric plays a prominent role in our development, and we derive a dual Szemerédi lemma as a consequence of our results. We describe next these contributions in more detail.
1.1 Our Contributions
To set the stage, we define a random quotient of a graph on $n$ vertices in terms of a uniformly random map $F_{k,n}\colon[n]\to[k]$. That is, each $i\in[n]$ gets sent independently to a uniform element $F_{k,n}(i)\in[k]$. Equivalently, this is a uniformly random ordered partition of $[n]$ into $k$ parts, given by the sequence of (random) fibers $F_{k,n}^{-1}(j)$ for $j\in[k]$. We denote by $\Delta^{n\times n}$ the set of $n\times n$ matrices with nonnegative entries that sum to one.
Definition 1.1.
A sequence of weighted graphs $(G_{n}\in\Delta^{n\times n})_{n=1}^{\infty}$ with edge weights that sum to one quotient-converges if for each $k\in\mathbb{N}$ the sequence of random graphs $(\rho(F_{k,n})G_{n}\in\Delta^{k\times k})_{n=1}^{\infty}$ converges weakly. A general sequence of nonzero weighted graphs $(G_{n}\in\mathbb{R}^{n\times n}_{+})_{n}$ quotient-converges if $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1}\in\Delta^{n\times n})_{n}$ quotient-converges.
Normalizing our graphs to have unit total edge weight does not change the structure of its hubs, which were defined above in terms of the fraction of edge weight incident on them. Unless specified otherwise, we will assume throughout that the total edge weight of our graphs equals one.
Note that $\rho(F_{n,k})^{\star}G_{n}$ is the induced subgraph on $k$ vertices sampled with replacement from the vertex set $[n]$ of $G_{n}$, in terms of which graphon convergence is defined [lovasz2012large, §9], giving a first sense in which our theory is dual to that of graphons. We remark that there is an alternative view of graphon limits via convergence of quotients [lovasz2012large, §12], but these latter quotients differ from ours. In particular, these latter quotients cannot detect hubs in a large graph, as we explain in Section 1.2 below.
1.1.1 Combinatorial Aspects of Quotient Convergence (Section 2)
As our first contribution, we analyze combinatorial aspects of quotient convergence in Section 2. Whereas previous notions of graph convergence were based on convergence of local “motifs” in the form of homomorphism densities (appropriately normalized), we show that our quotient convergence is equivalent to convergence of homomorphism numbers. More precisely, for a multigraph $H\in\mathbb{N}^{k\times k}$ possibly containing parallel edges and self-loops, and a general weighted graph $G\in\mathbb{R}^{n\times n}$ on $n$ vertices, the homomorphism number and homomorphism density of $H$ in $G$ are given by [lovasz2012large, Eq. (5.3)]
$\mathrm{hom}(H;G)=\sum_{f\colon[k]\to[n]}\prod_{i,j=1}^{k}G_{f(i),f(j)}^{H_{i,j}},\qquad t(H;G)=\frac{1}{n^{k}}\mathrm{hom}(H;G).$ (2)
The terminology comes from the fact that when $H$ and $G$ are simple graphs, then $\mathrm{hom}(H;G)$ is the number of graph homomorphisms from $H$ to $G$ and $t(H;G)$ counts the average number of times $H$ occurs in a uniformly random subgraph of $G$. In particular, convergence of the homomorphism densities $t(H;\cdot)$ is used to define limits of dense graphs [LOVASZ2006933, convergent_seqs1].
In analogy, we define the quotient density $t_{Q}(H;G)$ of a multigraph $H\in\mathbb{N}^{k\times k}$ in a general weighted graph $G\in\mathbb{R}^{n\times n}$ as the average number of times $H$ occurs in a uniformly random quotient of $G$:
$t_{Q}(H;G)=\frac{1}{k^{n}}\sum_{f\colon[n]\to[k]}\prod_{i,j=1}^{k}(\rho(f)G)_{i,j}^{H_{i,j}}.$ (3)
The role of homomorphism densities in graphons is played by quotient densities in our framework. We prove that a sequence of graphs is quotient-convergent if all of their quotient densities converge.
Theorem 1.2.
For a sequence of graphs $(G_{n}\in\Delta^{n\times n})_{n}$, the following are equivalent:
  1. 1.
    The sequence $(G_{n})$ is quotient-convergent.
  2. 2.
    For each multigraph $H\in\mathbb{N}^{k\times k}$, the sequence of quotient densities $(t_{Q}(H;G_{n}))_{n}$ converges.
  3. 3.
    For each multigraph $H\in\mathbb{N}^{k\times k}$, the sequence of homomorphism numbers $(\mathrm{hom}(H;G_{n}))_{n}$ converges.
The proof of this result is given in Section 2 based on expressing quotient densities in terms of (injective) homomorphism numbers. This expression involves a partial order on the collection of multigraphs induced by the quotient operation.
For simple graphs $(G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}})_{n}$, i.e., undirected graphs with weights $(G_{n})_{i,j}\in\{0,1\}$ and with no self-loops so $\mathrm{diag}(G_{n})=0$, we conclude that $(G_{n}/2|E(G_{n})|)_{n}$ is quotient-convergent if and only if $\left(\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}}\right)_{n}$ converges for all simple graphs $H\in\mathbb{S}^{k}_{\mathrm{sim}}$ (see Corollary 2.7). Thus, quotient-convergence of simple graphs is equivalent to convergence of a homomorphism-density-like quantity, but one that is differently normalized than usual. To the best of our knowledge, such a normalization based on the number of edges has not appeared previously in the literature, as normalizations based on the numbers of vertices $k$ and $n$ are used (see Section 1.2). As a consequence, we will see below in our discussion of limit objects (Corollary 3.7) that quotient-convergent sequences of both dense and degree-bounded sequences of simple graphs converge to the same (uninteresting) limit. In fact, quotient limits of simple graphs can only detect stars, which are vertices on which a positive fraction of edges are incident. More broadly, the utility of our framework lies in uncovering such hub structure in weighted graphs.
1.1.2 Limit Objects for Quotient Convergence (Section 3)
Our second contribution consists of constructing analytic limit objects together with a notion of distance between them that allows us to compare graphs of different sizes, akin to graphons and the cut metric for dense graph sequences. Note that appending isolated vertices to a graph $G\in\Delta^{n\times n}$ or relabeling its vertices does not change the distribution of its quotients $\rho(F_{k,n})G_{n}$ for each $k\in\mathbb{N}$. Therefore, the limit object we associate to $G$ must be the same as that associated to any other graph differing from it by extra isolated vertices and vertex relabeling. We describe our limit objects in terms of certain random exchangeable measures on the square, which we define next.
Definition 1.3 (Random exchangeable measures).
A random exchangeable probability measure on $0,1]^{2}$ is a random measure [^2] We endow the space of probability measures on $0,1]^{2}$ with the weak topology and consider Borel random variables on it. $\mathsf{M}$ satisfying $\mathsf{M}\overset{d}{=}\mathsf{M}\circ(\sigma,\sigma)$ for all (Lebesgue-)measure-preserving bijections $\sigma\colon[0,1]\to[0,1]$. We say that a sequence of random measures $(\mathsf{M}_{n})$ converges if it converges weakly viewed as a sequence of random variables taking values in the space of probability measures on $0,1]^{2}$ (see Section 1.3 for details).
Any finite graph $G\in\Delta^{n\times n}$ can be associated with the random exchangeable measure:
$\mathsf{M}_{G}=\sum_{i,j=1}^{n}G_{i,j}\delta_{(T_{i},T_{j})},\quad\textrm{where }T_{1},\ldots,T_{n}\overset{iid}{\sim}\mathrm{Unif}([0,1]).$ (4)
Note that relabeling the vertices of $G$ or appending isolated vertices to it yields the same random measure in this way. More generally, Kallenberg showed that all random exchangeable measures on $0,1]^{2}$ have the form [kallenberg1990exchangeable]:
$\mathsf{M}=\sum_{i,j\in\mathbb{N}}E_{i,j}\delta_{(T_{i},T_{j})}+\sum_{i\in\mathbb{N}}[\sigma_{i}(\delta_{T_{i}}\otimes\lambda)+\varsigma_{i}(\lambda\otimes\delta_{T_{i}})]+\theta\lambda^{2}+\vartheta\lambda_{D},$ (5)
where $(i)$ $\lambda$ is the uniform measure on $[0,1]$, $\lambda^{2}$ is the uniform measure on $0,1]^{2}$, and $\lambda_{D}$ is the uniform measure on the diagonal $\{(x,x):x\in[0,1]\}$; $(ii)$ $T_{1},T_{2},\ldots\overset{iid}{\sim}\mathrm{Unif}([0,1])$; and $(iii)$ $E_{i,j},\sigma_{i},\varsigma_{i},\theta,\vartheta\geq 0$ may potentially be random but must be independent of the $(T_{i})$ and must satisfy $\sum_{i,j}E_{i,j}+\sum_{i}(\sigma_{i}+\varsigma_{i})+\theta+\vartheta=1$ almost surely. Moreover, the collection of extremal random exchangeable measures, i.e., those that are not mixtures of other such random measures, are precisely the ones for which the coefficients $E,\sigma,\varsigma,\theta,\vartheta$ in (5) can be taken to be deterministic.
Definition 1.4.
A grapheur (graph edge measure) is an extremal random exchangeable measure, i.e., one of the form (5) with deterministic coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$. We denote the collection of grapheurs by
[…]
The space of grapheurs inherits the weak topology from the ambient space of random exchangeable measures.
The coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ describing a fixed grapheur $\mathsf{M}\in\mathcal{M}$ are essentially unique. Specifically, note that for any permutation $\pi\colon\mathbb{N}\to\mathbb{N}$, the coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ and $(\pi E\pi^{\top},\pi\sigma,\pi\varsigma,\theta,\vartheta)$ define the same random exchangeable distribution via (5). Further, adding zero rows to $E$ and corresponding zero entries to $\sigma$, or zero columns to $E$ and corresponding zero entries to $\varsigma$, does not change the resulting random measure. Kallenberg showed in [kallenberg1990exchangeable, Prop. 3] that these are the only possible ambiguities in the coefficients. In particular, note that $\mathsf{M}_{G}\in\mathcal{M}$ is a grapheur for any finite graph $G$, and that $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ for two such graphs $G,G^{\prime}$ if and only if they differ by isolated vertices and vertex relabeling.
In analogy with the cut metric between graphons, we endow $\mathcal{M}$ with the following metric between grapheurs that metrizes their convergence:
$W_{\square}(\mathsf{M}_{1},\mathsf{M}_{2})=\inf_{\begin{subarray}{c}\textrm{random }(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})\\ \mathsf{M}_{i}^{\prime}\overset{d}{=}\mathsf{M}_{i}\textrm{ for }i=1,2\end{subarray}}\ \sup_{\begin{subarray}{c}S,T\subseteq[0,1]\\ \textrm{intervals}\end{subarray}}\mathbb{E}|\mathsf{M}_{1}^{\prime}(S\times T)-\mathsf{M}_{2}^{\prime}(S\times T)|,$ (7)
where the infimum is over couplings $(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})$ of the two random measures $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$. Here $\mathsf{M}_{i}^{\prime}\overset{d}{=}\mathsf{M}_{i}$ means that the distribution of $\mathsf{M}_{i}^{\prime}$ is equal to that of $\mathsf{M}_{i}$, see Section 1.3 for more detail. We show that grapheurs are the limits of quotient-convergent sequences of graphs in the following precise sense.
Theorem 1.5.
A sequence of grapheurs $(\mathsf{M}_{n})\subseteq\mathcal{M}$ converges weakly to a grapheur $\mathsf{M}\in\mathcal{M}$ if and only if $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$. Furthermore, the following statements hold.
  1. 1.
    A sequence of graphs $(G_{n}\in\Delta^{n\times n})$ is quotient-convergent if and only if there is a grapheur $\mathsf{M}\in\mathcal{M}$ satisfying $\lim_{n}W_{\square}(\mathsf{M},\mathsf{M}_{G_{n}})=0$.
  2. 2.
    Any sequence of graphs $(G_{n}\in\Delta^{n\times n})$ has a quotient-convergent subsequence.
  3. 3.
    For every grapheur $\mathsf{M}\in\mathcal{M}$ there is a sequence $(G_{n}\in\Delta^{n\times n})$ satisfying $\lim_{n}W_{\square}(\mathsf{M},\mathsf{M}_{G_{n}})=0$.
In view of Theorem 1.5, we say that a sequence of graphs $(G_{n})$ converges to a grapheur $\mathsf{M}\in\mathcal{M}$ if $\lim_{n}\mathsf{M}_{G_{n}}=\mathsf{M}$ in $\mathcal{M}$. Theorem 1.5 states that any quotient-convergent sequence of graphs converges to a grapheur, and conversely each grapheur is the limit of a quotient-convergent sequence of graphs. Moreover, this theorem shows that the space of grapheurs endowed with the weak topology is compact, and that this compact topology is metrized by our $W_{\square}$ metric (7).
To prove Theorem 1.5, we relate random quotients of finite graphs to their associated grapheurs (4). For each grapheur $\mathsf{M}\in\mathcal{M}$ and $k\in\mathbb{N}$, define the random graph $\mathsf{G}_{k}[\mathsf{M}]\in\Delta^{k\times k}$ whose entries are given by:
$(\mathsf{G}_{k}[\mathsf{M}])_{i,j}=\mathsf{M}(I_{i}^{(n)}\times I_{j}^{(n)})$ (8)
for $i,j\in[n]$, where $I_{i}^{(n)}=[(i-1)/n,i/n)$ for $i\in[n]$. [^3] Note that $\sum_{i,j}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}=\mathsf{M}([0,1)^{2})=1$ almost surely by (5). In words, we partition the square $0,1]^{2}$ into a $k\times k$ uniform grid, and let the entries of $\mathsf{G}_{k}[\mathsf{M}]$ be the (random) measure under $\mathsf{M}$ of each square in this grid. We show in Section 3.1 that $\rho(F_{k,n})G\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}_{G}]$ for any finite graph $G\in\Delta^{n\times n}$, and that a sequence of graphs $(G_{n})$ converges to the grapheur $\mathsf{M}$ precisely when the random quotients $(\rho(F_{k,n})G_{n})_{k}$ converge weakly to $\mathsf{G}_{k}[\mathsf{M}]$ for each $k\in\mathbb{N}$. Similarly, we can extend quotient densities to grapheurs by setting
$t_{Q}(H;\mathsf{M})=\mathbb{E}\prod_{i,j=1}^{k}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}^{H_{i,j}}=\int\prod_{i,j\in[k]}\mu(I_{i}^{(k)}\times I_{j}^{(k)})^{H_{i,j}}\,d\mathsf{M}(\mu),$ (9)
for any multigraph $H\in\mathbb{N}^{k\times k}$, analogously to the extension of homomorphism densities to graphons [lovasz2012large, §7.2]. In terms of quotient densities, the sequence $(G_{n})$ converges to the grapheur $\mathsf{M}$ if and only if $t_{Q}(H;G_{n})\to t_{Q}(H;\mathsf{M})$ for all multigraphs $H$, see Section 3.1.
Having characterized our limit objects, we interpret in Section 3.2 each of the parameters $E,\sigma,\varsigma,\theta,\vartheta$ describing a grapheur (5) in terms of certain features of the large graphs converging to it. Informally, the array $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ specifies the connections between a discrete collection of hubs, with each entry corresponding to an edge with an asymptotically positive fraction of the total edge weight. The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars centered at hubs, with each entry corresponding to the fraction of out-degree or in-degree of a hub from all the edges incident on it with asymptotically vanishing edge weights. Lastly, the parameters $\theta$ and $\vartheta$ specify the remaining edge weight that comes from all edges between vertices with asymptotically vanishing degrees and their self-loops. In particular, we say that a grapheur $\mathsf{M}$ does not have hubs if $E=0$ and $\sigma=\varsigma=0$, and we say that a sequence $(G_{n})$ of graphs does not contain hubs asymptotically if it quotient-converges to such a grapheur.
Finally, we expand on the duality between graphons and grapheurs. We have seen previously that the sampling operators used to define each notion of graph limit, namely, those extracting random subgraphs and random quotients, respectively, are duals of each other. We further show in Section 3.3 that the space of graphons endowed with the cut metric is dual to the space $\mathcal{M}$ endowed with the $W_{\square}$ metric (7) in a precise sense.
Theorem 1.6 (Informal; see Theorem 3.10).
There is a random pairing $\langle\mathsf{M},W\rangle$ associating a scalar random variable to each grapheur $\mathsf{M}$ and graphon $W$ satisfying the following properties. We have $\lim_{n}\mathsf{M}_{n}=\mathsf{M}$ in the $W_{\square}$ metric if and only if $\langle\mathsf{M}_{n},W_{G}\rangle\to\langle\mathsf{M},W_{G}\rangle$ weakly for each step graphon $W_{G}$ associated to a finite simple graph $G$. We also have $\lim_{n}W_{n}\to W$ in the cut metric if and only if $\langle\mathsf{M}_{G},W_{n}\rangle\to\langle\mathsf{M}_{G},W\rangle$ weakly for each grapheur $\mathsf{M}_{G}$ associated to a finite undirected graph $G$ with unit total edge weight and no self-loops.
1.1.3 Edge Sampling and Property Testing (Section 4)
As our next contribution, we apply our framework to test properties of large graphs based on small summaries of them. Specifically, we focus on testing properties given by graph parameters that are continuous with respect to quotient convergence. While it would be natural to test such properties using random quotients, forming these requires access to the entire large graph, which limits their utility in practice. Motivated by this challenge, we develop an edge-based sampling perspective on quotient-convergence in Section 4.1. Specifically, we sample edges from a graph proportionally to their edge weights to form random graphs $\mathsf{G}^{(n)}$ with $n$ edges, and extend this sampling procedure to any grapheur $\mathsf{M}\in\mathcal{M}$. We detail the general sampling procedure in Section 4.1.
We prove the following convergence results for the sequence of random graphs $(\mathsf{G}^{(n)})$ sampled from a grapheur $\mathsf{M}$ back to $\mathsf{M}$, which are edge-based analogs of the second sampling lemma and Szemerédi’s regularity lemma for graphons [lovasz2012large, Lemma 10.16, Lemma 9.15]. Here and throughout the paper, a superscript $G^{(n)}$ denotes a graph on $n$ edges (and hence at most $2n$ vertices) while a subscript $G_{n}$ denotes a graph on $n$ vertices.
Theorem 1.7.
Fix any grapheur $\mathsf{M}\in\mathcal{M}$.
  1. fnum@enumiitem(Edge sampling lemma)
    For each $n\in\mathbb{N}$, let $\mathsf{G}^{(n)}$ be the random graph containing $n$ edges sampled from $\mathsf{M}$ as in Section 4. With probability at least $1-e^{-2\epsilon^{2}}$, we have that $W_{\square}(\mathsf{M}_{\mathsf{G}^{(n)}},\mathsf{M})\leq\frac{174+\epsilon}{\sqrt{n}}$.
  2. fnum@enumiitem(Edge-based Szemerédi regularity)
    For any $\epsilon>0$, there exists a graph $G^{(k(\epsilon))}$ on $k(\epsilon)=\lceil(174/\epsilon)^{2}\rceil$ edges satisfying $W_{\square}(\mathsf{M},\mathsf{M}_{G^{(k(\epsilon))}})\leq\epsilon$.
Theorem 1.7 allows us to test properties of large graphs by sampling a few of their edges at random. We formally define the graph parameters that can be tested using our framework in Section 4.3. These consist of graph parameters that are continuous with respect to quotient convergence, and include all polynomial graph parameters that do not change when appending isolated vertices to their inputs (Proposition 4.9). In particular, the graph parameter $f(G)=\|G\mathbbm{1}\|_{2}^{2}+\|G^{\top}\mathbbm{1}\|_{2}^{2}$ is quotient-testable in our framework, and can be used to test for the presence of hubs in a growing sequence of graphs. Specifically, if $(G_{n})$ is a quotient-convergent sequence, we show in Example 4.11 that $f(G_{n})\to 0$ precisely when $(G_{n})$ does not have hubs asymptotically, i.e., it converges to a limiting grapheur (5) with $E=0$ and $\sigma=\varsigma=0$. We then describe an edge-based sampling procedure for testing this graph parameter. In contrast, this parameter cannot be tested in the graphons framework; see Section 4.3.
1.1.4 Equipartition-Consistent Random Graph Models (Section 5)
As our final contribution, we observe that random quotients $(\mathsf{G}_{k}[\mathsf{M}])_{k}$ of any grapheur $\mathsf{M}$ are related to each other by equipartitions of their vertex sets. Specifically, we observe that $\rho(d_{k,nk})\mathsf{G}_{nk}[\mathsf{M}]\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}]$ for any equipartition $d_{k,nk}\colon[nk]\to[k]$, i.e., a map with equal-sized fibers. We call such sequences of random graphs equipartition-consistent random graph models, which play a role in our framework analogous to that played by consistent random graph models in graphon theory [lovasz2012random].
Definition 1.8 (Equipartition-consistent models).
A sequence of probability measures $(\mu_{k}\in\mathcal{P}(\Delta^{k\times k}))_{k\in\mathbb{N}}$ is an equipartition-consistent random graph model if $\rho(d_{k,nk})\mathsf{G}_{nk}\overset{d}{=}\mathsf{G}_{k}$ whenever $\mathsf{G}_{nk}\sim\mu_{nk}$ and $\mathsf{G}_{k}\sim\mu_{k}$ and for any equipartition $d_{k,nk}\colon[nk]\to[n]$.
Further, we show that any equipartition-consistent random graph model is derived from a mixture of grapheurs.
Theorem 1.9.
For any equipartition-consistent random graph model $(\mu_{k}\in\mathcal{P}(\Delta^{k\times k}))_{k}$, there is a unique random exchangeable measure $\mathsf{M}$ on $0,1]^{2}$ satisfying $\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}])$ for all $k\in\mathbb{N}$.
Recall that grapheurs were defined as extremal random exchangeable measures in Definition 1.4. By Choquet’s theorem, any random exchangeable measure is therefore a mixture of grapheurs. Finally, we show in Proposition 5.1 that the sequence $(\mathsf{G}_{k}[\mathsf{M}])$ of quotients of a grapheur $\mathsf{M}$ converge back to $\mathsf{M}$ in a certain sense.
1.2 Related Work
There are several well-studied notions of graph limits and random graph models in the literature. We review this body of work and discuss similarities and differences with the present paper.
Dense and degree-bounded graph limits.
The earliest two types of graph limits pertain to sequences of simple graphs that are dense, whose limits are given by graphons [LOVASZ2006933, convergent_seqs1], and degree-bounded, whose limits are given by graphings [benjamini_schramm, aldous2004objective, elek2007note]. We refer the reader to [lovasz2012large] for a survey of both limits. Both of these notions of graph limits are incomparable to ours, as both dense and degree-bounded sequences of simple graphs converge to the same limit in our framework, see Corollary 3.7. Further, dense graph limits can be characterized in terms of quotients [lovasz2012large, §12], called “right convergence” of graphons. However, these quotients are differently normalized compared to ours, and it is the set of quotients that converges in the graphon framework rather than their distribution as in our framework. Despite these distinctions, we highlight parallels between our development and the theory of graphons throughout our paper; in particular, grapheurs can be viewed as dual to graphons in a precise sense, see Section 3.3.
Edge-exchangeable random graphs.
A class of random graph models closely related to the ones we obtain by sampling edges from a grapheur is called edge-exchangeable [crane2016edge, Crane_Dempsey_2019, NIPS2016_1a0a283b, exchangeable_traits, janson2018edge]. These random graphs are obtained by adding random edges to a fixed infinite vertex set (often with repetition, yielding multigraphs), and can yield both sparse and dense random graphs [janson2018edge]. As explained in [janson2010graphons, Rmk. 4.4], every edge-exchangeable multigraph model can be obtained by sampling edges from a random exchangeable measure on $0,1]^{2}$. In fact, the random graphs we obtain by sampling edges in Section 4 are precisely these edge-exchangeable multigraphs, normalized to have unit edge weight, see Remark 4.2. From this perspective, convergence of grapheurs can be interpreted as convergence of the edge-exchangeable multigraph models associated to them, an interpretation we make precise in Proposition 4.1. Nevertheless, to our knowledge this notion of convergence of edge-exchangeable models has not been previously studied in the literature. In particular, the only metric between edge-exchangeable models we are aware of is an $\ell_{1}$ distance between edge probabilities [Crane_Dempsey_2019, Eq. (8)]. Convergence in this metric is too restrictive for our purposes, see Proposition 3.8. Further, our development of convergence via random quotients, which are not edge-exchangeable, allows us to take limits of general weighted graphs, to relate it to convergence of homomorphism densities and other graph parameters, and to test their values.
Graphex processes and limits.
The closest line of work to ours is the study of graphex processes and their associated graph limits and random graph models [caron_fox, veitch2015class, jmlr_graphon_proc, borgs2020identifiability, borgs2019sampling, multigraphex, JANSON2022103549]. Graphexes form a subset of random exchangeable measures on $\mathbb{R}^{2}_{\geq 0}$, and admit a characterization due to Kallenberg similarly to our grapheurs [kallenberg1990exchangeable]. Graphexes give rise to random graph models obtained by sampling edges from them [veitch2015class, borgs2019sampling, veitch2019sampling], analogously to our edge sampling procedure in Section 4.1. There are also several notions of limits for simple graphs and multigraphs, whereby such graphs converge to a limiting graphex [jmlr_graphon_proc, multigraphex]. These notions of limits are conjectured to be inequivalent [JANSON2022103549], and the space of graphexes with either topology is not compact. Some of these limits can be characterized by sampling random subgraphs (whose number of vertices is also random) [borgs2019sampling], while others can be characterized by convergence of differently-normalized homomorphism numbers [jmlr_graphon_proc, §2.5] (the authors of [jmlr_graphon_proc] normalize $\mathrm{hom}(H;G)$ by $(2|E(G)|)^{|V(H)|/2}$, whereas we normalize by $(2|E(G)|)^{|E(H)|}$). Graphex-based graph limits are again incomparable to ours. On the one hand, for a sequence of simple graphs our grapheur-based limits can only distinguish stars from a uniform component (Corollary 3.7), whereas graphex-based limits can capture more general limiting structures (see [JANSON2022103549, §9] or [multigraphex]). On the other hand, grapheur-based graph limits are defined for general sequences of nonnegatively-weighted graphs, for which graphex-based limits have not been defined, allowing us to test properties pertaining to the distribution of edge weight in the graph (see Section 4.3). In fact, grapheurs are particularly well-suited for property testing, since the space of grapheurs we study here is compact, in contrast to the space of graphexes; to our knowledge, there is no characterization of properties that can be tested using graphex-based limits. Another consequence of the compactness of the space of grapheurs is that our notion of graph limits can be characterized in several equivalent ways: by convergence of random quotients, convergence of random edge-sampled graphs, and convergence of homomorphism numbers and related graph parameters.
1.3 Notation and Preliminaries
We denote by $\mathbb{N}=\{0,1,2,\ldots\}$ the set of nonnegative integers. If $n\in\mathbb{N}$, then $n]=\{1,\ldots,n\}$. If $a\leq b$ are real numbers then $a,b]=\{x\in\mathbb{R}:a\leq x\leq b\}$ and $a,b)=[a,b]\setminus\{b\}$. We denote by $\lambda$ the uniform measure on $[0,1]$, by $\lambda^{2}$ the uniform measure on $0,1]^{2}$, and by $\lambda_{D}$ the uniform measure on the diagonal $D=\{(x,x):x\in[0,1]\}$. We denote by $\mathbbm{1}_{n}$ the vector of all-1’s of length $n$, and we omit the subscript $n$ when the size (possibly infinite) of the all-1’s vector is clear from context. We denote by $I_{n}$ the $n\times n$ identity matrix and by $\mathbbm{1}_{n\times n}=\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}$ the $n\times n$ all-1’s matrix.
We identify a (directed, weighted) graph with its adjacency matrix $G\in\mathbb{R}^{n\times n}_{\geq 0}$, where $G_{i,j}$ is the weight of the edge $i\to j$. The collection of all adjacency matrices of graphs with unit total edge weight is denoted by $\Delta^{n\times n}$. In this paper, all edge weights are nonnegative and sum to one unless stated otherwise. A graph is undirected if its adjacency matrix is symmetric. We denote the space of $n\times n$ symmetric matrices by $\mathbb{S}^{n}$, and the set of such matrices with nonnegative entries, which is the set of undirected graphs, by $\mathbb{S}^{n}_{\geq 0}$. An undirected graph $G$ is simple if $\mathrm{diag}(G)=0$ and $G_{i,j}\in\{0,1\}$ for all $i,j$, and we denote the set of simple graphs by $\mathbb{S}^{n}_{\mathrm{sim}}$. A graph $H$ is called a multigraph if it has integer edge weights $H\in\mathbb{N}^{k\times k}$, where we view $H_{i,j}$ as representing the number of parallel edges from $i$ to $j$. Note that a simple graph is a multigraph. We denote by $|E(G)|=\mathbbm{1}^{\top}G\mathbbm{1}/2$ the number of edges in a simple graph $G$. If $H$ is a multigraph, we write $\mathbbm{1}^{\top}H\mathbbm{1}=\|H\|_{1}$ to be the number of (directed) edges in $H$. We assume that multigraphs have no isolated vertices, defined as vertices with no self-loops and no edges incident on them. In other words, if $H\in\mathbb{N}^{k\times k}$ is a multigraph, then for every $i\in[k]$ there is a $j\in[k]$ such that $H_{i,j}>0$ or $H_{j,i}>0$. If $G\in\mathbb{R}^{n\times n}_{\geq 0}$ and $H\in\mathbb{N}^{k\times k}$ with $k\leq n$, we denote by $G^{H}=\prod_{i,j=1}^{k}G_{i,j}^{H_{i,j}}$ the monomial over $\mathbb{R}[G_{i,j}]$ defined by $H$. In this notation, we have $t_{Q}(H;G)=\mathbb{E}(\rho(F_{k,n})G)^{H}$ for a multigraph $H\in\mathbb{N}^{k\times k}$ and graph $G\in\Delta^{n\times n}$.
We denote by $\mathsf{S}_{n}$ the group of permutations of $[n]$, and by $\mathsf{S}_{\infty}$ the group of permutations of $\mathbb{N}$ that fixes all but finitely-many integers. If $\pi\in\mathsf{S}_{n}$ and $G\in\mathbb{R}^{n\times n}_{\geq 0}$ is a graph, we denote by $\pi G\in\mathbb{R}^{n\times n}_{\geq 0}$ the permuted graph with entries $(\pi G)_{i,j}=G_{\pi^{-1}(i),\pi^{-1}(j)}$. This corresponds to relabeling the vertices of $G$ according to $\pi$. Similarly, if $x\in\mathbb{R}^{n}$ is a vector we denote by $\pi x\in\mathbb{R}^{n}$ the permuted vector with entries $(\pi x)_{i}=x_{\pi^{-1}(i)}$. If $G\in\mathbb{R}^{\mathbb{N}\times\mathbb{N}}_{\geq 9}$ and $x\in\mathbb{R}^{\mathbb{N}}$ are infinite matrices and vectors, and $\pi\in\mathsf{S}_{\infty}$, we define $\pi G$ and $\pi x$ analogously.
If $\mu\in\mathcal{P}(S)$ is a probability measure on a compact Polish space $S$, we write $X\sim\mu$ and $\mathrm{Law}(X)=\mu$ to denote a random variable distributed according to $\mu$. If $X,Y$ are two random variables on $S$, we write $X\overset{d}{=}Y$ to denote equality in distribution, i.e., the equality $\mathrm{Law}(X)=\mathrm{Law}(Y)$. If $f$ is a function, we denote $\mathbb{E}_{\mu}f=\mathbb{E}_{X\sim\mu}f(X)$ its expectation with respect to $\mu$. If $(\mu_{n})$ is a sequence of measures in $\mathcal{P}(S)$, we say $(\mu_{n})$ converges weakly to $\mu$ and write $\mu_{n}\to\mu$ if $\mathbb{E}_{\mu_{n}}f\to\mathbb{E}_{\mu}f$ for all continuous $f$. All measures in this paper are supported in a single ambient compact set. In particular, all sequences of measures are tight.
A coupling of random variables $X_{1},\ldots,X_{k}$ taking values in Polish spaces $S_{1},\ldots,S_{k}$, respectively, is a probability distribution $\Gamma\in\mathcal{P}(S_{1}\times\cdots\times S_{k})$ such that if $(X_{1}^{\prime},\ldots,X_{k}^{\prime})\sim\Gamma$ then $X_{i}^{\prime}\overset{d}{=}X_{i}$ for all $i\in[k]$. In this paper we refer to a coupling of $X_{1},\ldots,X_{k}$ simply by the tuple of jointly-distributed random variables $(X_{1}^{\prime},\ldots,X_{k}^{\prime})$. If $X_{1},X_{2}$ are random graphs taking values in $\mathbb{R}^{k\times k}$, we define the Wasserstein-1 distance between them by
$W_{1}(X_{1},X_{2})=\inf_{\begin{subarray}{c}\textrm{random }(X_{1}^{\prime},X_{2}^{\prime})\\ X_{i}^{\prime}\overset{d}{=}X_{i}\end{subarray}}\mathbb{E}\|X_{1}^{\prime}-X_{2}^{\prime}\|_{1}=\sup_{\begin{subarray}{c}f\colon\mathbb{R}^{k\times k}\to\mathbb{R}\\ \textrm{1-Lipschitz}\end{subarray}}|\mathbb{E}f(X_{1})-\mathbb{E}f(X_{2})|,$
where the infimum is over couplings $(X_{1}^{\prime},X_{2}^{\prime})$ of $X_{1}$ and $X_{2}$, the norm $\|\cdot\|_{1}$ is the entrywise $\ell_{1}$ norm, and the supremum is over 1-Lipschitz functions with respect to this norm. Note that a sequence of random graphs $(X_{n})\subseteq\Delta^{k\times k}$ converges weakly to the random graph $X\in\Delta^{k\times k}$ if and only if $\lim_{n}W_{1}(X_{n},X)=0$ by [villani2008optimal, Thm. 6.9].
For a compact Polish space $S$, the space $\mathcal{P}(S)$ endowed with the weak topology is itself compact and Polish. If $\mathsf{M}$ is a random measure, that is, a random variable taking values in the Polish space $\mathcal{P}(S)$, we write $\mathbb{E}_{\mathsf{M}}f$ for the scalar random variable obtained by integrating a deterministic function $f$ against the random measure $\mathsf{M}$. In particular, we write $\mathsf{M}(B)=\mathbb{E}_{\mathsf{M}}\mathbbm{1}_{B}$ to denote the scalar random variable giving the (random) measure of a measurable set $B$, where $\mathbbm{1}_{B}(x)=1$ if $x\in B$ and zero otherwise. A sequence of random measures $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if $\mathrm{Law}(\mathsf{M}_{n})\to\mathrm{Law}(\mathsf{M})$ weakly, where these laws are probability measures over the space of probability measures. Equivalently, the sequence $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if the sequence of scalar random variables $\mathbb{E}_{\mathsf{M}_{n}}f$ converges weakly to $\mathbb{E}_{\mathsf{M}}f$ for all continuous functions $f$ [kallenberg2005probabilistic, Thm. A2.3].
2 Quotient Convergence: Global and Local Views
In this section we prove Theorem 1.2 relating quotient convergence to convergence of homomorphism numbers and quotient densities. Before presenting this proof, we give a few examples illustrating sequences of graphs that are quotient-convergent. As our first example, we show that the sequence of complete graphs is quotient-convergent.
Example 2.1 (Convergence of complete graphs).
To show that the sequence $(\mathbbm{1}_{n\times n})_{n\in\mathbb{N}}$ of complete graphs converges, consider the associated normalized graphs $(G_{n}=\mathbbm{1}_{n\times n}/n^{2}\in\Delta^{n\times n})$ and observe that the random quotients are given as $\rho(F_{k,n})G_{n}=\frac{1}{n^{2}}NN^{\top}$ for a multinomial random variable $N=(N_{1},\ldots,N_{k})\sim\mathrm{Multinom}(n,k,\mathbbm{1}_{k}/k)$. Since $\mathbb{E}\rho(F_{k,n})G_{n}=\frac{1}{k^{2}}\mathbbm{1}_{k\times k}+O(1/n)$ and $\mathrm{Var}[\rho(F_{k,n})G_{n}]_{i,j}=O(1/n)$ for all $i,j\in[k]$, we conclude that $\rho(F_{k,n})G_{n}$ converges weakly to the deterministic matrix $\frac{1}{k^{2}}\mathbbm{1}_{k\times k}$ as $n\to\infty$. Since this holds for all $k\in\mathbb{N}$, we conclude that $(G_{n})$ is quotient-convergent. Note that the quotient densities and homomorphism numbers of this sequence likewise converge. Indeed, for any $H\in\mathbb{N}^{k\times k}$ we have $t_{Q}(H;G_{n})=\mathbb{E}(\rho(F_{k,n})G_{n})^{H}\xrightarrow{n\to\infty}(\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top})^{H}=\frac{1}{k^{2\|H\|_{1}}}$. Similarly, $\mathrm{hom}(H;G_{n})=\frac{n^{k}}{n^{2\|H\|_{1}}}$, which converges to 1 if $H$ is a disjoint union of edges and to 0 otherwise. An analogous set of arguments show that the sequence $(I_{n})$ of $n$ self-loops is quotient-convergent.
As our next example, we consider sequences of (directed) star graphs and we show that these are quotient-convergent as well. Note that stars are sparse graphs, as their edge density converges to zero. Nevertheless, we will see in Section 3 that growing-sized stars have nontrivial limits in our framework, whereas the limits of stars in the context of graphons yield trivial limiting objects. Similarly, stars are not degree-bounded, as the degree of the star center increases with the size of the star, hence they do not have a limit in the framework of graphings. Stars do have limits in the graphex sense, though they differ from ours, see Section 1.2.
Example 2.2 (Convergence of stars).
The sequence of directed star graphs all of whose edges point away from the center may be expressed as $(e_{1}\mathbbm{1}_{n}^{\top})$ for $e_{1}=(1,0,\ldots,0)^{\top}$. This sequence quotient-converges, since random quotients $\rho(F_{k,n})G_{n}$ of the associated normalized graphs $G_{n}=\frac{1}{n}e_{1}\mathbbm{1}_{n}^{\top}$ are given by $\rho(F_{k,n})G_{n}=\frac{1}{n}e_{I}N^{\top}$ for a multinomial vector $N\sim\mathrm{Multinom}(n,k,\mathbbm{1}_{k}/k)$ and a uniformly random index $I\sim\mathrm{Unif}([k])$. Since $N/n$ weakly converges to the deterministic vector $\mathbbm{1}_{k}/k$, we conclude that $\rho(F_{k,n})G_{n}$ converges weakly to the randomly-labeled star $\frac{1}{k}e_{I}\mathbbm{1}_{k}^{\top}$, and hence that $(G_{n})$ is quotient-convergent. Similarly, the quotient densities and homomorphism numbers of this sequence of stars converge, sinces
$\lim_{n\to\infty}t_{Q}(H;G_{n})=\mathbb{E}\left(\frac{1}{k}e_{I}\mathbbm{1}_{k}^{\top}\right)^{H}=\begin{cases}\frac{1}{k^{1+\sum_{j}m_{j}}}&\textrm{if }H=e_{i}m^{\top}\textrm{ for }m\in\mathbb{N}^{k}\textrm{ and }i\in[k],\\ 0&\textrm{otherwise},\end{cases}$
and
$\lim_{n\to\infty}\mathrm{hom}(H;G_{n})=\begin{cases}\lim_{n\to\infty}\frac{n^{k-1}}{n^{\sum_{j}m_{j}}}&\textrm{if }H=e_{i}m^{\top},\\ 0&\textrm{otherwise},\end{cases}=\begin{cases}1&\textrm{if }H=e_{i}m^{\top}\textrm{ with }\sum_{j}m_{j}=k-1,\\ 0&\textrm{otherwise}.\end{cases}$
Analogous arguments show that the sequence of directed star graphs all of whose edges point towards the center quotient-converges, and so does the sequence of undirected stars.
As our final example of quotient-convergence, we consider an infinite weighted graph with edge weights summing to one and we show that any sequence of finite subgraphs obtained from this infinite graph is quotient-convergent.
Example 2.3 (Subgraphs of an infinite weighted graph).
Suppose $G_{\infty}=(G_{i,j})_{i,j\in\mathbb{N}}$ is an infinite two-dimensional array with $G_{i,j}\geq 0$ and $\sum_{i,j}G_{i,j}=1$, and consider the sequence of normalized truncations $G_{n}=\frac{1}{\sum_{i,j=1}^{n}G_{i,j}}(G_{i,j})_{i,j\in[n]}$. This sequence of subgraphs $(G_{n})$ is quotient-convergent. Indeed, let $F_{k}\colon\mathbb{N}\to[k]$ be a uniformly random map (whose images $F_{k}(i)$ are iid uniformly distributed over $[k]$), and consider the random graph $L_{k}=\rho(F_{k})G_{\infty}$ whose entries are given by (1). We claim that the sequence $(\rho(F_{k,n})G_{n})_{n}$ converges weakly to $L_{k}$. To this end, view $G_{n}$ as an infinite array by zero-padding it, and note that $(\rho(F_{k})G_{n},\rho(F_{k})G_{\infty})$ is a coupling of $(\rho(F_{k,n})G_{n},L_{k})$. Using this coupling, we obtain the bound
$W_{1}(\rho(F_{k,n})G_{n},L_{k})\leq\mathbb{E}\|\rho(F_{k})(G_{n}-G_{\infty})\|_{1}\leq\|G_{n}-G_{\infty}\|_{1}\xrightarrow{n\to\infty}0,$
where the second inequality follows from (1). Thus, we conclude that $(G_{n})$ is quotient-convergent, as claimed. Quotient densities and homomorphism numbers of this sequence again converge, with limits $\lim_{n}t_{Q}(H;G_{n})=\mathbb{E}L_{k}^{H}$ and $\lim_{n}\mathrm{hom}(H;G_{n})=\sum_{f\colon[k]\to\mathbb{N}}\prod_{i,j=1}^{k}G_{f(i),f(j)}^{H_{i,j}}$, which is finite since $(G_{i,j})$ is summable.
Finally, we remark that convex combinations of quotient-convergent sequences also converge.
Example 2.4 (Combinations of graphs).
Suppose $(G_{n})$ and $(G_{n}^{\prime})$ are quotient-convergent sequences of graphs. Then for any $\theta\in[0,1]$ the convex combination $(\theta G_{n}+(1-\theta)G_{n}^{\prime})$ of these graphs is quotient-convergent as well. Indeed, for each $k\in\mathbb{N}$, let $L_{k}$ and $L_{k}^{\prime}$ be the weak limits of $\rho(F_{k,n})G_{n}$ and $\rho(F_{k,n})G_{n}^{\prime}$, respectively. For each $n\in\mathbb{N}$, let $(\widetilde{L}_{k},\widetilde{F}_{k,n})$ and $(\widetilde{L}_{k}^{\prime},\widetilde{F}_{k,n}^{\prime})$ be couplings of $(L_{k},F_{k,n})$ and $(L_{k}^{\prime},F_{k,n})$, respectively, attaining the Wasserstein-1 distances $W_{1}(L_{k},\rho(F_{k,n})G_{n})$ and $W_{1}(L_{k}^{\prime},\rho(F_{k,n})G_{n})$. We define a coupling of $\theta L_{k}+(1-\theta)L_{k}^{\prime}$ and $\rho(F_{k,n})(\theta G_{n}+(1-\theta)G_{n}^{\prime})$ as follows. Draw $F_{k,n}$ uniformly at random, and draw $\widetilde{L}_{k}$ from the coupling $(\widetilde{L}_{k},\widetilde{F}_{k,n})$ conditioned on $\widetilde{F}_{k,n}=F_{k,n}$. Draw $\widetilde{L}_{k}^{\prime}$ similarly and independently. Then
$W_{1}(\theta L_{k}+(1-\theta)L_{k}^{\prime},\rho(F_{k,n})(\theta G_{n}+(1-\theta)G_{n}^{\prime}))\leq\mathbb{E}\|\theta\widetilde{L}_{k}+(1-\theta)\widetilde{L}_{k}^{\prime}-\rho(F_{k,n})(\theta G_{n}+(1-\theta G_{n}^{\prime}))\|_{1}$
$\quad\leq\theta W_{1}(L_{k},\rho(F_{k,n})G_{n})+(1-\theta)W_{1}(L_{k}^{\prime},\rho(F_{k,n})G_{n}^{\prime})\xrightarrow{n\to\infty}0.$
This proves the claim. An analogous argument shows that a convex combination of any finitely-many quotient-convergent sequences of graphs is also quotient-convergent.
Having seen a few examples of quotient convergence, we proceed to prove Theorem 1.2. We begin by showing the first part of that theorem on the equivalence between quotient convergence and convergence of quotient densities.
Proposition 2.5.
A sequence of graphs $(G_{n}\in\Delta^{n\times n})$ is quotient-convergent if and only if $(t_{Q}(H;G_{n}))_{n}$ converges for each $H\in\mathbb{N}^{k\times k}$ and $k\in\mathbb{N}$.
Proof.
[…]
This proves the first part of Theorem 1.2. To show the second part relating quotient convergence with convergence of homomorphism numbers, we express homomorphism numbers in terms of quotient densities and vice-versa. To this end, we begin by expressing quotient densities in terms of injective homomorphism numbers using [levin2025deFin, Prop. 6.8]. This expression is stated in terms of a partial order on multigraphs induced by quotients. Specifically, for two multigraphs $K\in\mathbb{N}^{n\times n},H\in\mathbb{N}^{m\times m}$ we say that $K$ is a refinement of $H$, denoted $K\leq_{R}H$, if $H$ is a quotient of $K$, i.e., if there exists a surjective map $f\colon[n]\to[m]$ such that $H=\rho(f)K$. We further let $R_{K,H}$ be the number of such maps $f$. Note that $R_{H,H}=|\mathrm{Aut}(H)|$ for any multigraph $H$. The poset of all multigraphs with the refinement order is a disjoint union of posets, one for each number of edges. The poset of all multigraphs with $n$ edges has a minimal element, namely, the disjoint union of $n$ edges, and a maximal element, namely, a single vertex with $n$ self-loops. We illustrate the poset of all multigraphs with two edges in Figure 3.
Figure 3: Hasse diagram for the refinement poset consisting of multigraphs with two edges. Here there is an upward path from $K$ to $H$ precisely when $K\leq_{R}H$.
The quotient density $t_{Q}(H;G)$ can be expressed in terms of a linear combination of injective homomorphism numbers over all refinements of $H$ [levin2025deFin, Prop. 6.8]:
$t_{Q}(H;G)=\sum_{K\leq_{R}H}|V(H)|^{-|V(K)|}\frac{H!}{K!}\frac{R_{K,H}}{R_{K,K}R_{H,H}}\mathrm{inj}(K;G),$ (10)
where $H!=\prod_{i,j=1}^{k}(H_{i,j}!)$ and similarly for $K!$. Here $\mathrm{inj}(K;G)$ is defined similarly to $\mathrm{hom}(K;G)$ in (2) but with a sum only over injective maps. The expression (10) states that $H$ is a subgraph of some quotient of $G$, i.e., $t_{Q}(H;G)>0$, if and only if some refinement of $H$ is a subgraph of $G$ itself, i.e., $\mathrm{inj}(K;G)>0$.
We are now ready to prove the second part of Theorem 1.2. The following argument is a simple modification of the one from [levin2025deFin, Prop. 6.8].
Proposition 2.6.
A sequence of graphs $(G_{n}\in\Delta^{n\times n})$ is quotient-convergent if and only if $(\mathrm{inj}(H;G_{n}))$ converges for each $H\in\mathbb{N}^{k\times k}$ and $k\in\mathbb{N}$, which is equivalent to $(\mathrm{hom}(H;G_{n}))$ converging for each $H\in\mathbb{N}^{k\times k}$ and $k\in\mathbb{N}$.
Proof.
For each $m\in\mathbb{N}$, define the matrix $M$ indexed by all multigraphs containing $m$ edges whose entries are given by $M_{H,K}=|V(H)|^{-|V(K)|}\frac{H!}{K!}\frac{R_{K,H}}{R_{K,K}R_{H,H}}$ if $K\leq_{R}H$ and $M_{H,K}=0$ otherwise. This is a lower-triangular matrix with respect to the refinement partial order, and has positive diagonal entries $M_{H,H}=(|V(H)|^{|V(H)|}|\mathrm{Aut}(H)|)^{-1}$, hence it is invertible. Thus, we can write
$\mathrm{inj}(H;G)=\sum_{H\leq_{R}K}(M^{-1})_{H,K}t_{Q}(K;G).$ (11)
Based on (11) and (10), we conclude that $(t_{Q}(H;G_{n}))$ converges for all multigraphs $H$ if and only if $(\mathrm{inj}(H;G_{n}))$ converges for all multigraphs $H$. Injective and non-injective homomorphism numbers can be similarly related, with one of the quantities being expressed in terms of linear combinations of the other as in [lovasz2012large, §5.2.3]. For example, we have [levin2025deFin, Eq. (73)] that
$\mathrm{hom}(H;G)=\sum_{H\leq_{R}K}\frac{R_{H,K}}{R_{K,K}}\mathrm{inj}(K;G),$ (12)
where the sum is over multigraphs that $H$ refines. This expression can be inverted as above to obtain injective homomorphism numbers in terms of non-injective homomorphism numbers. This allows us to conclude that convergence of $(\mathrm{inj}(H;G_{n}))$ for all multigraphs $H$ is equivalent to convergence of $(\mathrm{hom}(H;G_{n}))$ for all multigraphs $H$. ∎
Combined together, Propositions 2.5 and 2.6 complete the proof of Theorem 1.2. We illustrate the utility of these results next by expressing quotient convergence of simple graphs in terms of appropriately-normalized homomorphism numbers. The resulting expression allows us to conclude that all dense simple graphs are quotient-convergent; in fact, we prove that all the homomorphism numbers of dense simple graphs converge to the same limit (see also Corollary 3.7 below).
Corollary 2.7.
Let $(G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}})$ be a sequence of simple graphs.
  1. fnum@enumiitem(Convergence of simple graphs)
    The sequence $(G_{n})$ is quotient-convergent if and only if $\left(\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}}\right)$ converges for all simple and connected graphs $H\in\mathbb{S}^{k}_{\mathrm{sim}}$.
  2. fnum@enumiitem(Convergence of dense graphs)
    Suppose $\liminf_{n\to\infty}\frac{|E(G_{n})|}{n^{2}}>0$. Then $(G_{n})$ is quotient-convergent. In fact, for a simple and connected $H\in\mathbb{S}^{k}_{\mathrm{sim}}$, we have
    $\lim_{n\to\infty}\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}}=\begin{cases}1&\textrm{if }H=K_{2},\\ 0&\textrm{otherwise}\end{cases}$
    where $K_{2}$ denotes a graph with a single edge.
Proof.
For the first claim, if $H\in\mathbb{N}^{k\times k}$ we define a strictly upper-triangular $\widetilde{H}\in\{0,1\}^{n\times n}$ by $\widetilde{H}_{i,j}=1$ if $i<j$ and either $H_{i,j}>0$ or $H_{j,i}>0$, and $\widetilde{H}_{i,j}=0$ otherwise. Then $\mathrm{hom}(H;G_{n})=\mathrm{hom}(\widetilde{H},G_{n})$ for all simple graphs $G_{n}$ of all sizes. Therefore, we have
$\mathrm{hom}(H;\tfrac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}})=\mathrm{hom}(\widetilde{H};\tfrac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}})\cdot(2|E(G_{n})|)^{\mathbbm{1}^{\top}(\widetilde{H}-H)\mathbbm{1}}.$
Since $\mathbbm{1}^{\top}\widetilde{H}\mathbbm{1}\leq\mathbbm{1}^{\top}H\mathbbm{1}$, we conclude that $\mathrm{hom}(H;\tfrac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}})$ converges if $\mathrm{hom}(\widetilde{H};\tfrac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}})$ converges, since either $|E(G_{n})|\to\infty$ and $\mathrm{hom}(H;\tfrac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}})\to 0$ or $|E(G_{n})|$ is constant for all large $n$. In turn, defining the simple graph $\widetilde{H}_{\mathrm{sym}}=(\widetilde{H}+\widetilde{H}^{\top})/2$, we have
$\mathrm{hom}\left(\widetilde{H};\frac{G_{n}}{\mathbbm{1}^{\top}G_{n}\mathbbm{1}}\right)=\frac{\mathrm{hom}(\widetilde{H}_{\mathrm{sym}};G_{n})}{(2|E(G_{n})|)^{|E(\widetilde{H}_{\mathrm{sym}})|}}.$
Thus, the sequence $(G_{n})$ quotient-converges if and only if $(\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}})_{n}$ converges for all simple $H$. To see that it suffices to consider simple and connected graphs, note that if $H=H_{1}\sqcup\ldots\sqcup H_{\ell}$ is a disjoint union of connected graphs, then $\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}}=\prod_{i=1}^{\ell}\frac{\mathrm{hom}(H_{i};G_{n})}{|E(G_{n})|^{|E(H_{i})|}}$. This proves the first claim.
For the second claim, if $H=K_{2}$ then $\mathrm{hom}(H;G_{n})/|E(G_{n})|=1$ by definition. Otherwise, suppose $\frac{|E(G_{n})|}{n^{2}}\geq c>0$ for all $n$. Note that $\mathrm{hom}(H;G_{n})\leq 2|E(G_{n})|\cdot n^{|V(H)|-2}$, because the number of graph homomorphisms $V(H)\to V(G_{n})$ mapping all edges of $H$ to edges of $G_{n}$ is upper-bounded by the number of maps $V(H)\to V(G_{n})$ that send one particular edge in $H$ to an edge in $G_{n}$. Since $H$ is connected and contains more than one edge, we have $|V(H)|<2|E(H)|$, in which case
$\frac{\mathrm{hom}(H;G_{n})}{|E(G_{n})|^{|E(H)|}}\leq\frac{2n^{|V(H)|-2}}{|E(G_{n})|^{|E(H)|-1}}\leq\frac{2}{c^{|E(H)|-1}}\cdot n^{|V(H)|-2|E(H)|}\xrightarrow{n\to\infty}0.$
This proves the quotient-convergence of $(G_{n})$ by the first part of this corollary. ∎
We have now seen several examples and characterizations of quotient-convergent sequences. However, so far it is not clear to what limits such sequences converge. In the next section, we derive limits for quotient-convergent graphs in the form of random exchangeable measures on $0,1]^{2}$.
3 Limits of Quotient-Convergent Graph Sequences
This section is devoted to characterizing limits of quotient-convergent graph sequences and understanding their properties. In Section 3.1, we prove Theorem 1.5 characterizing limits of quotient-convergent graph sequences via grapheurs. In Section 3.2 we interpret the different components of limiting grapheurs (5) in terms of combinatorial aspects of the sequence of graphs converging to them. In Section 3.3 we present a (random) duality pairing between grapheurs and graphons, and prove the formal version of Theorem 1.6 given in Theorem 3.10. Finally, Section 3.4 contains some tedious but straightforward proofs omitted from Section 3.1 for clarity of exposition.
3.1 Characterizing Limits via Grapheurs
To prove Theorem 1.5, we show that the random quotients of size $k$ of a finite graph $G$ are equal in distribution to the random graph $\mathsf{G}_{k}[\mathsf{M}_{G}]$ of size $k$ obtained in (8) from the grapheur $\mathsf{M}_{G}$ associated in (4) to $G$.
Proposition 3.1.
For any $n,k\in\mathbb{N}$ and $G\in\Delta^{n\times n}$, we have $\rho(F_{k,n})G\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}_{G}]$.
Proof.
Recall from (4) that $\mathsf{M}_{G}=\sum_{i,j=1}^{n}G_{i,j}\delta_{(T_{i},T_{j})}$ where $T_{1},\ldots,T_{n}$ are iid uniform on $[0,1]$. With probability one, for each $i\in[n]$ there is a unique $j_{i}\in[k]$ such that $T_{i}\in I_{j_{i}}^{(k)}$. Moreover, the index $j_{i}$ is uniformly distributed in $[k]$ by uniformity of $T_{i}$, and the indices $j_{1},\ldots,j_{n}$ are independent of each other since the $T_{i}$ are independent. Define the random map $F\colon[n]\to[k]$ sending $F(i)=j_{i}$ and observe that $F\overset{d}{=}F_{k,n}$ is uniformly random, since each $i\in[n]$ is mapped to an independent and uniformly random element of $[k]$. Further, note that
$(\mathsf{G}_{k}[\mathsf{M}_{G}])_{i,j}=\mathsf{M}_{G}(I_{i}^{(k)}\times I_{j}^{(k)})=\sum_{u,v=1}^{n}G_{u,v}\mathbbm{1}[T_{u}\in I_{i}^{(k)},T_{v}\in I_{j}^{(k)}]=\sum_{\begin{subarray}{c}u\in F^{-1}(i)\\ v\in F^{-1}(j)\end{subarray}}G_{u,v}=(\rho(F)G)_{i,j}.$
Hence, we conclude that $\mathsf{G}_{k}[\mathsf{M}_{G}]=\rho(F)G\overset{d}{=}\rho(F_{k,n})G$, as claimed. ∎
Consequently, we also have $t_{Q}(H;G)=t_{Q}(H;\mathsf{M}_{G})$, where the latter is given by (9). Proposition 3.1 implies that a sequence $(G_{n})$ is quotient-convergent if and only if the sequence of random graphs $(\mathsf{G}_{k}[\mathsf{M}_{G_{n}}])_{n}$ constructed from the associated grapheurs $(\mathsf{M}_{G_{n}})$ converge weakly for each $k$. Since $\Delta^{k\times k}$ is compact, this is equivalent to convergence of $(\mathsf{G}_{k}[\mathsf{M}_{G_{n}}])_{n}$ in the Wasserstein-1 metric $W_{1}$ with respect to the entrywise $\ell_{1}$ norm on $\mathbb{R}^{k\times k}$ for each $k$ (see Section 1.3). Motivated by this observation, we relate our $W_{\square}$-metric (7) to this $W_{1}$-metric.
Proposition 3.2.
For any two grapheurs $\mathsf{M},\mathsf{N}\in\mathcal{M}$ and any $k\in\mathbb{N}$, we have
$W_{1}(\mathsf{G}_{k}[\mathsf{M}],\mathsf{G}_{k}[\mathsf{N}])\leq k^{2}W_{\square}(\mathsf{M},\mathsf{N})\quad\textrm{and }\quad W_{\square}(\mathsf{M},\mathsf{N})\leq W_{1}(\mathsf{G}_{k}[\mathsf{M}],\mathsf{G}_{k}[\mathsf{N}])+\frac{4}{k}.$
The proof is straightforward but somewhat technical, as it is based on approximating any rectangle appearing in our expression (7) for $W_{\square}$ by squares in a uniform grid. We defer the proof to Section 3.4. As an immediate consequence, we conclude that a sequence of graphs $(G_{n})$ is quotient-convergent if and only if the associated sequence of grapheurs $(\mathsf{M}_{G_{n}})$ is Cauchy in the $W_{\Box}$-metric (7). We now observe that the metric space $(\mathcal{M},W_{\square})$ is compact, because the metric $W_{\square}$ metrizes weak convergence in $\mathcal{M}$ and $\mathcal{M}$ is weakly compact; see Section 1.3 for a review of weak convergence.
Proposition 3.3.
The space $\mathcal{M}$ of grapheurs (6) is weakly compact. In particular, any Cauchy sequence $(\mathsf{M}_{n})\subseteq\mathcal{M}$ has a limit $\mathsf{M}\in\mathcal{M}$. Moreover, the following statements are equivalent for $(\mathsf{M}_{n})\subseteq\mathcal{M}$ and $\mathsf{M}\in\mathcal{M}$.
  1. 1.
    We have $\lim_{n}\mathsf{M}_{n}=\mathsf{M}$ weakly.
  2. 2.
    We have that $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$.
  3. 3.
    For each $k\in\mathbb{N}$, we have that $\lim_{n}\mathsf{G}_{k}[\mathsf{M}_{n}]=\mathsf{G}_{k}[\mathsf{M}]$ weakly.
  4. 4.
    For each multigraph $H\in\mathbb{N}^{k\times k}$, we have $\lim_{n}t_{Q}(H;\mathsf{M}_{n})=t_{Q}(H;\mathsf{M})$.
In particular, we have $\lim_{n}W_{\square}(\mathsf{M}_{G_{n}},\mathsf{M})=0$ for a sequence $(G_{n}\in\Delta^{n\times n})$ if and only if $\lim_{n}t_{Q}(H;G_{n})=t_{Q}(H;\mathsf{M})$ for all multigraphs $H\in\mathbb{N}^{k\times k}$, as can be seen by combining Proposition 3.1 and Proposition 3.3(4). Once again, the proof is straightforward but involved, and we defer it to Section 3.4. Combining the preceding propositions, we are ready to prove Theorem 1.5.
Proof (Theorem 1.5)..
To prove the first statement, note that the sequence $(G_{n})$ is quotient-convergent if and only if $(\mathsf{G}_{k}[\mathsf{M}_{G_{n}}])$ is Cauchy in $W_{1}$ for each $k\in\mathbb{N}$ by Definition 1.1. By Proposition 3.2, this is equivalent to $(\mathsf{M}_{G_{n}})$ being Cauchy in the $W_{\square}$-metric, as we remark previously. By Proposition 3.3, this is further equivalent to the existence of a limit $\mathsf{M}\in\mathcal{M}$ for $(\mathsf{M}_{G_{n}})$ in this metric, thus proving the first statement.
The second statement is a consequence of the compactness of $\mathcal{M}$, which follows from Proposition 3.3.
Finally, we prove the third statement by explicitly constructing a convergent sequence $(G_{n})$ of graphs for each grapheur $\mathsf{M}\in\mathcal{M}$. Suppose $\mathsf{M}$ is given by (5) with $E,\sigma,\varsigma,\theta,\vartheta\geq 0$ and iid uniform $(T_{i})_{i\in\mathbb{N}}$. Define the sequence of graphs $(G_{n}\in\Delta^{n\times n})$ with entries
$(G_{n})_{i,j}=\frac{1}{s_{n}}\left(E_{i,j}+\tfrac{1}{n}\sigma_{i}+\tfrac{1}{n}\varsigma_{j}+\tfrac{1}{n^{2}}\theta+\tfrac{1}{n}\vartheta\mathbbm{1}_{i=j}\right),\qquad s_{n}=\sum_{i,j=1}^{n}E_{i,j}+\sum_{i=1}^{n}(\sigma_{i}+\varsigma_{i})+\theta+\vartheta,$ (13)
where we note that $s_{n}\nearrow 1$ and we may assume $s_{n}>0$ for all $n$ by reordering the rows and columns of $E,\sigma,\varsigma$ if needed. Each summand in (13) quotient-converges to the corresponding summand in (5) as shown by combining Proposition 3.3 with the examples of Section 2. Moreover, by Example 2.4 the convex combination of these summands in (13) converges to the convex combination of the summands in (5), which is precisely the grapheur $\mathsf{M}$. Thus, the third statement in Theorem 1.5 is proved. ∎
In particular, the sequence of limits of random quotients of a convergent graph sequence fully determines its limit.
Corollary 3.4.
Two quotient-convergent sequences $(G_{n}^{(1)})$ and $(G_{n}^{(2)})$ have the same limit if and only if $\lim_{n}\rho(F_{k,n})G_{n}^{(1)}\overset{d}{=}\lim_{n}\rho(F_{k,n})G_{n}^{(2)}$ for all $k\in\mathbb{N}$.
Proof.
From Proposition 3.3, the two sequences have the same limit $\mathsf{M}$ if and only if $\lim_{n}\rho(F_{k,n})G_{n}^{(i)}\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}]$ for both $i=1,2$. ∎
We end this section by noting that valid polynomial inequalities in quotient densities precisely correspond to valid inequalities over grapheurs. We hope that grapheurs can be used to study such inequalities in future work, dually to the use of graphons in the study of inequalities in homomorphism densities [lovasz2012large, §16].
Corollary 3.5 (Inequalities in quotient densities).
Suppose $H_{1},\ldots,H_{k}$ are multigraphs and $f\colon\mathbb{R}^{k}\to\mathbb{R}$ is a continuous function. We have $f(t_{Q}(H_{1};G),\ldots,t_{Q}(H_{k};G))\geq 0$ for all $G\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$ if and only if $f(t_{Q}(H_{1};\mathsf{M}),\ldots,t_{Q}(H_{k};\mathsf{M}))\geq 0$ for all grapheurs $\mathsf{M}\in\mathcal{M}$.
Proof.
Since $t_{Q}(H;G)=t_{Q}(H;\mathsf{M}_{G})$, if $f(t_{Q}(H_{1};\mathsf{M}),\ldots,t_{Q}(H_{k};\mathsf{M}))\geq 0$ for all grapheurs $\mathsf{M}\in\mathcal{M}$ then setting $\mathsf{M}=\mathsf{M}_{G}$ for $G\in\Delta^{n\times n}$ yields $f(t_{Q}(H_{1};G),\ldots,t_{Q}(H_{k};G))\geq 0$. Conversely, suppose $f(t_{Q}(H_{1};G),\ldots,\allowbreak t_{Q}(H_{k};G))\geq 0$ for all $G\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$. For any $\mathsf{M}\in\mathcal{M}$, let $(G_{n})$ be a sequence of finite graphs that quotient-converges to $\mathsf{M}$. Then $t_{Q}(H;G_{n})\to t_{Q}(H;\mathsf{M})$ for all multigraphs $H$, hence by continuity of $f$ we have $f(t_{Q}(H_{1};\mathsf{M}),\ldots,t_{Q}(H_{k};\mathsf{M}))=\lim_{n}f(t_{Q}(H_{1};G_{n}),\ldots,t_{Q}(H_{k};G_{n}))\geq 0$. ∎
3.2 Structure of Grapheurs
Theorem 1.5 shows that limits of quotient-convergent sequences correspond to grapheurs in $\mathcal{M}$. Each such grapheur is described by a tuple $(E,\sigma,\varsigma,\theta,\vartheta)$ as in (5), and the goal of this section is to interpret each element of the tuple in terms of the combinatorics of the convergent sequence of graphs. Informally, the five parameters describing quotient limits correspond to the following properties.
  • The array $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ describes a discrete collection of hubs and the fraction of edge weight between them. Specifically, each entry corresponds to an edge in a sequence of graphs that carries an asymptotically positive fraction of the total edge weight.
  • The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars. Specifically, each entry corresponds to the fraction of in-degree or out-degree, respectively, of a vertex that comes from all edges with asymptotically vanishing edge weights.
  • The parameters $\theta$ and $\vartheta$ capture the remaining edge weight in the graph, coming from all edges between vertices with asymptotically vanishing degrees and their self-loops, respectively.
These properties can be seen in all the examples of Section 2. The following proposition proves these properties in general.
Proposition 3.6.
Suppose $(G_{n}\subseteq\Delta^{n\times n})$ quotient-converges to a grapheur $\mathsf{M}$ that is described by the tuple $(E,\sigma,\varsigma,\theta,\vartheta)$ as in (5). View each $G_{n}$ as an element of $\mathbb{R}^{\mathbb{N}\times\mathbb{N}}$ by zero-padding it with infinitely-many rows and columns. Then there exist permutations $\pi_{n}\in\mathsf{S}_{\infty}$ of $\mathbb{N}$ fixing all but finitely-many integers such that $\lim_{n}\|E-\pi_{n}G_{n}\|_{\infty}=0$, $\lim_{n}\|E\mathbbm{1}+\sigma-\pi_{n}G_{n}\mathbbm{1}\|_{\infty}=0$, and $\lim_{n}\|E^{\top}\mathbbm{1}+\varsigma-\pi_{n}G_{n}^{\top}\mathbbm{1}\|_{\infty}=0$. Furthermore, we have $\vartheta=\lim_{n}\mathrm{Tr}(G_{n})-\mathrm{Tr}(E)$.
Proposition 3.6 justifies the intuition outlined above: Up to a relabelling of the vertices of $G_{n}$, the entries of $E$ are entrywise limits of the edge weights of $G_{n}$, the entries of $\sigma,\varsigma$ are limits of the remaining degrees of vertices in $(G_{n})$ after removing edges with positive limiting weights, and the parameters $\vartheta,\theta$ capture the remaining edge weights in self loops and the rest of the edges, respectively. We prove Proposition 3.6 in Section 4.4 using an edge-sampling perspective on quotient convergence and a dual Szemerédi regularity lemma.
An immediate consequence of Proposition 3.6 is that all dense and degree-bounded sequences of simple graphs have the same limit.
Corollary 3.7.
Suppose $(G_{n}\in\mathbb{S}^{n})$ is a sequence of undirected graphs such that the sequence $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1})$ quotient-converges.
  1. 1.
    The limit is described by $(E,\sigma,\varsigma,\theta,\vartheta)$ with $\sigma=\varsigma$ and $E^{\top}=E$.
  2. 2.
    If each $G_{n}$ has no self-loops, then $\mathrm{diag}(E)=\vartheta=0$.
  3. 3.
    If $G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}}$ is a sequence of simple graphs with $|E(G_{n})|\to\infty$, and is either dense or degree-bounded, then the sequence $(\frac{G_{n}}{2|E(G_{n})|})$ quotient converges to the grapheur $\lambda^{2}$.
Proof.
The first two claims follow directly from Proposition 3.6, since for any permutation $\pi_{n}\in\mathsf{S}_{\infty}$ the matrix $\pi_{n}G_{n}$ is, respectively, symmetric and has zero diagonal if $G_{n}$ is symmetric and has no self-loops. For the third claim, since $G_{n}$ is symmetric, we have $G_{n}^{\top}\mathbbm{1}=G_{n}\mathbbm{1}=\deg(G_{n})$ is the vector of degrees of the vertices in $G_{n}$. If $(G_{n})$ is dense, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{c}{n}\xrightarrow{n\to\infty}0$ for some constant $c>0$. If $(G_{n})$ is degree-bounded, with all degrees bounded by $k$, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{k}{2|E(G_{n})|}\xrightarrow{n\to\infty}0$. In either case, Proposition 3.6 shows that $\lim_{n}\frac{G_{n}}{2|E(G_{n})|}=\lambda^{2}$ since $\mathrm{diag}(G_{n})=0$ for all $n$. ∎
We end this section by giving a sufficient condition for the limit of a quotient-convergent sequence $(G_{n})$ to consist entirely of discrete hubs, i.e., such that $\sigma=\varsigma=0$ and $\theta=\vartheta=0$ in (5). In particular, this sufficient condition implies that the random quotients $(\rho(F_{k,n})G_{n})$ and the grapheurs $(\mathsf{M}_{G_{n}})$ converge not just weakly but also in total variation. Such total variation convergence was shown for graphs sampled both from graphons [lovasz2012large, Cor. 10.25] and from graphexes [borgs2019sampling, §5], and was the primary mode of convergence studied for edge-exchangeable random graph models [Crane_Dempsey_2019]. However, this condition is too restrictive for our purposes in this paper; for example, it does not apply to the sequence of stars in Example 2.2. In the following result, if $\mu,\nu$ are two discrete measures on a space $S$ then their total variation distance is $\mathrm{dist}_{\mathrm{TV}}(\mu,\nu)=\frac{1}{2}\sum_{x\in S}|\mu(\{x\})-\nu(\{x\})|$.
Proposition 3.8.
Suppose $(G_{n}\in\Delta^{n\times n})$ and $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ satisfy $\lim_{n}\|\pi_{n}G_{n}-E\|_{1}=0$ for some permutations $\pi_{n}\in\mathsf{S}_{n}$. Then $(G_{n})$ is quotient-convergent to $\mathsf{M}=\sum_{i,j}E_{i,j}\delta_{(T_{i},T_{j})}$.
Proof.
Draw iid uniform $(T_{i})$ and consider the coupling $\mathsf{M}_{G_{n}}^{\prime}=\sum_{i,j}(\pi_{n}G_{n})_{i,j}\delta_{(T_{i},T_{j})}$, $\mathsf{M}^{\prime}=\sum_{i,j}E_{i,j}\delta_{(T_{i},T_{j})}$, using the same locations for both. Then with probability 1 we have $\mathrm{dist}_{\mathrm{TV}}(\mathsf{M}^{\prime},\mathsf{M}_{G_{n}}^{\prime})\leq\|\pi_{n}G_{n}-E\|_{1}/2\to 0$ since the atoms $(T_{i},T_{j})$ are distinct almost surely. By (8), we then have $W_{1}(\mathsf{G}_{k}[\mathsf{M}],\mathsf{G}_{k}[\mathsf{M}_{G_{n}}])\leq\mathbb{E}\|\mathsf{G}_{k}[\mathsf{M}^{\prime}]-\mathsf{G}_{k}[\mathsf{M}^{\prime}_{G_{n}}]\|_{1}\leq k^{2}\mathbb{E}\mathrm{dist}_{\mathrm{TV}}(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{G_{n}})\to 0$ for each $k$, proving quotient convergence. ∎
3.3 Duality between Grapheurs and Graphons
In this section, we characterize convergence of grapheurs in terms of graphons and vice-versa, exhibiting a facet of the duality between the two graph limit theories. Since graphon theory is most simply expressed in the undirected case (see [diaconis2007graph, §8] for the directed case), we restrict ourselves to this setting in the present section.
We begin by recalling a few basic notions from graphon theory. Let $\widetilde{\mathcal{W}}=\{W\colon[0,1]^{2}\to[0,1]\textrm{ measurable},\allowbreak\textrm{ symmetric}\}$ be the space of $[0,1]$-valued kernels. We define two kernels $W$ and $U$ to be equivalent, denoted $W\sim U$, if there exist (Lebesgue-)measure-preserving maps $\varphi,\psi\colon[0,1]\to[0,1]$ satisfying $W\circ(\varphi,\varphi)=U\circ(\psi,\psi)$. This is the weak isomorphism between kernels commonly defined in the dense graph limits literature [lovasz2012large, Cor. 10.35]. We denote by $\mathcal{W}=\widetilde{\mathcal{W}}$$\sim$ the quotient space under this equivalence relation. Elements of $\mathcal{W}$ are called graphons [lovasz2012large, §7.1]. A graphon $W\in\mathcal{W}$ defines random simple graphs of each size by setting $\mathbb{G}_{k}[W]=(\mathsf{B}_{T_{i},T_{j}}^{(W)})_{i,j\in[k]}$ where $\mathsf{B}_{T_{i},T_{j}}^{(W)}\sim\mathrm{Ber}(W(T_{i},T_{j}))$ are drawn independently for $i<j$ while $\mathsf{B}_{T_{j},T_{i}}^{(W)}=\mathsf{B}_{T_{i},T_{j}}^{(W)}$ for $i>j$ and $\mathsf{B}_{T_{i},T_{i}}^{(W)}=0$. Here $T_{1},\ldots,T_{k}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ as before. A sequence $(W_{n})\subseteq\mathcal{W}$ converges if for each $k$, the sequence of random graphs $(\mathbb{G}_{k}[W_{n}])_{k}$ converges weakly for each kk. [^4] Graphon convergence is usually defined by total variation convergence of these random graphs, which is equivalent to their weak convergence since they are supported on the finite set $\mathbb{S}^{k}_{\mathrm{sim}}$. This is equivalent to convergence of homomorphism densities and to convergence in cut metric [lovasz2012large, §10]. Finally, a finite graph $G\in\mathbb{S}^{n}_{\mathrm{sim}}$ defines a step graphon $W_{G}(x,y)=G_{i,j}$ if $x\in I_{i}^{(n)}\times I_{j}^{(n)}$ where $I_{i}^{(n)}=[(i-1)/n,i/n)$.
[…]
$\langle\mathsf{M},W\rangle=$ $\sum_{i,j\in\mathbb{N}}E_{i,j}\mathsf{B}_{T_{i},T_{j}}^{(W)}+2\sum_{i\in\mathbb{N}}\sigma_{i}\int_{0}^{1}W(T_{i},y)\,dy+\theta\int_{[0,1]^{2}}W(x,y)\,dx\,dy.$ (14)
Note that this random variable exists and is bounded by $|\langle\mathsf{M},W\rangle|\leq 1$ because both $(E_{i,j})$ and $(\sigma_{i})$ are summable. Also note that $\langle\mathsf{M},W_{G}\rangle=\mathbb{E}_{\mathsf{M}}W_{G}=\langle G,\mathsf{G}_{k}[\mathsf{M}]\rangle$ if $G\in\mathbb{S}^{k}_{\mathrm{sim}}$, since $W_{G}$ takes values in $\{0,1\}$ so $\mathsf{B}_{T_{i},T_{j}}^{(W_{G})}=W_{G}(T_{i},T_{j})$. Moreover, for finite graphs $G_{1}\in\Delta^{n\times n},~G_{2}\in\mathbb{S}^{m}_{\mathrm{sim}}$, we have
$\langle\mathsf{M}_{G_{1}},W_{G_{2}}\rangle=\langle\rho(F_{m,n})G_{1},G_{2}\rangle=\langle G_{1},\rho(F_{m,n})^{\star}G_{2}\rangle,$
which can equivalently be viewed as comparing a random quotient of $G_{1}$ to $G_{2}$ or comparing a random subgraph of $G_{2}$ to $G_{1}$ (see Section 1). We now prove that the random variable (14) only depends on the equivalence class of a graphon $W$.
Proposition 3.9.
For any $W,U\in\widetilde{\mathcal{W}}$, if $W\sim U$ then $\langle\mathsf{M},W\rangle\overset{d}{=}\langle\mathsf{M},U\rangle$ for any $\mathsf{M}\in\mathcal{M}_{\mathrm{sym}}$.
Proof.
Suppose $\mathsf{M}\in\mathcal{M}_{\mathrm{sym}}$ is defined by $(E,\sigma=\varsigma,\theta,\vartheta=0)$ as in (5). Since $W\sim U$, there are measure-preserving maps $\varphi,\psi\colon[0,1]\to[0,1]$ satisfying $W\circ(\varphi,\varphi)=U\circ(\psi,\psi)$. We now define a coupling of $\langle\mathsf{M},W\rangle$ and $\langle\mathsf{M},U\rangle$ under which the two are equal almost surely. Sample locations $(T_{i})$ iid uniformly on $[0,1]$, and note that both $(\varphi(T_{i})),(\psi(T_{i}))\overset{iid}{\sim}\mathrm{Unif}([0,1])$ because $\varphi$ and $\psi$ are measure-preserving. Note that, conditioned on the locations $(T_{i})$ above, we have $\mathsf{B}_{\varphi(T_{i}),\varphi(T_{j})}^{(W)}\overset{d}{=}\mathsf{B}_{\psi(T_{i}),\psi(T_{j})}^{(U)}$, since they are both Bernoulli random variables with success probability $W(\varphi(T_{i}),\varphi(T_{j}))=U(\psi(T_{i}),\psi(T_{j}))$. Letting $\mathsf{B}_{i,j}$ be a sample from the above distribution independently for each $i<j$, setting $\mathsf{B}_{i,j}=\mathsf{B}_{j,i}$ for $i>j$, and $\mathsf{B}_{i,i}=0$ for all $i$, we define the coupling of $\langle\mathsf{M},W\rangle$ and $\langle\mathsf{M},U\rangle$ by using (14) with these common samples $(\mathsf{B}_{i,j})$ for both, and note that under this coupling they are equal almost surely. Thus, we have $\langle\mathsf{M},W\rangle\overset{d}{=}\langle\mathsf{M},U\rangle$ as desired. ∎
Thus, the pairing (14) defines a pairing from $\mathcal{M}_{\mathrm{sym}}\times\mathcal{W}$ to bounded scalar random variables. Importantly, we can characterize convergence in $\mathcal{M}_{\mathrm{sym}}$ in terms of graphons and characterize graphon convergence in $\mathcal{W}$ in terms of grapheurs. The following is the formal version of Theorem 1.6 from Section 1.
Theorem 3.10.
We have $\lim_{n}\mathsf{M}_{n}=\mathsf{M}$ in $\mathcal{M}_{\mathrm{sym}}$ if and only if $\langle\mathsf{M}_{n},W_{G}\rangle\to\langle\mathsf{M},W_{G}\rangle$ weakly for all simple graphs $G$. We also have $\lim_{n}W_{n}\to W$ in $\mathcal{W}$ if and only if $\langle\mathsf{M}_{G},W_{n}\rangle\to\langle\mathsf{M}_{G},W\rangle$ weakly for all undirected graphs $G$ with unit edge weight and no self-loops.
Proof.
For the first claim, suppose $\langle\mathsf{M}_{n},W_{G}\rangle\to\langle\mathsf{M},W_{G}\rangle$ for all $G\in\mathbb{S}^{k}_{\mathrm{sim}}$ and all $k\in\mathbb{N}$. Therefore, $\mathbb{E}_{\mathsf{M}_{n}}W_{G}\to\mathbb{E}_{\mathsf{M}}W_{G}$ for any such $G$. Taking linear combinations of such $W_{G}$, we get $\mathbb{E}_{\mathsf{M}_{n}}W\to\mathbb{E}_{\mathsf{M}}W$ for any $W\colon[0,1]^{2}\to\mathbb{R}$ that is piecewise-constant on the $k\times k$ uniform grid and equals zero on the diagonal for some $k\in\mathbb{N}$. For any continuous and symmetric $f\colon[0,1]^{2}\to\mathbb{R}$, define $f_{k}=\sum_{i,j=1}^{k}f_{i,j}^{(k)}\mathbbm{1}_{I_{i}^{(k)}\times I_{j}^{(k)}}$ where if $i\neq j$ we set $f_{i,j}^{(k)}=f(\frac{i}{k}-\frac{1}{2k},\frac{j}{k}-\frac{1}{2k})$ to the value of $f$ at the center of the rectangle $I_{i}^{(k)}\times I_{j}^{(k)}$, and if $i=j$ we set $f_{i,i}^{(k)}=0$. Since $f$ is continuous on a compact set, it is uniformly continuous, so for any $\epsilon>0$ we have $|f(x)-f_{k}(x)|\leq\epsilon$ for all $x\in[0,1]^{2}\setminus\bigcup_{i=1}^{k}(I_{i}^{(k)}\times I_{i}^{(k)})$. Therefore, for any grapheur $\mathsf{M}^{\prime}\in\mathcal{M}_{\mathrm{sym}}$ we have
$\mathbb{E}_{\mathsf{M}^{\prime}}|f-f_{k}|\leq\epsilon+\mathbb{E}\mathsf{M}^{\prime}\left(\bigcup_{i=1}^{k}I_{i}^{(k)}\times I_{i}^{(k)}\right)=\epsilon+\frac{1}{k},$
for all large $k$ since $\mathbb{E}\mathsf{M}^{\prime}=\lambda^{2}$ for $\mathsf{M}^{\prime}\in\mathcal{M}_{\mathrm{sym}}$. Thus, we have
$W_{1}(\mathbb{E}_{\mathsf{M}_{n}}f,\mathbb{E}_{\mathsf{M}}f)\leq 2(\epsilon+k^{-1})+W_{1}(\mathbb{E}_{\mathsf{M}_{n}}f_{k},\mathbb{E}_{\mathsf{M}}f_{k})\xrightarrow{n\to\infty}2(\epsilon+k^{-1}),$
since $f_{k}$ is piecewise-constant on the $k\times k$ uniform grid and has zero values on its diagonal. We conclude by taking $k\to\infty$ that $\mathbb{E}_{\mathsf{M}_{n}}f\to\mathbb{E}_{\mathsf{M}}f$ weakly for any continuous $f\colon[0,1]^{2}\to\mathbb{R}$, and hence that $\mathsf{M}_{n}\to\mathsf{M}$ weakly. Conversely, suppose $\mathsf{M}_{n}\to\mathsf{M}$ weakly, and let $W=W_{G}$ for $G\in\mathbb{S}^{k}_{\mathrm{sim}}$. Then
$\langle\mathsf{M}_{n},W_{G}\rangle=\mathbb{E}_{\mathsf{M}_{n}}W_{G}=\langle G,\mathsf{G}_{k}[\mathsf{M}_{n}]\rangle\to\langle G,\mathsf{G}_{k}[\mathsf{M}]\rangle=\mathbb{E}_{\mathsf{M}}W_{G}=\langle\mathsf{M},W_{G}\rangle,$
weakly, since $\mathsf{G}_{k}[\mathsf{M}_{n}]\to\mathsf{G}_{k}[\mathsf{M}]$ weakly for each $k\in\mathbb{N}$. This proves the first claim.
For the second claim, suppose $\langle\mathsf{M}_{G},W_{n}\rangle\to\langle\mathsf{M}_{G},W\rangle$ weakly for all graphs $G\in\mathbb{S}^{k}\cap\Delta^{k\times k}$ with $\mathrm{diag}(G)=0$, i.e., undirected weighted graphs with no self-loops and total edge weight of 1. We then have
$\langle G,\mathbb{G}_{k}[W_{n}]\rangle=\langle\mathsf{M}_{G},W_{n}\rangle\to\langle\mathsf{M}_{G},W\rangle=\langle G,\mathbb{G}_{k}[W]\rangle,$
weakly. Therefore, we get $\langle G,\mathbb{G}_{k}[W_{n}]\rangle\to\langle G,\mathbb{G}_{k}[W]\rangle$ for all $G\in\mathbb{S}^{k}$ by linearity and the fact that $\mathbb{G}_{k}[W_{n}]$ all have zero diagonal. This implies $\mathbb{G}_{k}[W_{n}]\to\mathbb{G}_{k}[W]$ weakly by Cramér–Wold [kallenberg1997foundations, Cor. 6.5]. Since this holds for all $k\in\mathbb{N}$, we conclude that $W_{n}\to W$. Conversely, suppose $W_{n}\to W$. Then $\langle\mathsf{M}_{G},W_{n}\rangle=\langle G,\mathsf{G}_{k}[W_{n}]\rangle\to\langle G,\mathsf{G}_{k}[W]\rangle=\langle\mathsf{M}_{G},W\rangle$ for any $G\in\mathbb{S}^{k}\cap\Delta^{k\times k}$. ∎
We conjecture that the above duality pairing is, in fact, bi-continuous.
Conjecture 3.11.
If $(\mathsf{M}_{n})\subseteq\mathcal{M}_{\mathrm{sym}}$ and $(W_{n})\subseteq\mathcal{W}$ converge to $\mathsf{M}$ and $W$, respectively, then $\langle\mathsf{M}_{n},W_{n}\rangle\to\langle\mathsf{M},W\rangle$ weakly.
In particular, this conjecture implies that $\mathsf{M}_{n}\to\mathsf{M}$ if and only if $\langle\mathsf{M}_{n},W\rangle\to\langle\mathsf{M},W\rangle$ for all graphons $W\in\mathcal{W}$, and $W_{n}\to W$ if and only if $\langle\mathsf{M},W_{n}\rangle\to\langle\mathsf{M},W\rangle$ for all grapheurs $\mathsf{M}\in\mathcal{M}_{\mathrm{sym}}$, a strengthening of Theorem 3.10.
Theorem 3.10 exhibits a concrete, formal duality between quotient convergence and graphon convergence. This duality can be extended to general grapheurs and more general kernels on $0,1]^{2}$, but the extension of graphon convergence to this setting is more complicated [diaconis2007graph, §8]. We therefore leave this and other extensions of this duality for future work.
3.4 Missing Proofs from Section 3
In this section, we give all the proofs deferred from Section 3.1 and Section 3.2, namely, Propositions 3.2 and 3.3.
Proof of Proposition 3.2
Many of the proofs in this section are based on standard approximations of rectangles, and continuous functions by step functions on a uniform grid and vice-versa. Such approximations allow us to compare the measures of these objects under our random exchangeable measures using the following lemma.
Lemma 3.12.
Fix any grapheur $\mathsf{M}\in\mathcal{M}$.
  1. 1.
    There is $\theta\in[0,1]$ satisfying $\mathbb{E}\mathsf{M}=\theta\lambda^{2}+(1-\theta)\lambda_{D}$.
  2. 2.
    For any Borel set $B\subseteq[0,1]^{2}$, we have $\mathbb{E}\mathsf{M}(B)\leq\max\{\lambda^{2}(B),\lambda_{D}(B)\}$. In particular, if $\lambda^{2}(B)=\lambda_{D}(B)=0$, then $\mathsf{M}(B)=0$ almost surely.
Proof.
For the first claim, since $\mathsf{M}$ is exchangeable we conclude that $\mu=\mathbb{E}\mathsf{M}$ is a (deterministic) exchangeable measure on $0,1]^{2}$, meaning that $\mu\circ(\sigma,\sigma)=\mu$ for all (Lebesgue)-measure preserving bijections $\sigma\colon[0,1]\to[0,1]$. The space of such measures is two-dimensional and is spanned by $\lambda^{2},\lambda_{D}$ by [kallenberg2005probabilistic, Thm. 9.12], so $\mu=\theta\lambda+\vartheta\lambda_{D}$ for some $\theta,\vartheta\in\mathbb{R}$. Since $\mathsf{M}$ is a probability measure almost surely, its mean $\mu$ is a probability measure as well, necessitating $\theta+\vartheta=1$ and $\theta,\vartheta\geq 0$. This proves the first claim.
The second claim immediately follows from the first, noting in particular that if $\lambda^{2}(B)=\lambda_{D}(B)=0$ then the scalar random variable $\mathsf{M}(B)$ is nonnegative with a vanishing mean. ∎
To prove Proposition 3.2, we also need to relate couplings of random measures to couplings of random graphs derived from them.
Lemma 3.13.
For any $\mathsf{M}_{1},\mathsf{M}_{2}\in\mathcal{M}$ and coupling $(\mathsf{G}_{1},\mathsf{G}_{2})$ of $(\mathsf{G}_{k}[\mathsf{M}_{1}],\mathsf{G}_{k}[\mathsf{M}_{2}])$ for some $k\in\mathbb{N}$, there is a coupling $(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})$ of $(\mathsf{M}_{1},\mathsf{M}_{2})$ such that $(\mathsf{G}_{k}[\mathsf{M}_{1}^{\prime}],\mathsf{G}_{k}[\mathsf{M}_{2}^{\prime}])\overset{d}{=}(\mathsf{G}_{1},\mathsf{G}_{2})$.
Proof.
We apply the disintegration theorem [disintegration, Thm. A] to the continuous map $\pi_{k}\colon\mathcal{P}([0,1]^{2})\to\Delta^{k\times k}$ between compact metric spaces sending $\mu\mapsto(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}$. The disintegration theorem yields for each $\mathsf{M}\in\mathcal{M}$ a Borel collection of probability measures $(\lambda_{G}^{(\mathsf{M})})_{G\in\Delta^{k\times k}}$ for each $\mathsf{M}\in\mathcal{M}$ such that $\mathrm{supp}(\lambda_{G}^{(\mathsf{M})})\subseteq\pi_{k}^{-1}(G)$ and such that $\mathrm{Law}(\mathsf{M})=\mu_{k}^{(\mathsf{M})}\otimes\lambda_{G}^{(\mathsf{M})}$ in the sense that
$\mathbb{P}[\mathsf{M}\in B]=\int_{\Delta^{k\times k}}\int_{\pi_{k}^{-1}(G)}\mathbbm{1}_{B}(\mu)\,d\lambda_{G}^{(\mathsf{M})}(\mu)\,d\mu_{k}^{(\mathsf{M})}(s),$
for any Borel $B\subseteq\mathcal{P}([0,1]^{2})$. The measure $\lambda_{G}^{(\mathsf{M})}$ is the distribution of $\mathsf{M}$ conditioned on $\pi_{k}(\mathsf{M})=G$.
Let $\gamma=\mathrm{Law}((G_{1},G_{2}))$. We construct the coupling $\Gamma=\mathrm{Law}((\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime}))$ explicitly by setting
$\Gamma(B)=\int_{\Delta^{k\times k}\times\Delta^{k\times k}}\int_{\pi_{k}^{-1}(G_{2})}\int_{\pi_{k}^{-1}(G_{1})}\mathbbm{1}_{B}(\mu,\nu)\,d\lambda_{G_{1}}^{(\mathsf{M})}(\mu)\,d\lambda_{G_{2}}^{(\mathsf{N})}(\nu)\,d\gamma(G_{1},G_{2}).$
for each Borel $B\subseteq\mathcal{P}([0,1]^{2})\times\mathcal{P}([0,1]^{2})$. This distribution corresponds to first sampling $(\mathsf{G}_{1},\mathsf{G}_{2})\sim\gamma$ and then sampling $\mathsf{M}_{1}^{\prime}$ and $\mathsf{M}_{2}^{\prime}$ conditioned on $\mathsf{G}_{k}[\mathsf{M}_{1}^{\prime}]=\mathsf{G}_{1}$ and $\mathsf{G}_{k}[\mathsf{M}_{2}^{\prime}]=\mathsf{G}_{2}$. It is routine to check that this is a coupling of $(\mathsf{M}_{1},\mathsf{M}_{2})$ and that its pushforward under $(\pi_{k},\pi_{k})$ is the given coupling $\gamma$ of $(\mathsf{G}_{1},\mathsf{G}_{2})$, which is the claim we sought to prove. ∎
We can now prove Proposition 3.2.
Proof (Proposition 3.2)..
Define the intervals $I_{i}^{(n)}=[(i-1)/n,i/n)$ for $i\in[n]$ and let $(\mathsf{M}^{\prime},\mathsf{N}^{\prime})$ be a coupling of $(\mathsf{M},\mathsf{N})$ attaining $W_{\square}(\mathsf{M},\mathsf{N})$ in (7). Such a coupling exists since the infimum in (7) consists of minimizing a lower-semicontinuous function of the coupling over the compact space of all possible couplings. For each $k\in\mathbb{N}$, we have
$W_{1}(\mathsf{G}_{k}[\mathsf{M}],\mathsf{G}_{k}[\mathsf{N}])$ $\leq\sum_{i,j=1}^{k}\mathbb{E}|\mathsf{M}^{\prime}(I_{i}^{(k)}\times I_{j}^{(k)})-\mathsf{N}^{\prime}(I_{i}^{(k)}\times I_{j}^{(k)})|$
$\leq k^{2}W_{\square}(\mathsf{M},\mathsf{N}),$
giving the first claimed inequality.
For the second inequality, consider arbitrary intervals $S=[a_{1},a_{2}]$ and $T=[a_{3},a_{4}]$ for $a,b,c,d\in[0,1]$. For each $k\in\mathbb{N}$ and $j\in[4]$, let $i_{j}\in[k]$ be the unique index such that $a_{j}\in I_{i_{j}}^{(k)}$ (if $a_{j}=1$ set $i_{j}=k$). Note that $S\subseteq\bigcup_{i=i_{1}}^{i_{2}}I_{i}^{(k)}$ and $|\bigcup_{i=i_{1}}^{i_{2}}I_{i}^{(k)}\setminus S|\leq 2/k$, and similarly for $T$. Therefore, we have
$\lambda^{2}\left(\bigcup_{i=i_{1}}^{i_{2}}\bigcup_{j=i_{3}}^{i_{4}}I_{i}^{(k)}\times I_{j}^{(k)}\setminus S\times T\right)\leq\frac{4}{k},\quad\lambda_{D}\left(\bigcup_{i=i_{1}}^{i_{2}}\bigcup_{j=i_{3}}^{i_{4}}I_{i}^{(k)}\times I_{j}^{(k)}\setminus S\times T\right)\leq\frac{2\sqrt{2}}{k},$
hence by Lemma 3.12 we have
$\mathbb{E}|\mathsf{M}^{\prime}(S\times T)-\mathsf{N}^{\prime}(S\times T)|\leq\mathbb{E}\|\mathsf{G}_{k}[\mathsf{M}^{\prime}]-\mathsf{G}_{k}[\mathsf{N}^{\prime}]\|_{1}+\frac{4}{k}.$
Taking the supremum over $S,T$, we get
$W_{\square}(\mathsf{M}^{\prime},\mathsf{N}^{\prime})\leq\frac{4}{k}+\inf_{\begin{subarray}{c}\textrm{couplings}\\ (\mathsf{M}^{\prime},\mathsf{N}^{\prime})\end{subarray}}\mathbb{E}\|\mathsf{G}_{k}[\mathsf{M}^{\prime}]-\mathsf{G}_{k}[\mathsf{N}^{\prime}]\|_{1}=\frac{4}{k}+W_{1}(\mathsf{G}_{k}[\mathsf{M}],\mathsf{G}_{k}[\mathsf{N}]),$
where the last equality follows from Lemma 3.13. This is the second claimed inequality. ∎
Proof of Proposition 3.3
Next, we turn to proving Proposition 3.3, which states that $\mathcal{M}$ is weakly compact and that the metric $W_{\square}$ metrizes this weak topology. To prove that $\mathcal{M}$ is weakly compact, we rely on the following lemma and the characterization of $\mathcal{M}$ from [kallenberg2005probabilistic, §9] as the space of ergodic random exhangeable measures.
Lemma 3.14.
If $(\mu_{n})$ is a weakly convergent sequence of probability distributions on a Polish space $S$ with limit $\mu$, and if $B\subseteq S$ is a measurable subset satisfying $\mu_{n}(B)=1$ for all $n$, then $\mu(B)=1$.
Proof.
By Urysohn’s lemma, for any $\epsilon>0$ we can find a subset $B_{\epsilon}\supseteq B$ with $\mu(B_{\epsilon}\setminus B)\leq\epsilon$ and a continuous function $f\colon S\to[0,1]$ satisfying $f|_{B}=1$ and $f|_{S\setminus B_{\epsilon}}=0$. Since $f$ is continuous, we have $\mathbb{E}_{\mu}f=\lim_{n}\mathbb{E}_{\mu_{n}}f\geq\lim_{n}\mu_{n}(S)=1$. On the other hand, since $f\leq\mathbbm{1}_{B_{\epsilon}\setminus B}+\mathbbm{1}_{B}$ pointwise, we have $\mathbb{E}_{\mu}f\leq\mu(B)+\epsilon$. We conclude that $\mu(B)\geq 1$, and since $\mu$ is a probability measure, that $\mu(B)=1$, as desired. ∎
We are now to prove Proposition 3.3.
Proof (Proposition 3.3)..
We break down the proof into several steps.
Compactness of $\mathcal{M}$:
The space $\mathcal{P}([0,1]^{2})$ of probability measures on $0,1]^{2}$ is compact in the weak topology. Therefore, the space $\mathcal{P}(\mathcal{P}([0,1]^{2}))$ of random such measures is again compact in its own weak topology (see Section 1.3). Thus, it suffices to prove that $\mathcal{M}\subseteq\mathcal{P}(\mathcal{P}([0,1]^{2}))$ is closed in this weak topology.
By [kallenberg2005probabilistic, Prop. 9.1], the space of random exchangeable measures is precisely the collection of random measures on $0,1]^{2}$ invariant under the countable group $\widetilde{\mathsf{S}}_{\infty}$ consisting of permutations of the intervals $(I_{i}^{(n)})_{i\in[n]}$ for all $n\in\mathbb{N}$. By [kallenberg2005probabilistic, Lemma A1.2], we can characterize its extreme points by
$\mathcal{M}=\left\{\mathsf{M}\textrm{ random exchangeable}:\mathbb{P}[\mathsf{M}\in B]\in\{0,1\}\textrm{ whenever }B\subseteq\mathcal{P}([0,1]^{2})\textrm{ is }\widetilde{\mathsf{S}}_{\infty}\textrm{-invariant}\right\}.$
If $\mathsf{M}_{n}\to\mathsf{M}$ and $\mathrm{Law}(\mathsf{M}_{n})(B)\in\{0,1\}$ for all $n$, then passing to a subsequence we may assume $\mathrm{Law}(\mathsf{M}_{n})(B)=1$ for all $n$ or $\mathrm{Law}(\mathsf{M}_{n})(B)=0$ for all $n$. Applying Lemma 3.14 to either $B$ or its complement, respectively, yields $\mathrm{Law}(\mathsf{M})(B)\in\{0,1\}$. We thus conclude that $\mathcal{M}$ is weakly closed, and hence compact.
Equivalence of statements 2 and 3:
The equivalence of statements 2 and 3 follows from Proposition 3.2, since $\lim_{n}\mathsf{G}_{k}[\mathsf{M}_{n}]=\mathsf{G}_{k}[\mathsf{M}]$ weakly for a given $k\in\mathbb{N}$ if and only if $\lim_{n}W_{1}(\mathsf{G}_{k}[\mathsf{M}_{n}],\mathsf{G}_{k}[\mathsf{M}])=0$.
Equivalence of statements 3 and 4:
The equivalence of these two statements follows by the proof of Proposition 2.5. Specifically, the sequence $(\mathsf{G}_{k}[\mathsf{M}_{n}])$ converges weakly to $\mathsf{G}_{k}[\mathsf{M}]$ if and only if the moments of $(\mathsf{G}_{k}[\mathsf{M}_{n}])$ converge to those of $\mathsf{G}_{k}[\mathsf{M}]$, and these moments are precisely the quotient densities of grapheurs defined in (9).
Equivalence of statements 1 and 2:
Endow $\mathbb{R}^{2}$ with the $\ell_{\infty}$ norm, endow $\mathcal{P}([0,1]^{2})$ with the Wasserstein-1 metric with respect to this norm, and endow $\mathcal{M}$ with the Wasserstein-1 metric with respect to the Wasserstein-1 metric on $\mathcal{P}([0,1]^{2})$, which metrizes weak convergence of random measures [villani2008optimal, Thm. 6.9]. Explicitly, this metric is given by
$\mathrm{dist}_{W_{1}}(\mathsf{M}_{1},\mathsf{M}_{2})=\min_{\begin{subarray}{c}\textrm{random }(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})\\ \mathsf{M}_{i}^{\prime}\overset{d}{=}\mathsf{M}_{i}\textrm{ for }i=1,2\end{subarray}}\mathbb{E}_{(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})}\sup_{\begin{subarray}{c}f\colon[0,1]^{2}\to\mathbb{R}\\ f(0)=0\\ \textrm{ 1-Lipschitz in }\ell_{\infty}\end{subarray}}|\mathbb{E}_{\mathsf{M}_{1}^{\prime}}f-\mathbb{E}_{\mathsf{M}_{2}^{\prime}}f|,$
where the min over couplings is attained [villani2008optimal, Thm. 4.1]. It therefore suffices to prove that $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$ if and only if $\lim_{n}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})=0$. We proceed to prove this by standard approximations of continuous functions by step functions and vice-versa.
Suppose that $\lim_{n}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})=0$ and fix any $\epsilon>0$. Let $(\mathsf{M}_{n}^{\prime},\mathsf{M}^{\prime})$ be a coupling of $(\mathsf{M}_{n},\mathsf{M})$ attaining $\mathrm{dist}_{W_{1}}$. For any interval $S\subseteq[0,1]$, define $S_{\epsilon}=(S+[-\epsilon,\epsilon])\cap[0,1]$. Note that
$\lambda^{2}(S_{\epsilon}\times T_{\epsilon}\setminus S\times T)\leq 4\epsilon,\qquad\lambda_{D}(S_{\epsilon}\times T_{\epsilon}\setminus S\times T)\leq 2\sqrt{2}\epsilon.$
We can find an $\epsilon^{-1}$-Lipschitz $f_{\epsilon}\colon[0,1]^{2}\to[0,1]$ satisfying $f|_{S\times T}=1$ and $f|_{[0,1]^{2}\setminus S_{\epsilon}\times T_{\epsilon}}=0$. For example, set $f_{\epsilon}(x)=\frac{\mathrm{dist}(x,S_{\epsilon}\times T_{\epsilon})}{\mathrm{dist}(x,S\times T)+\mathrm{dist}(x,S_{\epsilon}\times T_{\epsilon})}$ where $\mathrm{dist}$ is with respect to the $\ell_{\infty}$ norm. For any $\mathsf{M}\in\mathcal{M}$, we have
$\mathbb{E}|\mathsf{M}(S\times T)-\mathbb{E}_{\mathsf{M}}f_{\epsilon}|\leq\mathbb{E}\mathsf{M}(S_{\epsilon}\times T_{\epsilon}\setminus S\times T)\leq 4\epsilon,$
by Lemma 3.12, hence
$|\mathbb{E}\mathsf{M}_{n}^{\prime}(S\times T)-\mathbb{E}\mathsf{M}^{\prime}(S\times T)|\leq 8\epsilon+|\mathbb{E}_{\mathsf{M}_{n}^{\prime}}f_{\epsilon}-\mathbb{E}_{\mathsf{M}^{\prime}}f_{\epsilon}|\leq 8\epsilon+\frac{1}{\epsilon}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M}).$
We conclude that $W_{\square}(\mathsf{M}_{n},\mathsf{M})\leq 8\epsilon+\epsilon^{-1}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})$, so taking $n\to\infty$ and then $\epsilon\to 0$ we conclude that $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$.
Conversely, suppose $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$. Once again, we note that the infimum over couplings in (7) is attained, since it consists of minimizing a lower-semicontinuous function of the coupling over the compact space of all possible couplings. For each $n$, let $(\mathsf{M}^{\prime},\mathsf{M}_{n}^{\prime})$ be a coupling attaining $W_{\square}(\mathsf{M}_{n},\mathsf{M})$. For each $N\in\mathbb{N}$, define the intervals $I_{i}^{(N)}=[(i-1)/N,i/N)$ for $i\in\{1,\ldots,N-1\}$ and $I_{N}^{(N)}=[1-1/N,1]$. For any 1-Lipschitz $f\colon[0,1]\to\mathbb{R}$ with $f(0)=0$, let
$f_{N}(x)=\sum_{i,j=1}^{N}f_{i,j}^{(N)}\mathbbm{1}_{I_{i}^{(N)}\times I_{j}^{(N)}}(x),$
where $f_{i,j}^{(N)}=f(\frac{i}{N}-\frac{1}{2N},\frac{j}{N}-\frac{1}{2N})$ is the value of $f$ on the center of the squares $I_{i}^{(N)}\times I_{j}^{(N)}$. Since $f$ is 1-Lipschitz in $\ell_{\infty}$ and the diameter of these squares is $1/2N$ in $\ell_{\infty}$, we have $|f(x)-f_{N}(x)|\leq\frac{1}{2N}$ for all $x\in[0,1]^{2}$. Furthermore, since $f(0)=0$ we have $|f_{i,j}^{(N)}|\leq 1$ for all $i,j$. Therefore, for any $\mathsf{M}\in\mathcal{M}$ we have
$|\mathbb{E}_{\mathsf{M}}f-\mathbb{E}_{\mathsf{M}}f_{N}|\leq\frac{1}{2N},$
and for our particular coupling $(\mathsf{M}^{\prime},\mathsf{M}_{n}^{\prime})$, we have
$|\mathbb{E}_{\mathsf{M}^{\prime}}f-\mathbb{E}_{\mathsf{M}_{n}^{\prime}}f|\leq\frac{1}{N}+|\mathbb{E}_{\mathsf{M}^{\prime}}f_{N}-\mathbb{E}_{\mathsf{M}_{n}^{\prime}}f_{N}|\leq\frac{1}{N}+\sum_{i,j=1}^{N}|\mathsf{M}^{\prime}(I_{i}^{(N)}\times I_{j}^{(N)})-\mathsf{M}_{n}^{\prime}(I_{i}^{(N)}\times I_{j}^{(N)})|.$
Since this inequality holds for any such function $f$, we conclude that
$\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})\leq\frac{1}{N}+\sum_{i,j=1}^{N}\mathbb{E}|\mathsf{M}^{\prime}(I_{i}^{(N)}\times I_{j}^{(N)})-\mathsf{M}_{n}^{\prime}(I_{i}^{(N)}\times I_{j}^{(N)})|\leq\frac{1}{N}+N^{2}W_{\square}(\mathsf{M}_{n},\mathsf{M}).$
Letting $n\to\infty$ and then $N\to\infty$, we conclude that $\lim_{n}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})=0$. ∎
4 Edge Sampling and Property Testing
In this section, we consider the problem of testing properties of large graphs by randomly sampling small summaries using our framework. For properties specified by graph parameters that are continuous with respect to quotient convergence, it would seem that random quotients are natural candidates for obtaining small summaries of large graphs. However, random quotients require access to the entire graph and are therefore impractical for large graphs. Moreover, we will see in Section 5 that random quotients do not concentrate very well, further obstructing property testing based on these summaries.
To remedy this situation, we present in this section an edge-based sampling procedure that yields a practical method for obtaining small summaries approximating large graphs of any size in the $W_{\square}$-metric with universal rates. This sampling procedure is based on an equivalent view of quotient convergence in terms of convergence of edge-exchangeable random graph models, the equivalence being furnished by the classical de Finetti’s theorem. We describe this perspective on quotient convergence in Section 4.1 and an edge-based analog of the Szemerédi regularity lemma in Section 4.2. We leverage these results to present a practical approach for property testing, which we formally define and study in Section 4.3. We conclude in Section 4.4 by presenting proofs from Section 4.2 that are omitted from notational clarity as well as the proof of Proposition 3.6 from Section 3.2.
4.1 Edge Sampling View of Quotient Convergence
We begin by describing a procedure to obtain finite random graphs from a grapheur by sampling edges with replacement. We then reinterpret quotient convergence in terms of convergence of the resulting random edge-exchangeable graphs.
Given a grapheur $\mathsf{M}$, we can obtain a sequence of weighted random graphs $(\mathsf{G}^{(n)}[\mathsf{M}])_{n}$ where $\mathsf{G}^{(n)}[\mathsf{M}]$ has $n$ edges as follows. We sample a measure $\mu\sim\mathsf{M}$ on $0,1]^{2}$, followed by sampling $n$ edges $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\overset{iid}{\sim}\mu$. Next we construct the weighted graph $\mathsf{G}^{(n)}[\mathsf{M}]$ on $N=|\{x_{i}\}\cup\{y_{i}\}|$ vertices by arbitrarily enumerating $\{t_{1},\ldots,t_{N}\}=\{x_{i}\}\cup\{y_{i}\}$ and setting the edge weights to:
$(\mathsf{G}^{(n)}[\mathsf{M}])_{i,j}=|\{k\in[n]:(x_{k},y_{k})=(t_{i},t_{j})\}|/n$
for $1\leq i,j\leq N$. Note that $|V(\mathsf{G}^{(n)}[\mathsf{M}])|\leq 2n$.
If the grapheur $\mathsf{M}$ is the limit of undirected graphs, we can further replace $\mathsf{G}^{(n)}[\mathsf{M}]$ by its symmetrization $\frac{\mathsf{G}^{(n)}[\mathsf{M}]+(\mathsf{G}^{(n)}[\mathsf{M}])^{\top}}{2}$ to obtain an undirected random graph; see Remark 4.6 below. When $\mathsf{M}=\mathsf{M}_{G}$ is the grapheur corresponding to a finite graph $G$ as in (4), the above sampling procedure amounts to sampling with replacement the edges of $G$ proportionally to their weight, and randomly labeling the resulting graph.
For a general grapheur $\mathsf{M}$ described by parameters $(E,\sigma,\varsigma,\theta,\vartheta)$, edge sampling from $\mathsf{M}$ as above corresponds to first sampling random locations $(T_{i})_{i}$ and then sampling edges as follows.
  • With probability $E_{i,j}$, sample the edge $(T_{i},T_{j})$.
  • With probability $\sigma_{i}$, we sample a new vertex $T^{\prime}\sim\mathrm{Unif}([0,1])$ and sample the edge $(T_{i},T^{\prime})$. With probability $\varsigma_{i}$, we analogously sample $(T^{\prime},T_{i})$.
  • With probability $\theta$ we sample two new vertices $T_{1}^{\prime},T_{2}^{\prime}\sim\mathrm{Unif}([0,1])$ and add the edge $(T_{1}^{\prime},T_{2}^{\prime})$. Similarly, with probability $\vartheta$ we add the edge $(T^{\prime\prime},T^{\prime\prime})$ for a new location $T^{\prime\prime}\sim\mathrm{Unif}([0,1])$.
After repeating the above procedure $n$ times, we get a random exchangeable multigraph containing $n$ edges (possibly with repetition), and normalize the resulting graph to have total edge weight one.
We now observe that convergence of edge-sampled graphs is equivalent to quotient convergence.
Proposition 4.1.
We have $\lim_{n}\mathsf{M}_{n}=\mathsf{M}$ in $\mathcal{M}$ if and only if $(\mathsf{G}^{(k)}[\mathsf{M}_{n}])_{n}\subseteq\Delta^{2k\times 2k}$ converges weakly to $\mathsf{G}^{(k)}[\mathsf{M}]$ for all $k\in\mathbb{N}$.
Proof.
The grapheurs $(\mathsf{M}_{n})$ converge weakly if and only if the random exchangeable arrays
$\begin{bmatrix}x_{1}^{(n)}&\cdots&x_{k}^{(n)}\\ y_{1}^{(n)}&\cdots&y_{k}^{(n)}\end{bmatrix}\in[0,1]^{2\times k},$ (15)
obtained by first sampling $\mu_{n}\sim\mathsf{M}_{n}$ and then sampling $(x_{i},y_{i})\overset{iid}{\sim}\mu_{n}$ converge weakly by [kallenberg1973characterization, Thm. 1.1], since the distributions of (15) are precisely the finite-dimensional distributions of $\mathsf{M}_{n}$. The exchangeable array (15) is precisely the list of edges of the multigraph $k\mathsf{G}^{(k)}[\mathsf{M}_{n}]$ listed in arbitrary order after assigning its vertices iid labels in $\mathrm{Unif}([0,1])$. Formally, we have $\mathsf{M}_{\mathsf{G}^{(k)}[\mathsf{M}_{n}]}\overset{d}{=}\frac{1}{k}\sum_{i=1}^{k}\delta_{(x_{i}^{(n)},y_{i}^{(n)})}$ for all $n,k\in\mathbb{N}$. Thus, the sequence (15) converges weakly for each $k\in\mathbb{N}$ precisely when the sequence of random graphs $(\mathsf{G}^{(k)}[\mathsf{M}_{n}])_{n}$ converges weakly. ∎
In particular, a sequence of graphs $(G_{n}\in\Delta^{n\times n})$ quotient-converges precisely when the sequence of random graphs $(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}])$—obtained by independently sampling $k$ edges from each $G_{n}$ proportionally to their edge weights and randomly labeling the vertices—converges weakly for each $k\in\mathbb{N}$.
Remark 4.2 (Edge-exchangeable random graphs).
The random graphs we obtain by sampling edges from a grapheur appeared previously in the literature as edge-exchangeable random multigraph models [janson2018edge]. As explained in [janson2018edge, Rmk. 4.4], every edge-exchangeable random multigraph model is obtained from a random exchangeable measure on $0,1]^{2}$ as above after labelling the vertices by iid labels from $\mathrm{Unif}([0,1])$. Nevertheless, to our knowledge graph limits based on these edge-exchangeable random graphs as in Proposition 4.1 have not been previously studied in the literature, see also Section 1.2.
As noted in [janson2018edge], if we repeat the above edge sampling process infinitely-many times, each edge $(T_{i},T_{j})$ corresponding to a positive edge weight $E_{i,j}>0$ will be repeated infinitely-many times. In contrast, all the other edges sampled above will only appear once. In particular, the vertices $T^{\prime},T_{1}^{\prime},T_{2}^{\prime}$ sampled in this process will only appear in one edge. Such vertices are called “blips” in [borgs2019sampling, janson2018edge], and the isolated edges and self-loops are called “dust”.
4.2 Edge-Based Szemerédi Regularity
We now provide bounds on the rate of convergence for $W_{\square}(\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M})$ to zero as $n$ grows large, both in expectation and with high probability. In bounding the quantity $W_{\square}(\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M})$, it is important to distinguish between the randomness in the grapheur $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}$ coming from the random atoms $T_{i}$ and randomness coming from $\mathsf{G}^{(n)}[\mathsf{M}]$. In our next result, we fix a realization of $\mathsf{G}^{(n)}[\mathsf{M}]$, compute the distance $W_{\square}(\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M})$ depending on this realization (and hence random), and seek to bound both its expectation and its tails with respect to the randomness in $\mathsf{G}^{(n)}[\mathsf{M}]$.
In fact, our results provide convergence rates for the larger metric $\overline{W_{\square}}(\mathsf{M},\mathsf{N})$ defined by
$W_{\square}(\mathsf{M},\mathsf{N})\leq\overline{W_{\square}}(\mathsf{M},\mathsf{N})=\inf_{\begin{subarray}{c}\textrm{random }(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})\\ \mathsf{M}_{i}^{\prime}\overset{d}{=}\mathsf{M}_{i}\textrm{ for }i=1,2\end{subarray}}\ \mathbb{E}\left[\sup_{\begin{subarray}{c}S,T\subseteq[0,1]\\ \textrm{intervals}\end{subarray}}|\mathsf{M}_{1}^{\prime}(S\times T)-\mathsf{M}_{2}^{\prime}(S\times T)|\right],$
where the expectation is taken with respect to the coupling $(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})$ of $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$. Note that $\overline{W_{\square}}$ differs from $W_{\square}$ in (7) by moving the supremum over axis-aligned rectangles into the expectation.
Theorem 4.3.
For any grapheur $\mathsf{M}\in\mathcal{M}$ and each $n\in\mathbb{N}$, let $\mathsf{G}^{(n)}[\mathsf{M}]$ be a random graph obtained by sampling $n$ edges from $\mathsf{M}$. Then
$\mathbb{E}_{\mathsf{G}^{(n)}[\mathsf{M}]}\left[W_{\square}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})\right]\leq\mathbb{E}_{\mathsf{G}^{(n)}[\mathsf{M}]}\left[\overline{W_{\square}}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})\right]\leq\frac{174}{\sqrt{n}}.$
Moreover, $W_{\square}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})\leq\overline{W_{\square}}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})\leq\frac{174+\epsilon}{\sqrt{n}}$ with probability at least $1-\exp(-2\epsilon^{2})$ over the random graph $\mathsf{G}^{(n)}[\mathsf{M}]$.
The proof of Theorem 4.3 relies on two key lemmas. The first involves the construction of a particular coupling of $\mathsf{M}$, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}$, and $\mathsf{G}^{(n)}[\mathsf{M}]$.
Lemma 4.4.
There is a coupling $(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}},\mathsf{G}^{(n)})$ of $\mathsf{M}$, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}$, and $\mathsf{G}^{(n)}[\mathsf{M}]$ satisfying the following two properties. First, in this coupling $\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$ is obtained by sampling $\mu\sim\mathsf{M}^{\prime}$, then sampling $n$ iid points from $\mu$, and finally forming the associated empirical measure. Second, conditioned on $\mathsf{G}^{(n)}=G$, the pair $(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}})$ is a coupling of $\mathsf{M}$ and $\mathsf{M}_{G}$.
The second step in the proof of Theorem 4.3 involves using Vapnik-Chervonenkis theory to uniformly bound the discrepancy between a measure and its empirical sample on axis-aligned rectangles.
Lemma 4.5.
For any measure $\mu\in\mathcal{P}([0,1]^{2})$, let $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ be the random empirical measure obtained by sampling $X_{1},\ldots,X_{n}\overset{iid}{\sim}\mu$. Then
$\mathbb{E}_{\mu_{n}}\sup_{\begin{subarray}{c}S,T\subseteq[0,1]\\ \textrm{intervals}\end{subarray}}|\mu(S\times T)-\mu_{n}(S\times T)|\leq\frac{174}{\sqrt{n}}.$
The proofs of these two lemmas are straightforward but technical, and we defer them to Section 4.4 for clarity of exposition. Using the lemmas, we can now prove Theorem 4.3.
Proof (Theorem 4.3)..
Let $\mathcal{AAR}=\{[a,b]\times[c,d]:a,b,c,d\in[0,1]\}$ be the class of axis-aligned rectangles contained in $0,1]^{2}$, and let $(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}},\mathsf{G}^{(n)})$ be the coupling from Lemma 4.4. Define
$f(\mathsf{G}^{(n)}[\mathsf{M}])=f((x_{1},y_{1}),\ldots,(x_{n},y_{n}))=\overline{W_{\square}}(\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M}),$
which is a function of the edges $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ sampled iid from $\mu\sim\mathsf{M}$ to create $\mathsf{G}^{(n)}[\mathsf{M}]$. First note that
$\mathbb{E}_{\mathsf{G}^{(n)}[\mathsf{M}]}f(\mathsf{G}^{(n)}[\mathsf{M}])=\mathbb{E}_{\mathsf{G}^{(n)}}f(\mathsf{G}^{(n)})\leq\mathbb{E}_{(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}},\mathsf{G}^{(n)})}\sup_{R\in\mathcal{AAR}}|\mathsf{M}^{\prime}(R)-\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}(R)|\leq\frac{174}{\sqrt{n}},$
where the equality follows from $\mathsf{G}^{(n)}[\mathsf{M}]\overset{d}{=}\mathsf{G}^{(n)}$, the first inequality follows by considering the coupling $(\mathsf{M}^{\prime},\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}})|\mathsf{G}^{(n)}$ of $\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}$, and the last inequality follows by Lemma 4.5 and the fact that $\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$ is an empirical measure obtained from $n$ iid samples from $\mathsf{M}^{\prime}$.
Second, note that $f$ is $1/n$-Lipschitz in each $(x_{i},y_{i})$. Indeed, if $\mathsf{G}^{(n,i)}[\mathsf{M}]$ differs from $\mathsf{G}^{(n)}[\mathsf{M}]$ only in the $i$th location $(x_{i},y_{i})$, and we glue the two couplings $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})$ attaining the corresponding $\overline{W_{\square}}$ distances as in [villani2008optimal, §1], then
$|f(\mathsf{G}^{(n,i)}[\mathsf{M}])-f(\mathsf{G}^{(n)}[\mathsf{M}])|\leq\mathbb{E}_{(\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})}\sup_{R\in\mathcal{AAR}}|\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}(R)-\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}(R)|\leq\frac{1}{n},$
since the only difference between $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}$ and $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}$ is a weight of $1/n$ placed at one of their atoms. Since $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are sampled iid from a fixed measure $\mu$ (after it was sampled once from $\mathsf{M}$), the bounded-difference concentration inequality [wainwright2019high, Cor. 2.21] implies that $f(\mathsf{G}^{(n)}[\mathsf{M}])\leq\mathbb{E}f(\mathsf{G}^{(n)}[\mathsf{M}])+\epsilon/\sqrt{n}$ with probability at least $1-\exp(-2\epsilon^{2})$, proving the theorem. ∎
Using Theorem 4.3, we can now prove Theorem 1.7 from Section 1.
Proof (Theorem 1.7)..
The first claim follows from Theorem 4.3. The second claim follows from the fact that $\mathbb{E}W_{\square}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})\leq\frac{174}{\sqrt{n}}$, which implies that some realization $G^{(n)}$ of $\mathsf{G}^{(n)}[\mathsf{M}]$, satisfies $W_{\square}(\mathsf{M},\mathsf{M}_{G^{(n)}})\leq\frac{174}{\sqrt{n}}$. Setting $n=k(\epsilon)=\lceil(174/\epsilon)^{2}\rceil$, we obtain the second claim. ∎
We conclude this section by remarking that the above discussion applies to undirected random graphs sampled from symmetric grapheurs.
Remark 4.6 (Undirected graphs).
If $\mathsf{M}$ is the limit of undirected graphs, then by Corollary 3.7 it is symmetric, i.e., we have $\mathsf{M}(S\times T)=\mathsf{M}(T\times S)$ for any measurable $S,T\subseteq[0,1]$. In that case, we can sample undirected graphs from $\mathsf{M}$ by first sampling $\mathsf{G}^{(n)}[\mathsf{M}]$ as above and symmetrizing it to get $\mathsf{G}_{\mathrm{sym}}^{(n)}[\mathsf{M}]=\frac{\mathsf{G}^{(n)}[\mathsf{M}]+(\mathsf{G}^{(n)}[\mathsf{M}])^{\top}}{2}$. For any $S,T\subseteq[0,1]$, we have
$|\mathsf{M}(S\times T)-\mathsf{M}_{\mathsf{G}_{\mathrm{sym}}^{(n)}[\mathsf{M}]}(S\times T)|$ $=\frac{1}{2}|\mathsf{M}(S\times T)+\mathsf{M}(T\times S)-\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}(S\times T)-\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}(T\times S)|$
$\leq\max\{|\mathsf{M}(S\times T)-\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}(S\times T)|,|\mathsf{M}(T\times S)-\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}(T\times S)|\},$
hence $W_{\square}(\mathsf{M},\mathsf{M}_{\mathsf{G}_{\mathrm{sym}}^{(n)}[\mathsf{M}]})\leq W_{\square}(\mathsf{M},\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and similarly for $\overline{W_{\square}}$. Thus, all the results of this section apply to the undirected random graph $\mathsf{G}^{(n)}_{\mathrm{sym}}[\mathsf{M}]$.
4.3 Property Testing
We apply the preceding edge-based sampling procedure to test properties of large graphs based on small summaries. We begin by formally defining the class of graph parameters we can test using our framework. Here a graph parameter is a function $f\colon\bigsqcup_{n\in\mathbb{N}}\Delta^{n\times n}\to\mathbb{R}$ on graphs of all sizes that is permutation-invariant, in the sense that $f(\pi_{n}G_{n})=f(G_{n})$ for any permutation $\pi_{n}\in\mathsf{S}_{n}$ of the vertex labels.
Definition 4.7 (Quotient-testable graph parameters).
A quotient-testable graph parameter is a graph parameter $f\colon\bigsqcup_{n\in\mathbb{N}}\Delta^{n\times n}\to\mathbb{R}$ satisfying the following properties:
  1. fnum@enumiitem(Isolate-indifference)
    $f(G_{N})=f(G_{n})$ whenever $G_{N}\in\Delta^{N\times N}$ differs from $G_{n}\in\Delta^{n\times n}$ by isolated vertices;
  2. fnum@enumiitem(Continuity)
    The sequence $(f(G_{n}))$ converges whenever $(G_{n}\in\Delta^{n\times n})$ is quotient-convergent.
We proceed to give several equivalent characterizations of quotient-testable parameters.
Proposition 4.8 (Testability as continuity).
The following are equivalent for a graph parameter $f$.
  1. 1.
    The parameter $f$ is quotient-testable.
  2. 2.
    There is a function $\bar{f}\colon\mathcal{M}\to\mathbb{R}$ satisfying $\bar{f}(\mathsf{M}_{G})=f(G)$ for all $G\in\bigsqcup_{n}\Delta^{n\times n}$. Such $\bar{f}$ is then unique and continuous.
  3. 3.
    For any $\epsilon>0$ there is a $\delta=\delta(\epsilon)>0$ such that if $W_{\square}(\mathsf{M}_{G_{n}},\mathsf{M}_{G_{m}})\leq\delta$ then $|f(G_{n})-f(G_{m})|\leq\epsilon$ for any two graphs $G_{n},G_{m}$ of any two sizes.
Proof.
We prove these statements imply each other in order, with the last statement implying the first.
Suppose $f$ is quotient-testable. Define $\mathsf{M}_{<\infty}=\{\mathsf{M}_{G}:G\in\Delta^{n\times n}\textrm{ for some }n\}\to\mathbb{R}$ to be the grapheurs in $\mathsf{M}$ arising from finite graphs and define $\bar{f}\colon\mathsf{M}_{<\infty}\to\mathbb{R}$ by $\bar{f}(\mathsf{M}_{G})=f(G)$. This is well-defined because $f$ is permutation-invariant and isolate-indifferent, as $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ if and only if $G$ and $G^{\prime}$ differ by isolated vertices and node relabeling [kallenberg1990exchangeable, Prop. 3]. Further, $\bar{f}$ is continuous by continuity of $f$, hence it extends uniquely to a continuous function on the closure $\mathcal{M}$ of $\mathcal{M}_{<\infty}$. This proves statement 2.
Suppose $f$ extends to a continuous $\bar{f}\colon\mathcal{M}\to\mathbb{R}$. Since $\mathcal{M}$ is compact and this compact topology is metrized by the $W_{\square}$ metric by Proposition 3.3, the function $\bar{f}$ is uniformly continuous in the $W_{\square}$ metric. This proves statement 3.
Finally, assume statement 3 holds. Since $W_{\square}(\mathsf{M}_{G_{N}},\mathsf{M}_{G_{n}})=0$ if $G_{n}$ and $G_{N}$ differ by isolated vertices, we conclude that $f$ is isolate-indifferent. Since quotient-convergent sequences $(G_{n})_{n}$ are Cauchy in the $W_{\square}$-metric by Theorem 1.5, we conclude that $(f(G_{n}))_{n}\subseteq\mathbb{R}$ is Cauchy and hence converges. ∎
As a first set of examples, quotient densities and homomorphism numbers are quotient-testable. More generally, we show next that a large class of polynomial graph parameters are testable, because all such polynomial parameters are functions of fixed-dimensional random quotients of their input graphs.
Proposition 4.9 (Polynomial parameters).
Suppose $f$ is an isolate-indifferent graph parameter such that $f|_{\Delta^{n\times n}}$ is given by a polynomial function for each $n$, and that all these polynomials have degree $d\in\mathbb{N}$. Then $f$ is quotient-testable.
Further, a graph parameter $f$ is quotient-testable if and only if for any $\epsilon>0$ there exists a polynomial quotient-testable parameter $p$ satisfying $|f(G)-p(G)|\leq\epsilon$ for all graphs $G$ of all sizes.
Proof.
For the first claim, it was shown in [levin2025deFin, Thm. 5.2] that there exists a polynomial $q_{k}$ over $\mathbb{R}^{k\times k}$ satisfying $f(G_{n})=\mathbb{E}q_{k}(\rho(F_{k,n})G_{n})$ for all $G_{n}\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$. Since $q_{k}$ is $L$-Lipschitz on $\Delta^{k\times k}$ for some $L>0$, we have by Proposition 3.2 that
$|f(G_{n})-f(G_{m})|$ $=|\mathbb{E}q_{k}(\rho(F_{k,n}G_{n}))-\mathbb{E}q_{k}(\rho(F_{k,m})G_{m})|\leq LW_{1}(\rho(F_{k,n})G_{n},\rho(F_{k,m})G_{m})$
$\leq Lk^{2}W_{\square}(\mathsf{M}_{G_{n}},\mathsf{M}_{G_{m}}),$
for any two graphs $G_{n},G_{m}$ of any two sizes. Thus, the function $f$ is testable by Proposition 4.8(3), as can be seen by setting $\delta(\epsilon)=\frac{\epsilon}{Lk^{2}}$.
For the second claim, if for any $\epsilon>0$ there is a testable parameter $p$ satisfying $|f(G)-p(G)|\leq\epsilon$ for all $G$, then $f$ is a uniform limit of $W_{\square}$-continuous functions and hence is $W_{\square}$-continuous itself. By Proposition 4.8, the parameter $f$ is therefore testable. Conversely, if $f$ is testable then it extends to a continuous $\bar{f}\colon\mathcal{M}\to\mathbb{R}$ by Proposition 4.8 again. Note that the collection of polynomial testable parameters is an algebra, contains constant functions, and separates points by Corollary 3.4. Since $\mathcal{M}$ is compact, Stone–Weierestrass implies that $\bar{f}$ is the uniform limit of polynomial testable parameters. ∎
We combine the insights of Sections 4.1-4.2 along with the above characterization of quotient-testable parameters to show that these can indeed be tested by edge sampling.
Theorem 4.10.
Consider a graph parameter $f\colon\bigsqcup_{n}\Delta^{n\times n}\to\mathbb{R}$ whose restrictions $f|_{\Delta^{n\times n}}$ are continuous (in the usual topology on $\Delta^{n\times n}$) for each $n$. Then $f$ is quotient-testable if and only if for every $\epsilon>0$, there exists $k\in\mathbb{N}$ such that for any finite graph $G$ of any size we have $|f(G)-f(\mathsf{G}^{(k)}[\mathsf{M}_{G}])|\leq\epsilon$ with probability at least $1-\epsilon$. In this case, we can take $k(\epsilon)=\left\lceil\left(\frac{174+\sqrt{\log(\epsilon^{-1/2})}}{\delta(\epsilon)}\right)^{2}\right\rceil$ where $\delta(\epsilon)>0$ is as in Proposition 4.8(3).
Proof.
Suppose $f$ is quotient-testable. For any $\epsilon>0$, let $\delta(\epsilon)$ be as in Proposition 4.8(3) and set $k=k(\epsilon)$ as in the theorem statement. Then, noting that $\epsilon=e^{-2\widetilde{\epsilon}^{2}}$ where $\widetilde{\epsilon}=\sqrt{\log(\epsilon^{-1/2})}$, with probability at least $1-\epsilon$ we have $W_{\square}(\mathsf{M}_{G},\mathsf{M}_{\mathsf{G}^{(k)}})\leq\frac{174+\widetilde{\epsilon}}{\sqrt{k}}\leq\delta(\epsilon)$ and hence $|f(G)-f(\mathsf{G}^{(k)}[\mathsf{M}_{G}])|\leq\epsilon$ by Theorem 4.3.
We turn to proving the converse. If $G_{1}$ and $G_{2}$ differ by isolated vertices, then $\mathsf{M}_{G_{1}}=\mathsf{M}_{G_{2}}$ and hence $\mathsf{G}^{(k)}[\mathsf{M}_{G_{1}}]\overset{d}{=}\mathsf{G}^{(k)}[\mathsf{M}_{G_{2}}]$. Therefore, for any $\epsilon>0$ let $k\in\mathbb{N}$ be as in the theorem statement and let $\mathsf{G}^{(k)}$ be a sample from this common distribution. Then with probability $1-2\epsilon$ we have $|f(G_{1})-f(\mathsf{G}^{(k)})|\leq\epsilon$ and $|f(\mathsf{G}^{(k)})-f(G_{2})|\leq\epsilon$, so in particular there is a realization $G^{(k)}$ of $\mathsf{G}^{(k)}$ satisfying both inequalities. We then have $|f(G_{1})-f(G_{2})|\leq|f(G_{1})-f(G^{(k)})|+|f(G^{(k)})-f(G_{2})|\leq 2\epsilon$. Since $\epsilon>0$ was arbitrary, we conclude that $f$ is isolate-indifferent. A similar argument shows that $|f|$ is bounded. Indeed, since $|f|$ restricted to the compact set $\Delta^{2k\times 2k}$ is continuous, it is bounded by some $B_{2k}>0$. For any other graph $G$ of any size, some realization $G^{(k)}$ of $\mathsf{G}^{(k)}[\mathsf{M}_{G}]$ satisfies $|f(G)-f(G^{(k)})|\leq\epsilon$, and hence $|f(G)|\leq B_{2k}+\epsilon$ for all graphs $G$. Denoting $B=\sup_{G}|f(G)|$, we then have $|f(G)-\mathbb{E}f(\mathsf{G}^{(k)}[\mathsf{M}_{G}])|\leq\mathbb{E}|f(G)-f(\mathsf{G}^{(k)}[\mathsf{M}_{G}])|\leq\epsilon(1-\epsilon)+2B\epsilon$. Finally, if $(G_{n})$ is quotient-convergent, then $(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}])_{n}$ is weakly-convergent, and by continuity of $f|_{\Delta^{2k\times 2k}}$, the sequence $(\mathbb{E}f(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}]))_{n}$ is convergent. Therefore, we have
$|f(G_{n})-f(G_{m})|\leq 2\epsilon(1-\epsilon+2B)+|\mathbb{E}f(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}])-\mathbb{E}f(\mathsf{G}^{(k)}[\mathsf{M}_{G_{m}}])|\leq 3\epsilon(1-\epsilon+2B)$
for all large $n,m$, showing that $(f(G_{n}))$ is convergent. ∎
We remark that the graphon analog of the characterization of quotient-testable parameters in Theorem 4.10 was used as the definition of testable parameters in [convergent_seqs1, Def. 2.11].
The upshot of this result is that a continuous graph parameter being quotient-testable is equivalent to the parameter being testable using edge sampling. In particular, Theorem 4.10 gives a simple recipe for estimating the value of a testable parameter $f(G)$ on an arbitrarily-large graph $G$: sample $k$ edges iid from $G$ proportionally to their edge weights, and compute the parameter $f(G^{(k)})$ of the resulting small graph $G^{(k)}$. The number of edges needed is given explicitly as $k=\left\lceil\left(\frac{174+\sqrt{\log(\epsilon^{-1/2})}}{\delta(\epsilon)}\right)^{2}\right\rceil$ in terms of the desired accuracy $\epsilon$ and the modulus of continuity $\delta(\epsilon)$ of $f$ in the $W_{\square}$-metric. When $f$ is $L$-Lipschitz continuous in the $W_{\square}$-metric, we can set $\delta(\epsilon)=\epsilon/L$.
Many graph parameters of interest are indeed Lipschitz-continuous, including all polynomial parameters as the proof of Proposition 4.9 demonstrates. Thus, the key quantity that is needed to bound the required number of edges to test the values of such parameters is their Lipschitz constants. We can explicitly bound the Lipschitz constants of quotient densities. Specifically, for a multigraph $H\in\mathbb{N}^{k\times k}$ we have
$|t_{Q}(H;\mathsf{M}_{1})-t_{Q}(H;\mathsf{M}_{2})|=|\mathbb{E}\mathsf{G}_{k}[\mathsf{M}_{1}]^{H}-\mathbb{E}\mathsf{G}_{k}[\mathsf{M}_{2}]^{H}|\leq k^{2}\|H\|_{\infty}W_{\square}(\mathsf{M}_{1},\mathsf{M}_{2}),$
since the polynomial $G\mapsto G^{H}$ is $\|H\|_{\infty}$-Lipschitz with respect to the $\ell_{1}$ norm on $\Delta^{k\times k}$. By using (10) and (12), we can similarly derive Lipschitz constants for homomorphism numbers. In the next example, we derive the Lipschitz constant of a particular graph parameter that can be used to test the presence of hubs.
Example 4.11 (Testing for hubs).
Consider the graph parameter
$f(G)=\mathrm{hom}(S_{2}^{(\mathrm{out})};G)+\mathrm{hom}(S_{2}^{(\mathrm{in})};G)=\|G\mathbbm{1}\|_{2}^{2}+\|G^{\top}\mathbbm{1}\|_{2}^{2},$
where $S_{2}^{(\mathrm{out})},S_{2}^{(\mathrm{in})}\in\mathbb{N}^{3\times 3}$ are outward- and inward-pointing stars with two leaves, respectively. Define the polynomial $q_{2}$ on $2\times 2$ matrices by
$q_{2}(G)=(G_{1,1}+G_{1,2}-G_{2,1}-G_{2,2})^{2}+(G_{1,1}+G_{2,1}-G_{1,2}-G_{2,2})^{2},$
and observe that $f(G_{n})=\mathbb{E}q_{2}(\rho(F_{2,n})G_{n})$ for any $G_{n}\in\Delta^{n\times n}$ and any $n\in\mathbb{N}$. Indeed, more generally for any grapheur $\mathsf{M}$ with random locations $(T_{i})$, let $R_{i}\overset{iid}{\sim}\mathrm{Unif}(\{\pm 1\})$ be iid Rademacher random variables and note that
$\mathbb{E}q_{2}(\mathsf{G}_{2}[\mathsf{M}])$ $=\mathbb{E}\left[\sum_{i\in\mathbb{N}}(E\mathbbm{1}+\sigma)_{i}\left(\mathbbm{1}_{T_{i}<1/2}-\mathbbm{1}_{T_{i}>1/2}\right)\right]^{2}+\mathbb{E}\left[\sum_{i\in\mathbb{N}}(E^{\top}\mathbbm{1}+\varsigma)_{i}\left(\mathbbm{1}_{T_{i}<1/2}-\mathbbm{1}_{T_{i}>1/2}\right)\right]^{2}$
$=\mathbb{E}\left[\sum_{i\in\mathbb{N}}(E\mathbbm{1}+\sigma)_{i}R_{i}\right]^{2}+\mathbb{E}\left[\sum_{i\in\mathbb{N}}(E^{\top}\mathbbm{1}+\varsigma)_{i}R_{i}\right]^{2}=\|E\mathbbm{1}+\sigma\|_{2}^{2}+\|E^{\top}\mathbbm{1}+\varsigma\|_{2}^{2},$
where $\mathbbm{1}_{T_{i}<1/2}=1$ if $T_{i}<1/2$ and zero otherwise and similarly for $\mathbbm{1}_{T_{i}>1/2}$, and we note that $(\mathbbm{1}_{T_{i}<1/2}-\mathbbm{1}_{T_{i}>1/2})_{i\in\mathbb{N}}\overset{d}{=}(R_{i})_{i\in\mathbb{N}}$. We conclude that $f$ can be extended to all grapheurs $\mathsf{M}$ by $\bar{f}(\mathsf{M})=\mathbb{E}q_{2}(\mathsf{G}_{2}[\mathsf{M}])$. This extension is Lipschitz-continuous. Explicitly, since $|\partial_{G_{i,j}}q_{2}(G)|=2|G_{1,1}-G_{2,2}|$ if $i=j$ and $2|G_{i,j}-G_{j,i}|$ if $i\neq j$, we conclude that $q_{2}$ is 2-Lipschitz on $\Delta^{2\times 2}$ with respect to $\|\cdot\|_{1}$, and hence that
$|\bar{f}(\mathsf{M}_{1})-\bar{f}(\mathsf{M}_{2})|=|\mathbb{E}q_{2}(\mathsf{G}_{2}[\mathsf{M}_{1}])-\mathbb{E}q_{2}(\mathsf{G}_{2}[\mathsf{M}_{2}])|\leq 2W_{1}(\mathsf{G}_{2}[\mathsf{M}_{1}],\mathsf{G}_{2}[\mathsf{M}_{2}])\leq 8W_{\square}(\mathsf{M}_{1},\mathsf{M}_{2}),$
for any two grapheurs $\mathsf{M}_{1},\mathsf{M}_{2}\in\mathcal{M}$ by Proposition 3.2. Finally, observe that $\bar{f}(\mathsf{M})=0$ if and only if $E=0$ and $\sigma=\varsigma=0$, hence if and only if $\mathsf{M}$ has no hubs. In particular, $W_{\square}(\mathsf{M}_{G},\mathsf{M})\geq f(G)/8$ for any grapheur $\mathsf{M}$ with no hubs.
We can now use the above parameter to test for the existence of significant hubs in a graph. Specifically, suppose we are given an arbitrarily-large graph $G$, and we sample $k=\left\lceil 64\left(\frac{174+\sqrt{\log(\epsilon^{-1/2})}}{\epsilon}\right)^{2}\right\rceil$ edges from $G$ to obtain a graph $G^{(k)}$. Then with probability $1-\epsilon$, we have $W_{\square}(\mathsf{M}_{G},\mathsf{M})\geq\frac{f(G^{(k)})-\epsilon}{8}$ for any grapheur $\mathsf{M}$ with no hubs. In this sense, large values of $f(G^{(k)})$ indicate the likely presence of significant hubs in $G$.
We observe that testing for hubs also has an appealing interpretation in terms of quotients. A grapheur $\mathsf{M}$ has no hubs if and only if there is a $\theta\in[0,1]$ satisfying $\mathsf{G}_{m}[\mathsf{M}]=\theta\frac{1}{m^{2}}\mathbbm{1}\mathbbm{1}^{\top}+(1-\theta)\frac{1}{m}I_{m}$ (a deterministic matrix) for all $m\in\mathbb{N}$. Thus, testing for hubs also amounts to testing whether the random quotients of a graph $G$ deviate from a uniform distribution of edge weights, i.e., quotients of the above form. See Figures 1-2 for example. The parameter of Example 4.11 yields a lower bound on this deviation, as by Proposition 3.2 we have
$\min_{\theta\in[0,1]}\mathbb{E}\left\|\theta\tfrac{1}{m^{2}}\mathbbm{1}\mathbbm{1}^{\top}+(1-\theta)\tfrac{1}{m}I_{m}-\mathsf{G}_{m}[\mathsf{M}_{G}]\right\|_{1}\geq\frac{f(G)}{8}-\frac{4}{m}.$
For example, the values of the parameter $f$ on the full graphs in Figures 1-2 are $0.0073$ and $0.010$, respectively. Estimating this parameter by sampling $k=10^{3}$ edges, we obtain values of $0.0084$ and $0.012$, respectively. For comparison, the value of this parameter on an Erdős–Rényi random graph with edge probability $1/2$ on 1,500 vertices is $0.001$.
4.4 Missing Proofs from Section 4
In this section we give all the proofs missing from Section 4.
Proofs from Section 4.2
In this section, we prove the two lemmas used in Section 4.2 to prove Theorem 4.3. We begin by proving Lemma 4.4 by constructing an explicit coupling between $\mathsf{M}$, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}$, and $\mathsf{G}^{(n)}[\mathsf{M}]$ with the claimed properties.
Proof (Lemma 4.4)..
Sample $\mu\sim\mathsf{M}$ and set $\mathsf{M}^{\prime}=\mu$. Next, sample $(x_{i},y_{i})\overset{iid}{\sim}\mu$, let $\{t_{1},\ldots,t_{N}\}=\{x_{i}\}\cup\{y_{i}\}$, form $\mathsf{G}^{(n)}$ as in Section 4.1, and set
$\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}=\sum_{i,j=1}^{N}(\mathsf{G}^{(n)})_{i,j}\delta_{(t_{i},t_{j})}.$ (16)
We proceed to argue that this coupling satisfies all the claimed properties.
By construction, we have $\mathsf{M}^{\prime}\overset{d}{=}\mathsf{M}$ and $\mathsf{G}^{(n)}\overset{d}{=}\mathsf{G}^{(n)}[\mathsf{M}]$. Next, because $\mathsf{M}$ is exchangeable, we claim that $t_{1},\ldots,t_{N}\overset{iid}{\sim}\mathrm{Unif}([0,1])$. Indeed, $(x_{i},y_{i})\overset{d}{=}(\sigma(x_{i}),\sigma(y_{i}))$ for all $\sigma\in S_{[0,1]}$, hence $(x_{i},y_{i})\sim\theta\lambda^{2}+(1-\theta)\lambda_{D}$. Therefore, for each $i$ either $x_{i}=y_{i}\sim\mathrm{Unif}([0,1])$ or $x_{i},y_{i}\overset{iid}{\sim}\lambda^{2}$. Combining this with the fact that the $(x_{i},y_{i})$ are independent for different $i$, we obtain the claim. Therefore, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}\overset{d}{=}\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}=\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$, the latter being an empirical measure drawn from $\mu\sim\mathsf{M}$. Finally, we have $\mathsf{M}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}_{\mathsf{G}^{(n)}}$ as can be seen from (16), as $\mathsf{G}^{(n)}$ only depends on the multiplicities of $\{(t_{i},t_{j})\}$ rather than their locations, and the multiplicities and locations of a random exchangeable measure are independent of each other [kallenberg1990exchangeable, Thm. 2]. ∎
We now turn to proving Lemma 4.5, combining standard results from VC theory.
Proof (Lemma 4.5)..
Fix any $\mu\in\mathcal{P}([0,1]^{2})$. By the proof of [wainwright2019high, Thm. 4.10], we have
$\underset{\begin{subarray}{c}\mu^{(n)}\\ \textrm{emp. meas.}\sim\mu\end{subarray}}{\mathbb{E}}\sup_{R\in\mathcal{AAR}}|\mu_{n}(R)-\mu(R)|\leq 2R_{n}(\mathcal{F}_{\mathcal{AAR}}),$
where $R_{n}(\mathcal{F})=\underset{\begin{subarray}{c}X_{i}\overset{iid}{\sim}\mu\\ \varepsilon_{i}\overset{iid}{\sim}\mathrm{Unif}(\{\pm 1\})\end{subarray}}{\mathbb{E}}\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|$ is the Rademacher complexity of a class of functions $\mathcal{F}$ and we set $\mathcal{F}_{\mathcal{AAR}}=\{\mathbbm{1}_{R}:R\in\mathcal{AAR}\}$. We further have by [wainwright2019high, Ex. 5.24] that
$R_{n}(\mathcal{F})\leq\frac{24}{\sqrt{n}}\int_{0}^{\mathrm{diam}(\mathcal{F})}\mathbb{E}\sqrt{\log N(t;\mathcal{F};\|\cdot\|_{\mu_{n}})}\,dt,$
where $N(t;\mathcal{F};\|\cdot\|_{\mu_{n}})$ is the $t$-covering number of $\mathcal{F}$ in the (random) metric $\|f-g\|_{\mu_{n}}=\sqrt{\mathbb{E}_{\mu_{n}}(f-g)^{2}}$ defined by the empirical measure $\mu_{n}$. If $\mathcal{F}$ is class of $\{0,1\}$-valued functions, then $\mathrm{diam}(\mathcal{F})\leq 1$. Finally, we have by [HAUSSLER1995217, Cor. 1] that
$N(t;\mathcal{F}_{\mathcal{AAR}};\|\cdot\|_{\mu})\leq e(V+1)\left(\frac{2e}{t}\right)^{V},$
for any measure $\mu$, where $V=\mathrm{VC}(\mathcal{AAR})=4$ is the VC-dimension of the class of axis-aligned rectangles. Combining the above results, we conclude that
$\underset{\begin{subarray}{c}\mu^{(n)}\\ \textrm{emp. meas.}\sim\mu\end{subarray}}{\mathbb{E}}\sup_{R\in\mathcal{AAR}}|\mu_{n}(R)-\mu(R)|\leq\frac{48}{\sqrt{n}}\int_{0}^{1}\sqrt{\log(5e)+4\log(2e/t)}\,dt\leq\frac{174}{\sqrt{n}},$
where we numeriucally evaluated the last integral to 12 digits of accuracy to get $3.624000076562\leq 3.625$. ∎
Proof of Proposition 3.6 from Section 3.2
Using Theorem 4.3 and its proof, we can now prove Proposition 3.6 by comparing both the quotient-convergent sequence $(G_{n})$ and its limit $\mathsf{M}$ to their edge-sampled random graphs.
Proof (Proposition 3.6)..
Since $(G_{n})$ quotient-converges to $\mathsf{M}$, we have $W_{1}(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}],\mathsf{G}^{(k)}[\mathsf{M}])\to 0$ as $n\to\infty$ for each $k\in\mathbb{N}$. Pick $k\geq(174/\epsilon)^{2}$, sample $(G_{n}^{(k)},G_{\mathsf{M}}^{(k)})$ from a coupling of $\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}]$ and $\mathsf{G}^{(k)}[\mathsf{M}]$ attaining $W_{1}(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}],\mathsf{G}^{(k)}[\mathsf{M}])$, and note that with probability $1-4e^{-2\epsilon^{2}}-\frac{W_{1}(\mathsf{G}^{(k)}[\mathsf{M}_{G_{n}}],\mathsf{G}^{(k)}[\mathsf{M}])}{\epsilon}$ we have
$\sup_{R\in\mathcal{AAR}}|\mu_{G_{n}^{(k)}}(R)-\mu_{G_{n}}(R)|\leq\epsilon,\quad\sup_{R\in\mathcal{AAR}}|\mu_{G_{\mathsf{M}}^{(k)}}(R)-\mu_{\mathsf{M}}(R)|\leq\epsilon,\quad\textrm{and}\quad\|G_{n}^{(k)}-G_{\mathsf{M}}^{(k)}\|_{1}\leq\epsilon,$ (17)
for $(\mu_{G_{\mathsf{M}}^{(k)}},\mu_{\mathsf{M}})\sim(\mathsf{M}_{G_{\mathsf{M}}^{(k)}}^{\prime},\mathsf{M}^{\prime})$ and $(\mu_{G_{n}^{(k)}},\mu_{G_{n}})\sim(\mathsf{M}_{G_{n}^{(k)}}^{\prime},\mathsf{M}_{G_{n}}^{\prime})$ sampled from the coupling of Lemma 4.4 conditioned on $G_{\mathsf{M}}^{(k)}$ and $G_{n}^{(k)}$, respectively. Picking $n$ large enough, the above occurs with positive probability, so we can find deterministic $G_{n}^{(k)}$ and $G_{\mathsf{M}}^{(k)}$ satisfying (17) for all large $n$. Explicitly, by Lemma 4.4 there are locations $t_{i}\in[0,1]$ and a permutation $\pi_{n}\in\mathsf{S}_{n}$ satisfying
$\mu_{G_{n}}=\sum_{i,j=1}^{n}(G_{n})_{i,j}\delta_{(t_{i},t_{j})},\quad\mu_{G_{n}^{(k)}}=\sum_{i,j=1}^{n}(\pi_{n}G_{n}^{(k)})_{i,j}\delta_{(t_{i},t_{j})}.$
By considering rectangles in (17) centered at each of the atoms of $\mu_{G_{n}}$ and $\mu_{G_{n}^{(k)}}$, and around rows and columns of these two measures, we conclude that
$\|G_{n}-\pi_{n}G_{n}^{(k)}\|_{\infty}\leq\epsilon,\quad\|G_{n}\mathbbm{1}-\pi_{n}G_{n}^{(k)}\mathbbm{1}\|_{\infty}\leq\epsilon,\quad\textrm{and}\quad\|G_{n}^{\top}\mathbbm{1}-\pi_{n}(G_{n}^{(k)})^{\top}\mathbbm{1}\|_{\infty}\leq\epsilon.$ (18)
Similarly, we can find a permutation $\widetilde{\pi}_{M}\in\mathsf{S}_{M}$ for large enough $M\in\mathbb{N}$ and atoms $(t_{i})_{i\in\mathbb{N}}$ and $(t_{i}^{\prime})_{i=1}^{M-m}$ for some $m\leq M$ satisfying
$\mu_{\mathsf{M}}=\sum_{i,j\in\mathbb{N}}E_{i,j}\delta_{(t_{i},t_{j})}+\sum_{i\in\mathbb{N}}\sigma_{i}(\delta_{t_{i}}\otimes\lambda)+\sum_{i\in\mathbb{N}}\varsigma_{i}(\lambda\otimes\delta_{t_{i}})+\theta\lambda^{2}+\vartheta\lambda_{D},$
$\mu_{G_{\mathsf{M}}^{(k)}}=\sum_{i,j=1}^{m}(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})_{i,j}\delta_{(t_{i},t_{j})}+\sum_{i=1}^{m}\sum_{j=m+1}^{M}(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})_{i,j}\delta_{(t_{i},t_{j}^{\prime})}+\sum_{i=m+1}^{M}\sum_{j=1}^{m}(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})_{i,j}\delta_{(t_{i}^{\prime},t_{j})}$
$\qquad\qquad+\sum_{i,j=m+1}^{M}(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})_{i,j}\delta_{(t_{i}^{\prime},t_{j}^{\prime})}.$
Here we have separated the edges sampled from each summand in $\mathsf{M}$. Moreover, with probability 1 we have $(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})_{i,j}\in\{0,1/k\}$ if $i>m$ or $j>m$. Note that $k^{-1}\leq(\epsilon/174)^{2}\leq\epsilon$ for all small $\epsilon>0$. By considering similar rectangles in (17), we therefore
$\|E-\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)}\|_{\infty}\leq\epsilon,\quad\|E\mathbbm{1}+\sigma-\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)}\mathbbm{1}\|_{\infty}\leq\epsilon,\quad\textrm{and}\quad\|E^{\top}\mathbbm{1}+\varsigma-(\widetilde{\pi}_{M}G_{\mathsf{M}}^{(k)})^{\top}\mathbbm{1}\|_{\infty}\leq\epsilon.$ (19)
Finally, since $\|G_{n}^{(k)}-G_{\mathsf{M}}^{(k)}\|_{1}\leq\epsilon$, we conclude that
$\|G_{n}^{(k)}-G_{\mathsf{M}}^{(k)}\|_{\infty}\leq\epsilon,\quad\|G_{n}^{(k)}\mathbbm{1}-G_{\mathsf{M}}^{(k)}\mathbbm{1}\|_{\infty}\leq\epsilon,\quad\textrm{and}\quad\|(G_{n}^{(k)})^{\top}\mathbbm{1}-(G_{\mathsf{M}}^{(k)})^{\top}\mathbbm{1}\|_{\infty}\leq\epsilon.$ (20)
Combining (18), (19), and (20), we obtain
$\|\widetilde{\pi}_{M}\pi_{n}^{-1}G_{n}-E\|_{\infty}\leq 3\epsilon,\quad\|\widetilde{\pi}_{M}\pi_{n}^{-1}G_{n}\mathbbm{1}-(E\mathbbm{1}+\sigma)\|_{\infty}\leq 3\epsilon,\quad\textrm{and}\quad\|\widetilde{\pi}_{M}\pi_{n}^{-1}G_{n}^{\top}\mathbbm{1}-(E^{\top}\mathbbm{1}+\varsigma)\|_{\infty}\leq 3\epsilon.$
Since $\epsilon>0$ was arbitrary, our first three claims are proved. For the last claim, observe that $\lim_{n}\mathrm{Tr}(G_{n})=\lim_{n}\mathrm{hom}([1],G_{n})$ exists as it is the limit of the homomorphism numbers of a single self-loop. By Corollary 3.4 and the proof of Proposition 2.6, the limit of the homomorphism numbers is the same for any sequence of graphs quotient-converging to $\mathsf{M}$. By considering the sequence of graphs (13) derived from $\mathsf{M}$, we conclude that $\lim_{n}\mathrm{Tr}(G_{n})=\mathrm{Tr}(E)+\vartheta$, proving the last claim. ∎
5 Equipartition-Consistent Random Graph Models
Each grapheur $\mathsf{M}$ yields a sequence of random graphs $(\mathsf{G}_{k}[\mathsf{M}])$ via the construction (8). In Section 5.1, we show that these graphs are related to each other via equipartitions, and hence form an equipartition-consistent random graph model (Definition 1.8). We then prove that realizations of these random graphs $(\mathsf{G}_{k}[\mathsf{M}])$ converge back to the grapheur $\mathsf{M}$ in a suitable fashion. Finally, in Section 5.2 we prove Theorem 1.9, stating that any equipartition-consistent random graph model is a mixture of ones obtained from grapheurs as above.
5.1 Equipartition Consistency
We begin by arguing that $(\mathsf{G}_{k}[\mathsf{M}])$ is equipartition-consistent. Note that $d_{k,nk}\circ F_{nk,m}\overset{d}{=}F_{k,m}$ for any $k,n,m\in\mathbb{N}$ by [levin2025deFin, Lemma 4.13], since each $i\in[m]$ is mapped independently by $F_{nk,m}$ to one of the fibers of $d_{k,nk}$ with the same probability. Therefore, for any $G\in\Delta^{n\times n}$ the sequence of distributions $(\mathrm{Law}(\rho(F_{k,n})G_{n}))_{k}$ is indeed an equipartition-consistent model. If $(G_{m})$ is a sequence of finite graphs converging to a grapheur $\mathsf{M}$, then $\rho(F_{k,m})G_{m}\to\mathsf{G}_{k}[\mathsf{M}]$ and $\rho(F_{nk,m})G_{m}\to\mathsf{G}_{nk}[\mathsf{M}]$ weakly, hence
$\mathrm{Law}(\rho(d_{k,nk})\mathsf{G}_{nk}[\mathsf{M}])=\lim_{m}\mathrm{Law}(\rho(d_{k,nk}\circ F_{nk,m})G_{m})=\lim_{m}\mathrm{Law}(\rho(F_{k,m})G_{m})=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]),$
showing that $(\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$ is equipartition-consistent for any grapheur $\mathsf{M}$.
We now ask whether samples from an equipartition-consistent model $(\mathsf{G}_{k}[\mathsf{M}])$ obtained from a grapheur $\mathsf{M}\in\mathcal{M}$ converge back to $\mathsf{M}$. The answer depends on how this sampling is performed.
Proposition 5.1.
Fix a grapheur $\mathsf{M}\in\mathcal{M}$ and consider the associated equipartition-consistent random graph model $(\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$.
  1. 1.
    Draw $\mu\sim\mathsf{M}$ and let $\mathsf{G}_{k}=(\mu(I_{i}^{(n)}\times I_{j}^{(n)}))_{i,j\in[k]}$ for each $k\in\mathbb{N}$. Then $G_{k}\sim\mu_{k}$ for all $k\in\mathbb{N}$, and $(G_{k})$ is quotient-convergent to $\mathsf{M}$ almost surely.
  2. 2.
    Suppose $\mathsf{M}=\delta_{(T_{1},T_{2})}$ is the grapheur associated to a single edge. Draw $G_{k}\sim\mu_{k}$ independently of each other. Then $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely, so $(G_{k})$ does not converge to $\mathsf{M}$.
Proof.
For the first claim, fix any $\epsilon>0$ and let $N\in\mathbb{N}$ be sufficiently large so that $\sum_{i>N\textrm{ or }j>N}E_{i,j}+\sum_{i>N}(\sigma_{i}+\varsigma_{i})\leq\epsilon$. Drawing $\mu\sim\mathsf{M}$ amounts to sampling locations $T_{i}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ in (5), and these samples are distinct almost surely. Therefore, there exists $K\geq N\in\mathbb{N}$ such that $T_{1},\ldots,T_{N}$ lie in distinct intervals $(I_{i}^{(k)})_{i\in[k]}$ for all $k\geq K$. Therefore, for all $k\geq K$ we can find a permutation $\pi_{k}\in\mathsf{S}_{k}$ such that $\pi_{k}G_{k}=\rho(\iota_{k,N})(E_{i,j}+\sigma_{i}/k+\varsigma_{j}/k)_{i,j\in[N]}+N_{k}+\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+\vartheta\frac{1}{k}I_{k}$ where $\|N_{k}\|_{1}\leq\epsilon$. Using the coupling $(\mathsf{M}_{G_{k}}^{\prime},\mathsf{M}^{\prime})$ obtained by drawing random locations $(T_{i})$, setting $\mathsf{M}^{\prime}$ to (5), and setting $\mathsf{M}^{\prime}_{G_{k}}=\sum_{i,j=1}^{N}(\pi_{k}G_{k})_{i,j}\delta_{(T_{i},T_{j})}$, we conclude that $\limsup_{k\to\infty}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\leq\epsilon$. Since $\epsilon>0$ was arbitrary, we obtain the claimed convergence.
For the second claim, note that with probability $1/k$ we have $T_{1},T_{2}\in I_{i}^{(k)}$ for the same $i\in[k]$, in which case $G_{k}=\mathrm{diag}(0,\ldots,1,\ldots,0)$ and $W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ as can be seen by setting $S=[0,1/2]$ and $T=[1/2,1]$ in (7). Since these events are independent and the sum of their probabilities $\sum_{k}k^{-1}$ diverges, Borel–Cantelli implies that $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely. ∎
In words, the quotients must be sampled in a manner that is related across different sizes for the samples to converge to the underlying grapheur; if we instead sample random quotients independently for each size, the resulting sequence almost surely does not converge to the underlying grapheur. In contrast, if we sample growing-sized random graphs from a graphon independently for different sizes, then the resulting sequence does converge back to the graphon almost surely as a consequence of Szemerédi’s regularity lemma [lovasz2012large, Lemma 11.8]. Proposition 5.1(2) thus implies that the analog of Theorem 1.7 fails for random quotients. More precisely, for any rate of convergence $(R(k)\in\mathbb{R}_{\geq 0})_{n\in\mathbb{N}}$ with $R(k)\to 0$, we have $\sum_{k\in\mathbb{N}}\mathbb{P}[W_{\square}(\mathsf{M},\mathsf{G}_{k}[\mathsf{M}])>R(k)]=\infty$ by Borel–Cantelli. We leave the precise analysis of convergence rates of random quotients for future work.
5.2 Proof of Theorem 1.9
Our goal in this section is to prove Theorem 1.9, showing that every equipartition-consistent random graph model arises from a unique random exchangeable measure. We deduce the existence of the desired random measure from results on inverse limits of random histograms [orbanz2011projective], and proving that this random measure is exchangeable proceeds via standard approximation arguments. To apply the results of [orbanz2011projective], we require the following lemma on the sequence of means of an equipartition-consistent model.
Lemma 5.2.
For any equipartition-consistent random graph model $(\mu_{k})$, there exists $\theta\in[0,1]$ satisfying $\mathbb{E}_{X\sim\mu_{k}}X=\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+(1-\theta)\frac{1}{k}I_{k}$ for all $k\in\mathbb{N}$.
Proof.
Note that $M_{n}=\mathbb{E}_{X\sim\mu_{n}}X\in\mathbb{S}^{n}$ is a deterministic matrix invariant under simultaneous permutations of its rows and columns, because $\mu_{n}$ is exchangeable. Any such matrix has the same on-diagonal and off-diagonal entries, hence there exist $\theta_{n},\vartheta_{n}\in\mathbb{R}$ satisfying $M_{n}=\theta_{n}\frac{1}{n}\mathbbm{1}_{n\times n}+\vartheta_{n}\frac{1}{n}I_{n}$. Since $\mathbbm{1}^{\top}X\mathbbm{1}=1$ almost surely, we similarly have $\mathbbm{1}^{\top}M_{n}\mathbbm{1}=\theta_{n}+\vartheta_{n}=1$. Also, since $X\geq 0$ almost surely, we must have $\theta_{n}\in[0,1]$. Finally, since $\rho(d_{n,N})X_{N}\overset{d}{=}X_{n}$ whenever $n|N$, we must have
$M_{n}=\rho(d_{n,N})M_{N}=\theta_{N}\tfrac{1}{n^{2}}\mathbbm{1}_{n\times n}+(1-\theta_{N})\tfrac{1}{n}I_{n}.$
Since $\frac{1}{n^{2}}\mathbbm{1}_{n\times n}$ and $\frac{1}{n}I_{n}$ are linearly independent for all $n\geq 2$, we must have $\theta_{N}=\theta_{n}$ whenever $n|N$, and therefore for all $n,N\in\mathbb{N}$. This proves the claim. ∎
Combining the above lemma with [orbanz2011projective], we are ready to prove Theorem 1.9.
Proof (Theorem 1.9)..
Uniqueness is clear since squares of the form $I_{i}^{(k)}\times I_{j}^{(k)}$ generate the Borel $\sigma$-algebra on $0,1]^{2}$. To prove existence we appeal to [orbanz2011projective, Thm. 1.1], which states that there exists a random probability measure $\mathsf{M}$ on $0,1]^{2}$ (not guaranteed to be exchangeable) satisfying $\mathsf{M}(I_{i}^{(k)}\times I_{j}^{(k)})\sim\mu_{k}$ for all $k$ if there is a determinisitic measure $\mu\in\mathcal{P}([0,1]^{2})$ satisfying $(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}=\mathbb{E}_{X\sim\mu_{k}}X$ for all $k\in\mathbb{N}$. Such a deterministic measure exists by Lemma 5.2, since
$((\theta\lambda^{2}+(1-\theta)\lambda_{D})(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}=\theta\tfrac{1}{k^{2}}\mathbbm{1}_{k\times k}+(1-\theta)\tfrac{1}{k}I_{k}.$
We proceed to prove that the above $\mathsf{M}$ is exchangeable. To do so, it suffices to prove that for any continuous function $f\colon[0,1)^{2}\to\mathbb{R}$ and any measure-preserving bijection $\sigma$, we have $\mathbb{E}_{\mathsf{M}}f\overset{d}{=}\mathbb{E}_{\mathsf{M}}f\circ(\sigma,\sigma)$. By [kallenberg2005probabilistic, Prop. 9.1], it further suffices to prove this equality for $\sigma\in\mathsf{S}_{m}$, a permutation of the $m$ intervals $I_{1}^{(m)},\ldots,I_{m}^{(m)}$ acting by $\sigma((i-1)/m+x)=(\sigma(i)-1)/m+x$ for $x\in[0,1/m)$ for $i\in[m]$.
For any continuous function $f\colon[0,1]^{2}\to\mathbb{R}$, let $f_{n}=\sum_{i,j=1}^{n}f_{i,j}^{(n)}\mathbbm{1}_{I_{i}^{(n)}\times I_{j}^{(n)}}$ where $f_{i,j}=f(\frac{i}{n}-\frac{1}{2n},\frac{j}{n}-\frac{1}{2n})$ is the value of $f$ on the centers of the squares $I_{i}^{(n)}\times I_{j}^{(n)}$, whose diameter in $\ell_{\infty}$ is $1/2n$. Since $f$ is continuous on a compact set, it is uniformly continuous, hence for any $\epsilon>0$ we have $|f(x)-f_{n}(x)|\leq\epsilon$ for all $x\in[0,1)^{2}$ and all large $n$. Note that
$\mathbb{E}_{\mathsf{M}}f_{n}=\sum_{i,j=1}^{n}f_{i,j}^{(n)}\mathcal{M}(I_{i}^{(n)}\times I_{j}^{(n)})=\sum_{i,j=1}^{n}f_{i,j}^{(n)}(\mathsf{G}_{n}[\mathsf{M}])_{i,j}.$
Therefore, for any $m,k\in\mathbb{N}$ and $\sigma_{m}\in\mathsf{S}_{m}$, we view $\sigma_{m}$ as an interval permutation and as a permutation on $[mk]$ letters permuting consecutive intervals of length $k$, in which case we get
$\mathbb{E}_{\mathsf{M}}f_{mk}\circ(\sigma_{m},\sigma_{m})=\sum_{i,j=1}^{m}f_{i,j}^{(mk)}(G_{mk}^{(\mathsf{M})})_{\sigma^{-1}(i),\sigma^{-1}(j)}\overset{d}{=}\sum_{i,j=1}^{m}f_{i,j}^{(mk)}(G_{mk}^{(\mathsf{M})})_{i,j}=\mathbb{E}_{\mathsf{M}}f_{mk}.$
Since this holds for all $k\in\mathbb{N}$, we have
$W_{1}(\mathbb{E}_{\mathsf{M}}f,\mathbb{E}_{\mathsf{M}}f\circ(\sigma_{m},\sigma_{m}))\leq$ $W_{1}(\mathbb{E}_{\mathsf{M}}f,\mathbb{E}_{\mathsf{M}}f_{mk})+W_{1}(\mathbb{E}_{\mathsf{M}}f_{mk},\mathbb{E}_{\mathsf{M}}f_{mk}\circ(\sigma_{m},\sigma_{m}))$
$+W_{1}(\mathbb{E}_{\mathsf{M}}f_{mk}\circ(\sigma_{m},\sigma_{m}),\mathbb{E}_{\mathsf{M}}f\circ(\sigma_{m},\sigma_{m}))\leq 2\epsilon,$
for all large $k$, where the last inequality follows since $\mathsf{M}$ is supported on $0,1)^{2}=\bigcup_{i,j}I_{i}^{(n)}\times I_{j}^{(n)}$. Since $\epsilon>0$ was arbitrary, this proves the exchangeability of $\mathsf{M}$. ∎
6 Conclusions and Future Directions
We have introduced a new notion of limits of growing graphs based on convergence of their fixed-size random quotients. We showed that this notion of convergence can be equivalently characterized in terms of convergence of graph parameters such as homomorphism numbers and quotient densities. We characterized limits of quotient-convergent graph sequences via grapheurs, which are certain random exchangeable measures on $0,1]^{2}$. The limiting grapheur of a quotient-convergent sequence of graphs describes the asymptotic distribution of edge weight in the graphs. We then presented another equivalent view of quotient convergence via edge sampling, allowing us to prove a dual analog of Szemerédi’s regularity lemma and to test properties of arbitrarily-large graphs by sampling edges from them. The number of edges needed is indpendent of the size of the graph, only depending on the desired accuracy and the Lipschitz constant of the graph parameter in question. Finally, we studied equipartition-consistent random graph models, showing that any such model corresponds to the sequence of random quotients associated to a mixture of grapheurs. We conclude by mentioning several questions suggested by our work.
  1. item(Optimal Szemerédi rate)
    Is the rate of $O(1/\sqrt{n})$ for approximating a grapheur by a graph on $n$ edges optimal?
  2. item(Rates for random quotients)
    Is there a Szemerédi-type regularity lemma for random quotients, i.e., does $\mathbb{E}_{\mathsf{G}_{k}[\mathsf{M}]}W_{\square}(\mathsf{M}_{\mathsf{G}_{k}[\mathsf{M}]},\mathsf{M})$ converge to zero at a universal rate independent of $\mathsf{M}$? If not, how does this rate depend on $\mathsf{M}$?
  3. item(Extending graphon duality)
    Is Conjecture 3.11 on the bi-continuity of our duality pairing true? Can the duality with graphons in Section 3.3 be extended to limits of non-simple graphs? Can we relate the cut metric and our $W_{\square}$-metric via this duality?
  4. item(Inequalities in quotient densities)
    Corollary 3.5 shows that valid inequalities in quotient densities (or equivalently, homomorphism numbers) that hold for graphs of all sizes correspond to inequalities over grapheurs. Can we exploit the structure of the space of grapheurs to prove such inequalities, in analogy with the use of graphon theory to prove homomorphism density inequalities?
    Acknowledgements
    Eitan Levin and Venkat Chandrasekaran were supported in part by AFOSR grant FA9550-23-1-0070 and by NSF grant DMS-2502377.

### v2

@@
 We call our limits grapheurs and show that these are dual to graphons in a precise sense.
-Grapheurs are well-suited to modeling hubs and connections between them in large graphs; previous notions of graph limits based on subgraph densities fail to adequately model such global structures as subgraphs are inherently local.
-Using our framework, we present an edge-based sampling approach for testing properties pertaining to hubs in large graphs.
-This method relies on an edge-based analog of the Szemerédi regularity lemma, whereby we show that sampling a small number of edges from a large graph approximately preserves its quotients.
+Grapheurs are well-suited to modeling global structure in large graphs such as hubs and connections between them; previous notions of graph limits based on subgraph densities fail to adequately model such global structure as subgraphs are inherently local.
+Using our framework, we characterize properties of large graphs that are continuous with respect to our limits and present an edge-based sampling approach for testing them.
+This method relies on an edge-based analog of the Szemerédi regularity lemma, whereby we show that sampling a constant number of edges from an arbitrarily-large graph approximately preserves its quotients.
 Finally, we observe that the random quotients of a graph are related to each other by equipartitions, and we conclude with a characterization of such random graph models.
@@
 1 Introduction
-Analyzing the structure of large graphs is a challenge that arises in numerous problem domains such as bioinformatics, network tomography, operations research, and the social sciences [newman_networks, network_tomography, social_network, JACKSON2002265, bassett2017network].
+Analyzing the structure of large graphs is a challenge that arises in numerous problem domains such as bioinformatics, network tomography, operations research, and the social sciences [36, 13, 42, 22, 2].
 Graphons and other related notions of graph limits in the literature employ subgraph densities to summarize and compare large graphs.
-However, subgraph-based measures are inherently local and they fail to suitably account for global structures such as hubs and connections between them, which are commonly encountered in real-world networks [OKELLY1998171, hub_and_spoke, power2013evidence, hubs_in_drugs, van2013network].
+However, subgraph-based measures are inherently local and they fail to suitably account for global structures such as hubs and connections between them, which are commonly encountered in real-world networks [37, 9, 40, 21, 43].
 For example, most subgraphs of a star, the prototypical example of a hub, do not contain any edges.
@@
 Here we view $\rho(f_{k,n}):\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{k\times k}$ as a ‘linear representation’ of $f_{k,n}$.
-In words, the vertices of the quotient $\rho(f_{k,n})G$ correspond to the parts in the partition defined by $f_{k,n}$, and the edge weight between parts $i,j\in[k]$ is given by sums of all edge weights in $G$ between vertices in parts $f_{k,n}^{-1}(i)$ and $f_{k,n}^{-1}(j)$.
+In words, the vertices of the quotient $\rho(f_{k,n})G$ correspond to the parts in the partition defined by $f_{k,n}$, and the edge weight between parts $i,j\in[k]$ is given by the sum of all edge weights in $G$ between vertices in $f_{k,n}^{-1}(i)$ and $f_{k,n}^{-1}(j)$.
 Thus, a quotient of $G$ represents the edge weights between each pair of parts in a vertex partition.
 As an illustration, we give in Figures 1-2 examples of small quotients of two networks.
-The first, depicted in Figure 1, is a traffic network representing the fraction of trips starting and ending at each zone of the Gold Coast, Australia [traffic_networks].
+The first, depicted in Figure 1, is a traffic network representing the fraction of trips starting and ending at each zone of the Gold Coast, Australia [41].
 The second, depicted in Figure 2, is a brain connectome, the network of connections between neurons of an adult male C.
-elegans roundworm [cook2019whole, worms_brain].
+elegans roundworm [14, 35].
 The structure of each quotient indicates some of the features of the underlying large network.
@@
 Figure 1: Traffic network for the Gold Coast, Australia, and an associated random quotient.
-The city is divided into 1068 zones, and the weight of an edge is the fraction of trips from a dataset with given start and end zones [traffic_networks].
+The city is divided into 1068 zones, and the weight of an edge is the fraction of trips from a dataset with given start and end zones [41].
 Here we plot the first 300 of these zones and their interconnections, displaying the width of an edge proportionally to its weight.
-We color the vertices according to their image in the quotient, and color and edge connecting two vertices with the same image by the color of that image.
+We color the vertices according to their image in the quotient, and color an edge connecting two vertices with the same image by the color of that image.
 The nonuniformity in the edge weights of the quotient indicates the presence of hubs that we can see in the plot of the entire network.
 Figure 2: Network of chemical connections among the 575 neurons of adult male C.
-elegans (a species of roundworm), and its random quotient.
-Edges are directed from pre-synaptic to post-synaptic cells, and their weight is the total number of EM serial sections of connectivity [cook2019whole, worms_brain].
+elegans (a species of roundworm), and a random quotient.
+Edges are directed from pre-synaptic to post-synaptic cells, and their weight is the total number of EM serial sections of connectivity [14, 35].
 The nonuniform distribution of self-loops and the unequal in-degrees and out-degrees among the four vertices in the quotient again indicate the presence of hubs.
 In this paper, we define a new notion of convergence for sequences of growing-sized graphs based on convergence of their random quotients.
-In analogy with graphons [lovasz2012large, §1], we define suitable limit objects in terms of random measures for any quotient-convergent sequence of graphs.
+In analogy with graphons [34, §1], we define suitable limit objects in terms of random measures for any quotient-convergent sequence of graphs.
 Finally, we show that finite (random) graphs sampled from our limit objects are quotient-convergent to the limit almost surely.
@@
 In particular, we shall see that random quotients and random subgraphs, in terms of which graphon convergence is defined, are random linear maps which are adjoints of each other.
-Further, a dual analog of the cut metric plays a prominent role in our development, and we derive a dual Szemerédi lemma as a consequence of our results.
+Further, a dual analog of the cut metric plays a prominent role in our development, and we derive an edge-based analog of the Szemerédi regularity lemma as a consequence of our results.
 We describe next these contributions in more detail.
 1.1 Our Contributions
-To set the stage, we define a random quotient of a graph on $n$ vertices in terms of a uniformly random map $F_{k,n}\colon[n]\to[k]$.
-That is, each $i\in[n]$ gets sent independently to a uniform element $F_{k,n}(i)\in[k]$.
+To set the stage, we define a random quotient of a graph on $n$ vertices in terms of a uniformly random map $F_{k,n}\colon[n]\to[k]$ sending each $i\in[n]$ independently to a uniformly-random element $F_{k,n}(i)\in[k]$.
 Equivalently, this is a uniformly random ordered partition of $[n]$ into $k$ parts, given by the sequence of (random) fibers $F_{k,n}^{-1}(j)$ for $j\in[k]$.
@@
 Unless specified otherwise, we will assume throughout that the total edge weight of our graphs equals one.
-Note that $\rho(F_{n,k})^{\star}G_{n}$ is the induced subgraph on $k$ vertices sampled with replacement from the vertex set $[n]$ of $G_{n}$, in terms of which graphon convergence is defined [lovasz2012large, §9], giving a first sense in which our theory is dual to that of graphons.
-We remark that there is an alternative view of graphon limits via convergence of quotients [lovasz2012large, §12], but these latter quotients differ from ours.
+Denoting by $\rho(F_{n,k})^{\star}\colon\mathbb{R}^{n\times n}\to\mathbb{R}^{k\times k}$ the adjoint of $\rho(F_{n,k})$ with respect to the Frobenius inner product, note that $\rho(F_{n,k})^{\star}G_{n}$ is the induced subgraph on $k$ vertices sampled with replacement from the vertex set $[n]$ of $G_{n}$, in terms of which graphon convergence is defined [34, §9].
+This is a first sense in which our theory is dual to that of graphons.
+We remark that there is an alternative view of graphon limits via convergence of quotients [34, §12], but these latter quotients differ from ours.
 In particular, these latter quotients cannot detect hubs in a large graph, as we explain in Section 1.2 below.
@@
 Whereas previous notions of graph convergence were based on convergence of local “motifs” in the form of homomorphism densities (appropriately normalized), we show that our quotient convergence is equivalent to convergence of homomorphism numbers.
-More precisely, for a multigraph $H\in\mathbb{N}^{k\times k}$ possibly containing parallel edges and self-loops, and a general weighted graph $G\in\mathbb{R}^{n\times n}$ on $n$ vertices, the homomorphism number and homomorphism density of $H$ in $G$ are given by [lovasz2012large, Eq.
+More precisely, for a multigraph $H\in\mathbb{N}^{k\times k}$ possibly containing parallel edges and self-loops, and a general weighted graph $G\in\mathbb{R}^{n\times n}$ on $n$ vertices, the homomorphism number and homomorphism density of $H$ in $G$ are given by [34, Eq.
 (5.3)]
@@
 The terminology comes from the fact that when $H$ and $G$ are simple graphs, then $\mathrm{hom}(H;G)$ is the number of graph homomorphisms from $H$ to $G$ and $t(H;G)$ counts the average number of times $H$ occurs in a uniformly random subgraph of $G$.
-In particular, convergence of the homomorphism densities $t(H;\cdot)$ is used to define limits of dense graphs [LOVASZ2006933, convergent_seqs1].
+In particular, convergence of the homomorphism densities $t(H;\cdot)$ is used to define limits of dense graphs [32, 4].
 In analogy, we define the quotient density $t_{Q}(H;G)$ of a multigraph $H\in\mathbb{N}^{k\times k}$ in a general weighted graph $G\in\mathbb{R}^{n\times n}$ as the average number of times $H$ occurs in a uniformly random quotient of $G$:
 $t_{Q}(H;G)=\frac{1}{k^{n}}\sum_{f\colon[n]\to[k]}\prod_{i,j=1}^{k}(\rho(f)G)_{i,j}^{H_{i,j}}.$ (3)
-The role of homomorphism densities in graphons is played by quotient densities in our framework. We prove that a sequence of graphs is quotient-convergent if all of their quotient densities converge.
+The role of homomorphism densities in graphon theory is played by quotient densities in our framework.
+We prove that a sequence of graphs is quotient-convergent if all of their quotient densities converge.
 Theorem 1.2.
@@
     For each multigraph $H\in\mathbb{N}^{k\times k}$, the sequence of homomorphism numbers $(\mathrm{hom}(H;G_{n}))_{n}$ converges.
-The proof of this result is given in Section 2 based on expressing quotient densities in terms of (injective) homomorphism numbers.
+The proof of this result is given in Section 2, and is based on expressing quotient densities in terms of (injective) homomorphism numbers.
 This expression involves a partial order on the collection of multigraphs induced by the quotient operation.
@@
 1.1.2 Limit Objects for Quotient Convergence (Section 3)
-Our second contribution consists of constructing analytic limit objects together with a notion of distance between them that allows us to compare graphs of different sizes, akin to graphons and the cut metric for dense graph sequences.
+Our second contribution consists of constructing analytic limit objects for quotient-convergent sequences together with a notion of distance between them that allows us to compare graphs of different sizes, akin to graphons and the cut metric for dense graph sequences.
 Note that appending isolated vertices to a graph $G\in\Delta^{n\times n}$ or relabeling its vertices does not change the distribution of its quotients $\rho(F_{k,n})G_{n}$ for each $k\in\mathbb{N}$.
@@
 Note that relabeling the vertices of $G$ or appending isolated vertices to it yields the same random measure in this way.
-More generally, Kallenberg showed that all random exchangeable measures on $0,1]^{2}$ have the form [kallenberg1990exchangeable]:
+More generally, Kallenberg showed that all random exchangeable measures on $0,1]^{2}$ have the form [28]:
 $\mathsf{M}=\sum_{i,j\in\mathbb{N}}E_{i,j}\delta_{(T_{i},T_{j})}+\sum_{i\in\mathbb{N}}[\sigma_{i}(\delta_{T_{i}}\otimes\lambda)+\varsigma_{i}(\lambda\otimes\delta_{T_{i}})]+\theta\lambda^{2}+\vartheta\lambda_{D},$ (5)
@@
 Moreover, the collection of extremal random exchangeable measures, i.e., those that are not mixtures of other such random measures, are precisely the ones for which the coefficients $E,\sigma,\varsigma,\theta,\vartheta$ in (5) can be taken to be deterministic.
+We call such extremal random measures grapheurs, and show that these are precisely limits of quotient-convergent graph sequences.
 Definition 1.4.
@@
 Further, adding zero rows to $E$ and corresponding zero entries to $\sigma$, or zero columns to $E$ and corresponding zero entries to $\varsigma$, does not change the resulting random measure.
-Kallenberg showed in [kallenberg1990exchangeable, Prop.
+Kallenberg showed in [28, Prop.
 3] that these are the only possible ambiguities in the coefficients.
@@
   2. 2.
-    Any sequence of graphs $(G_{n}\in\Delta^{n\times n})$ has a quotient-convergent subsequence.
+    The space of grapheurs $\mathcal{M}$ is compact. In particular, any sequence of graphs $(G_{n}\in\Delta^{n\times n})$ has a quotient-convergent subsequence.
   3. 3.
@@
 Similarly, we can extend quotient densities to grapheurs by setting
-$t_{Q}(H;\mathsf{M})=\mathbb{E}\prod_{i,j=1}^{k}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}^{H_{i,j}}=\int\prod_{i,j\in[k]}\mu(I_{i}^{(k)}\times I_{j}^{(k)})^{H_{i,j}}\,d\mathsf{M}(\mu),$ (9)
-for any multigraph $H\in\mathbb{N}^{k\times k}$, analogously to the extension of homomorphism densities to graphons [lovasz2012large, §7.2].
+$t_{Q}(H;\mathsf{M})=\mathbb{E}\prod_{i,j=1}^{k}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}^{H_{i,j}}=\int\prod_{i,j=1}^{k}\mu(I_{i}^{(k)}\times I_{j}^{(k)})^{H_{i,j}}\,d\mathsf{M}(\mu),$ (9)
+for any multigraph $H\in\mathbb{N}^{k\times k}$, analogously to the extension of homomorphism densities to graphons [34, §7.2].
 In terms of quotient densities, the sequence $(G_{n})$ converges to the grapheur $\mathsf{M}$ if and only if $t_{Q}(H;G_{n})\to t_{Q}(H;\mathsf{M})$ for all multigraphs $H$, see Section 3.1.
@@
 Informally, the array $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ specifies the connections between a discrete collection of hubs, with each entry corresponding to an edge with an asymptotically positive fraction of the total edge weight.
-The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars centered at hubs, with each entry corresponding to the fraction of out-degree or in-degree of a hub from all the edges incident on it with asymptotically vanishing edge weights.
+The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars centered at hubs, with each entry corresponding to the fraction of out-degree or in-degree of a hub from all the edges incident to it with asymptotically vanishing edge weights.
 Lastly, the parameters $\theta$ and $\vartheta$ specify the remaining edge weight that comes from all edges between vertices with asymptotically vanishing degrees and their self-loops.
@@
 We have seen previously that the sampling operators used to define each notion of graph limit, namely, those extracting random subgraphs and random quotients, respectively, are duals of each other.
-We further show in Section 3.3 that the space of graphons endowed with the cut metric is dual to the space $\mathcal{M}$ endowed with the $W_{\square}$ metric (7) in a precise sense.
+We further show in Section 3.3 that the space of graphons endowed with the cut metric is dual to the space $\mathcal{M}$ endowed with the $W_{\square}$ metric (7) as follows.
 Theorem 1.6 (Informal; see Theorem 3.10).
@@
 As our next contribution, we apply our framework to test properties of large graphs based on small summaries of them.
-Specifically, we focus on testing properties given by graph parameters that are continuous with respect to quotient convergence.
+Specifically, we focus on testing properties given by graph parameters that are continuous with respect to quotient convergence, including quotient densities, homomorphism numbers, clustering and centrality coefficients, see Section 4.3.
 While it would be natural to test such properties using random quotients, forming these requires access to the entire large graph, which limits their utility in practice.
@@
 We detail the general sampling procedure in Section 4.1.
-We prove the following convergence results for the sequence of random graphs $(\mathsf{G}^{(n)})$ sampled from a grapheur $\mathsf{M}$ back to $\mathsf{M}$, which are edge-based analogs of the second sampling lemma and Szemerédi’s regularity lemma for graphons [lovasz2012large, Lemma 10.16, Lemma 9.15].
+We prove the following convergence results for the sequence of random graphs $(\mathsf{G}^{(n)})$ sampled from a grapheur $\mathsf{M}$ back to $\mathsf{M}$, which are edge-based analogs of the second sampling lemma and Szemerédi’s regularity lemma for graphons [34, Lemma 10.16, Lemma 9.15].
 Here and throughout the paper, a superscript $G^{(n)}$ denotes a graph on $n$ edges (and hence at most $2n$ vertices) while a subscript $G_{n}$ denotes a graph on $n$ vertices.
@@
 Specifically, we observe that $\rho(d_{k,nk})\mathsf{G}_{nk}[\mathsf{M}]\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}]$ for any equipartition $d_{k,nk}\colon[nk]\to[k]$, i.e., a map with equal-sized fibers.
-We call such sequences of random graphs equipartition-consistent random graph models, which play a role in our framework analogous to that played by consistent random graph models in graphon theory [lovasz2012random].
+We call such sequences of random graphs equipartition-consistent random graph models, which play a role in our framework analogous to that played by consistent random graph models in graphon theory [33].
 Definition 1.8 (Equipartition-consistent models).
@@
 For any equipartition-consistent random graph model $(\mu_{k}\in\mathcal{P}(\Delta^{k\times k}))_{k}$, there is a unique random exchangeable measure $\mathsf{M}$ on $0,1]^{2}$ satisfying $\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}])$ for all $k\in\mathbb{N}$.
-Recall that grapheurs were defined as extremal random exchangeable measures in Definition 1.4.
-By Choquet’s theorem, any random exchangeable measure is therefore a mixture of grapheurs.
-Finally, we show in Proposition 5.1 that the sequence $(\mathsf{G}_{k}[\mathsf{M}])$ of quotients of a grapheur $\mathsf{M}$ converge back to $\mathsf{M}$ in a certain sense.
+Recall that grapheurs are extremal random exchangeable measures by Definition 1.4.
+By Choquet’s theorem, any random exchangeable measure is thus a mixture of grapheurs.
+Finally, we show in Proposition 5.1 that the sequence $(\mathsf{G}_{k}[\mathsf{M}])$ of quotients of a grapheur $\mathsf{M}$ converges back to $\mathsf{M}$ in a certain sense.
 1.2 Related Work
@@
 Dense and degree-bounded graph limits.
-The earliest two types of graph limits pertain to sequences of simple graphs that are dense, whose limits are given by graphons [LOVASZ2006933, convergent_seqs1], and degree-bounded, whose limits are given by graphings [benjamini_schramm, aldous2004objective, elek2007note].
-We refer the reader to [lovasz2012large] for a survey of both limits.
+The earliest two types of graph limits pertain to sequences of simple graphs that are dense, whose limits are given by graphons [32, 4], and degree-bounded, whose limits are given by graphings [3, 1, 18].
+We refer the reader to [34] for a survey of both limits.
 Both of these notions of graph limits are incomparable to ours, as both dense and degree-bounded sequences of simple graphs converge to the same limit in our framework, see Corollary 3.7.
-Further, dense graph limits can be characterized in terms of quotients [lovasz2012large, §12], called “right convergence” of graphons.
+Further, dense graph limits can be characterized in terms of quotients [34, §12], called “right convergence” of graphons.
 However, these quotients are differently normalized compared to ours, and it is the set of quotients that converges in the graphon framework rather than their distribution as in our framework.
@@
 Edge-exchangeable random graphs.
-A class of random graph models closely related to the ones we obtain by sampling edges from a grapheur is called edge-exchangeable [crane2016edge, Crane_Dempsey_2019, NIPS2016_1a0a283b, exchangeable_traits, janson2018edge].
-These random graphs are obtained by adding random edges to a fixed infinite vertex set (often with repetition, yielding multigraphs), and can yield both sparse and dense random graphs [janson2018edge].
-As explained in [janson2010graphons, Rmk.
+A class of random graph models closely related to the ones we obtain by sampling edges from a grapheur is called edge-exchangeable [15, 16, 10, 11, 24].
+These random graphs are obtained by adding random edges to a fixed infinite vertex set (often with repetition, yielding multigraphs), and can yield both sparse and dense random graphs [24].
+As explained in [23, Rmk.
 4.4], every edge-exchangeable multigraph model can be obtained by sampling edges from a random exchangeable measure on $0,1]^{2}$.
@@
 Nevertheless, to our knowledge this notion of convergence of edge-exchangeable models has not been previously studied in the literature.
-In particular, the only metric between edge-exchangeable models we are aware of is an $\ell_{1}$ distance between edge probabilities [Crane_Dempsey_2019, Eq.
+In particular, the only metric between edge-exchangeable models we are aware of is an $\ell_{1}$ distance between edge probabilities [16, Eq.
 (8)].
@@
 Graphex processes and limits.
-The closest line of work to ours is the study of graphex processes and their associated graph limits and random graph models [caron_fox, veitch2015class, jmlr_graphon_proc, borgs2020identifiability, borgs2019sampling, multigraphex, JANSON2022103549].
-Graphexes form a subset of random exchangeable measures on $\mathbb{R}^{2}_{\geq 0}$, and admit a characterization due to Kallenberg similarly to our grapheurs [kallenberg1990exchangeable].
-Graphexes give rise to random graph models obtained by sampling edges from them [veitch2015class, borgs2019sampling, veitch2019sampling], analogously to our edge sampling procedure in Section 4.1.
-There are also several notions of limits for simple graphs and multigraphs, whereby such graphs converge to a limiting graphex [jmlr_graphon_proc, multigraphex].
-These notions of limits are conjectured to be inequivalent [JANSON2022103549], and the space of graphexes with either topology is not compact.
-Some of these limits can be characterized by sampling random subgraphs (whose number of vertices is also random) [borgs2019sampling], while others can be characterized by convergence of differently-normalized homomorphism numbers [jmlr_graphon_proc, §2.5] (the authors of [jmlr_graphon_proc] normalize $\mathrm{hom}(H;G)$ by $(2|E(G)|)^{|V(H)|/2}$, whereas we normalize by $(2|E(G)|)^{|E(H)|}$).
+The closest line of work to ours is the study of graphex processes and their associated graph limits and random graph models [12, 45, 6, 5, 7, 8, 25].
+Graphexes form a subset of random exchangeable measures on $\mathbb{R}^{2}_{\geq 0}$, and admit a characterization due to Kallenberg similarly to our grapheurs [28].
+Graphexes give rise to random graph models obtained by sampling edges from them [45, 7, 44], analogously to our edge sampling procedure in Section 4.1.
+There are also several notions of limits for simple graphs and multigraphs, whereby such graphs converge to a limiting graphex [6, 8].
+These notions of limits are conjectured to be inequivalent [25], and the space of graphexes with either topology is not compact.
+Some of these limits can be characterized by sampling random subgraphs (whose number of vertices is also random) [7], while others can be characterized by convergence of differently-normalized homomorphism numbers [6, §2.5] (the authors of [6] normalize $\mathrm{hom}(H;G)$ by $(2|E(G)|)^{|V(H)|/2}$, whereas we normalize by $(2|E(G)|)^{|E(H)|}$).
 Graphex-based graph limits are again incomparable to ours.
-On the one hand, for a sequence of simple graphs our grapheur-based limits can only distinguish stars from a uniform component (Corollary 3.7), whereas graphex-based limits can capture more general limiting structures (see [JANSON2022103549, §9] or [multigraphex]).
+On the one hand, for a sequence of simple graphs our grapheur-based limits can only distinguish stars from a uniform component (Corollary 3.7), whereas graphex-based limits can capture more general limiting structures (see [25, §9] or [8]).
 On the other hand, grapheur-based graph limits are defined for general sequences of nonnegatively-weighted graphs, for which graphex-based limits have not been defined, allowing us to test properties pertaining to the distribution of edge weight in the graph (see Section 4.3).
@@
 where the infimum is over couplings $(X_{1}^{\prime},X_{2}^{\prime})$ of $X_{1}$ and $X_{2}$, the norm $\|\cdot\|_{1}$ is the entrywise $\ell_{1}$ norm, and the supremum is over 1-Lipschitz functions with respect to this norm.
-Note that a sequence of random graphs $(X_{n})\subseteq\Delta^{k\times k}$ converges weakly to the random graph $X\in\Delta^{k\times k}$ if and only if $\lim_{n}W_{1}(X_{n},X)=0$ by [villani2008optimal, Thm.
+Note that a sequence of random graphs $(X_{n})\subseteq\Delta^{k\times k}$ converges weakly to the random graph $X\in\Delta^{k\times k}$ if and only if $\lim_{n}W_{1}(X_{n},X)=0$ by [46, Thm.
 6.9].
@@
 A sequence of random measures $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if $\mathrm{Law}(\mathsf{M}_{n})\to\mathrm{Law}(\mathsf{M})$ weakly, where these laws are probability measures over the space of probability measures.
-Equivalently, the sequence $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if the sequence of scalar random variables $\mathbb{E}_{\mathsf{M}_{n}}f$ converges weakly to $\mathbb{E}_{\mathsf{M}}f$ for all continuous functions $f$ [kallenberg2005probabilistic, Thm.
+Equivalently, the sequence $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if the sequence of scalar random variables $\mathbb{E}_{\mathsf{M}_{n}}f$ converges weakly to $\mathbb{E}_{\mathsf{M}}f$ for all continuous functions $f$ [26, Thm.
 A2.3].
@@
 To show the second part relating quotient convergence with convergence of homomorphism numbers, we express homomorphism numbers in terms of quotient densities and vice-versa.
-To this end, we begin by expressing quotient densities in terms of injective homomorphism numbers using [levin2025deFin, Prop.
+To this end, we begin by expressing quotient densities in terms of injective homomorphism numbers using [31, Prop.
 6.8].
@@
 Figure 3: Hasse diagram for the refinement poset consisting of multigraphs with two edges. Here there is an upward path from $K$ to $H$ precisely when $K\leq_{R}H$.
-The quotient density $t_{Q}(H;G)$ can be expressed in terms of a linear combination of injective homomorphism numbers over all refinements of $H$ [levin2025deFin, Prop. 6.8]:
+The quotient density $t_{Q}(H;G)$ can be expressed in terms of a linear combination of injective homomorphism numbers over all refinements of $H$ [31, Prop. 6.8]:
 $t_{Q}(H;G)=\sum_{K\leq_{R}H}|V(H)|^{-|V(K)|}\frac{H!}{K!}\frac{R_{K,H}}{R_{K,K}R_{H,H}}\mathrm{inj}(K;G),$ (10)
@@
 The expression (10) states that $H$ is a subgraph of some quotient of $G$, i.e., $t_{Q}(H;G)>0$, if and only if some refinement of $H$ is a subgraph of $G$ itself, i.e., $\mathrm{inj}(K;G)>0$.
-We are now ready to prove the second part of Theorem 1.2. The following argument is a simple modification of the one from [levin2025deFin, Prop. 6.8].
+We are now ready to prove the second part of Theorem 1.2. The following argument is a simple modification of the one from [31, Prop. 6.8].
 Proposition 2.6.
@@
 Based on (11) and (10), we conclude that $(t_{Q}(H;G_{n}))$ converges for all multigraphs $H$ if and only if $(\mathrm{inj}(H;G_{n}))$ converges for all multigraphs $H$.
-Injective and non-injective homomorphism numbers can be similarly related, with one of the quantities being expressed in terms of linear combinations of the other as in [lovasz2012large, §5.2.3].
-For example, we have [levin2025deFin, Eq.
+Injective and non-injective homomorphism numbers can be similarly related, with one of the quantities being expressed in terms of linear combinations of the other as in [34, §5.2.3].
+For example, we have [31, Eq.
 (73)] that
@@
 By Proposition 3.3, this is further equivalent to the existence of a limit $\mathsf{M}\in\mathcal{M}$ for $(\mathsf{M}_{G_{n}})$ in this metric, thus proving the first statement.
-The second statement is a consequence of the compactness of $\mathcal{M}$, which follows from Proposition 3.3.
+The second statement follows from Proposition 3.3.
 Finally, we prove the third statement by explicitly constructing a convergent sequence $(G_{n})$ of graphs for each grapheur $\mathsf{M}\in\mathcal{M}$.
@@
 We end this section by noting that valid polynomial inequalities in quotient densities precisely correspond to valid inequalities over grapheurs.
-We hope that grapheurs can be used to study such inequalities in future work, dually to the use of graphons in the study of inequalities in homomorphism densities [lovasz2012large, §16].
+We hope that grapheurs can be used to study such inequalities in future work, dually to the use of graphons in the study of inequalities in homomorphism densities [34, §16].
 Corollary 3.5 (Inequalities in quotient densities).
@@
 Proposition 3.6 justifies the intuition outlined above: Up to a relabelling of the vertices of $G_{n}$, the entries of $E$ are entrywise limits of the edge weights of $G_{n}$, the entries of $\sigma,\varsigma$ are limits of the remaining degrees of vertices in $(G_{n})$ after removing edges with positive limiting weights, and the parameters $\vartheta,\theta$ capture the remaining edge weights in self loops and the rest of the edges, respectively.
-We prove Proposition 3.6 in Section 4.4 using an edge-sampling perspective on quotient convergence and a dual Szemerédi regularity lemma.
+We prove Proposition 3.6 in Section 4.4 using an edge-sampling perspective on quotient convergence and an edge-based Szemerédi regularity lemma.
 An immediate consequence of Proposition 3.6 is that all dense and degree-bounded sequences of simple graphs have the same limit.
 Corollary 3.7.
-Suppose $(G_{n}\in\mathbb{S}^{n})$ is a sequence of undirected graphs such that the sequence $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1})$ quotient-converges.
+Suppose $(G_{n}\in\mathbb{S}^{n}_{\geq 0})$ is a sequence of undirected graphs such that the sequence $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1})$ quotient-converges to a grapheur described by $(E,\sigma,\varsigma,\theta,\vartheta)$.
   1. 1.
-    The limit is described by $(E,\sigma,\varsigma,\theta,\vartheta)$ with $\sigma=\varsigma$ and $E^{\top}=E$.
+    We have $\sigma=\varsigma$ and $E^{\top}=E$.
   2. 2.
@@
   3. 3.
-    If $G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}}$ is a sequence of simple graphs with $|E(G_{n})|\to\infty$, and is either dense or degree-bounded, then the sequence $(\frac{G_{n}}{2|E(G_{n})|})$ quotient converges to the grapheur $\lambda^{2}$.
+    If $G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}}$ is a sequence of simple graphs with $|E(G_{n})|\to\infty$, then $E=0$.
+If $(G_{n})$ is either dense or degree-bounded, then the sequence $(\frac{G_{n}}{2|E(G_{n})|})$ quotient converges to the grapheur $\lambda^{2}$.
 Proof.
 The first two claims follow directly from Proposition 3.6, since for any permutation $\pi_{n}\in\mathsf{S}_{\infty}$ the matrix $\pi_{n}G_{n}$ is, respectively, symmetric and has zero diagonal if $G_{n}$ is symmetric and has no self-loops.
-For the third claim, since $G_{n}$ is symmetric, we have $G_{n}^{\top}\mathbbm{1}=G_{n}\mathbbm{1}=\deg(G_{n})$ is the vector of degrees of the vertices in $G_{n}$.
+For the third claim, since each $G_{n}$ is simple, we have $G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1}\in\{0,1/n\}$ which converges entrywise to zero.
+Thus $E=0$ by Proposition 3.6.
+Since $G_{n}$ is symmetric, we have $G_{n}^{\top}\mathbbm{1}=G_{n}\mathbbm{1}=\deg(G_{n})$ is the vector of degrees of the vertices in $G_{n}$.
 If $(G_{n})$ is dense, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{c}{n}\xrightarrow{n\to\infty}0$ for some constant $c>0$.
@@
 In particular, this sufficient condition implies that the random quotients $(\rho(F_{k,n})G_{n})$ and the grapheurs $(\mathsf{M}_{G_{n}})$ converge not just weakly but also in total variation.
-Such total variation convergence was shown for graphs sampled both from graphons [lovasz2012large, Cor.
-10.25] and from graphexes [borgs2019sampling, §5], and was the primary mode of convergence studied for edge-exchangeable random graph models [Crane_Dempsey_2019].
+Such total variation convergence was shown for graphs sampled both from graphons [34, Cor.
+10.25] and from graphexes [7, §5], and was the primary mode of convergence studied for edge-exchangeable random graph models [16].
 However, this condition is too restrictive for our purposes in this paper; for example, it does not apply to the sequence of stars in Example 2.2.
@@
 In this section, we characterize convergence of grapheurs in terms of graphons and vice-versa, exhibiting a facet of the duality between the two graph limit theories.
-Since graphon theory is most simply expressed in the undirected case (see [diaconis2007graph, §8] for the directed case), we restrict ourselves to this setting in the present section.
+Since graphon theory is most simply expressed in the undirected case (see [17, §8] for the directed case), we restrict ourselves to this setting in the present section.
 We begin by recalling a few basic notions from graphon theory.
@@
 We define two kernels $W$ and $U$ to be equivalent, denoted $W\sim U$, if there exist (Lebesgue-)measure-preserving maps $\varphi,\psi\colon[0,1]\to[0,1]$ satisfying $W\circ(\varphi,\varphi)=U\circ(\psi,\psi)$.
-This is the weak isomorphism between kernels commonly defined in the dense graph limits literature [lovasz2012large, Cor.
+This is the weak isomorphism between kernels commonly defined in the dense graph limits literature [34, Cor.
 10.35].
 We denote by $\mathcal{W}=\widetilde{\mathcal{W}}$$\sim$ the quotient space under this equivalence relation.
-Elements of $\mathcal{W}$ are called graphons [lovasz2012large, §7.1].
+Elements of $\mathcal{W}$ are called graphons [34, §7.1].
 A graphon $W\in\mathcal{W}$ defines random simple graphs of each size by setting $\mathbb{G}_{k}[W]=(\mathsf{B}_{T_{i},T_{j}}^{(W)})_{i,j\in[k]}$ where $\mathsf{B}_{T_{i},T_{j}}^{(W)}\sim\mathrm{Ber}(W(T_{i},T_{j}))$ are drawn independently for $i<j$ while $\mathsf{B}_{T_{j},T_{i}}^{(W)}=\mathsf{B}_{T_{i},T_{j}}^{(W)}$ for $i>j$ and $\mathsf{B}_{T_{i},T_{i}}^{(W)}=0$.
@@
 [^4] Graphon convergence is usually defined by total variation convergence of these random graphs, which is equivalent to their weak convergence since they are supported on the finite set $\mathbb{S}^{k}_{\mathrm{sim}}$.
-This is equivalent to convergence of homomorphism densities and to convergence in cut metric [lovasz2012large, §10].
+This is equivalent to convergence of homomorphism densities and to convergence in cut metric [34, §10].
 Finally, a finite graph $G\in\mathbb{S}^{n}_{\mathrm{sim}}$ defines a step graphon $W_{G}(x,y)=G_{i,j}$ if $x\in I_{i}^{(n)}\times I_{j}^{(n)}$ where $I_{i}^{(n)}=[(i-1)/n,i/n)$.
@@
 Therefore, we get $\langle G,\mathbb{G}_{k}[W_{n}]\rangle\to\langle G,\mathbb{G}_{k}[W]\rangle$ for all $G\in\mathbb{S}^{k}$ by linearity and the fact that $\mathbb{G}_{k}[W_{n}]$ all have zero diagonal.
-This implies $\mathbb{G}_{k}[W_{n}]\to\mathbb{G}_{k}[W]$ weakly by Cramér–Wold [kallenberg1997foundations, Cor.
+This implies $\mathbb{G}_{k}[W_{n}]\to\mathbb{G}_{k}[W]$ weakly by Cramér–Wold [29, Cor.
 6.5].
@@
 Theorem 3.10 exhibits a concrete, formal duality between quotient convergence and graphon convergence.
-This duality can be extended to general grapheurs and more general kernels on $0,1]^{2}$, but the extension of graphon convergence to this setting is more complicated [diaconis2007graph, §8].
+This duality can be extended to general grapheurs and more general kernels on $0,1]^{2}$, but the extension of graphon convergence to this setting is more complicated [17, §8].
 We therefore leave this and other extensions of this duality for future work.
@@
 For the first claim, since $\mathsf{M}$ is exchangeable we conclude that $\mu=\mathbb{E}\mathsf{M}$ is a (deterministic) exchangeable measure on $0,1]^{2}$, meaning that $\mu\circ(\sigma,\sigma)=\mu$ for all (Lebesgue)-measure preserving bijections $\sigma\colon[0,1]\to[0,1]$.
-The space of such measures is two-dimensional and is spanned by $\lambda^{2},\lambda_{D}$ by [kallenberg2005probabilistic, Thm.
+The space of such measures is two-dimensional and is spanned by $\lambda^{2},\lambda_{D}$ by [26, Thm.
 9.12], so $\mu=\theta\lambda+\vartheta\lambda_{D}$ for some $\theta,\vartheta\in\mathbb{R}$.
@@
 Proof.
-We apply the disintegration theorem [disintegration, Thm.
+We apply the disintegration theorem [39, Thm.
 A] to the continuous map $\pi_{k}\colon\mathcal{P}([0,1]^{2})\to\Delta^{k\times k}$ between compact metric spaces sending $\mu\mapsto(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}$.
@@
 Next, we turn to proving Proposition 3.3, which states that $\mathcal{M}$ is weakly compact and that the metric $W_{\square}$ metrizes this weak topology.
-To prove that $\mathcal{M}$ is weakly compact, we rely on the following lemma and the characterization of $\mathcal{M}$ from [kallenberg2005probabilistic, §9] as the space of ergodic random exhangeable measures.
+To prove that $\mathcal{M}$ is weakly compact, we rely on the following lemma and the characterization of $\mathcal{M}$ from [26, §9] as the space of ergodic random exhangeable measures.
 Lemma 3.14.
@@
 Thus, it suffices to prove that $\mathcal{M}\subseteq\mathcal{P}(\mathcal{P}([0,1]^{2}))$ is closed in this weak topology.
-By [kallenberg2005probabilistic, Prop.
+By [26, Prop.
 9.1], the space of random exchangeable measures is precisely the collection of random measures on $0,1]^{2}$ invariant under the countable group $\widetilde{\mathsf{S}}_{\infty}$ consisting of permutations of the intervals $(I_{i}^{(n)})_{i\in[n]}$ for all $n\in\mathbb{N}$.
-By [kallenberg2005probabilistic, Lemma A1.2], we can characterize its extreme points by
+By [26, Lemma A1.2], we can characterize its extreme points by
 $\mathcal{M}=\left\{\mathsf{M}\textrm{ random exchangeable}:\mathbb{P}[\mathsf{M}\in B]\in\{0,1\}\textrm{ whenever }B\subseteq\mathcal{P}([0,1]^{2})\textrm{ is }\widetilde{\mathsf{S}}_{\infty}\textrm{-invariant}\right\}.$
@@
 Equivalence of statements 1 and 2:
-Endow $\mathbb{R}^{2}$ with the $\ell_{\infty}$ norm, endow $\mathcal{P}([0,1]^{2})$ with the Wasserstein-1 metric with respect to this norm, and endow $\mathcal{M}$ with the Wasserstein-1 metric with respect to the Wasserstein-1 metric on $\mathcal{P}([0,1]^{2})$, which metrizes weak convergence of random measures [villani2008optimal, Thm.
+Endow $\mathbb{R}^{2}$ with the $\ell_{\infty}$ norm, endow $\mathcal{P}([0,1]^{2})$ with the Wasserstein-1 metric with respect to this norm, and endow $\mathcal{M}$ with the Wasserstein-1 metric with respect to the Wasserstein-1 metric on $\mathcal{P}([0,1]^{2})$, which metrizes weak convergence of random measures [46, Thm.
 6.9].
@@
 $\mathrm{dist}_{W_{1}}(\mathsf{M}_{1},\mathsf{M}_{2})=\min_{\begin{subarray}{c}\textrm{random }(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})\\ \mathsf{M}_{i}^{\prime}\overset{d}{=}\mathsf{M}_{i}\textrm{ for }i=1,2\end{subarray}}\mathbb{E}_{(\mathsf{M}_{1}^{\prime},\mathsf{M}_{2}^{\prime})}\sup_{\begin{subarray}{c}f\colon[0,1]^{2}\to\mathbb{R}\\ f(0)=0\\ \textrm{ 1-Lipschitz in }\ell_{\infty}\end{subarray}}|\mathbb{E}_{\mathsf{M}_{1}^{\prime}}f-\mathbb{E}_{\mathsf{M}_{2}^{\prime}}f|,$
-where the min over couplings is attained [villani2008optimal, Thm.
+where the min over couplings is attained [46, Thm.
 4.1].
@@
 $\begin{bmatrix}x_{1}^{(n)}&\cdots&x_{k}^{(n)}\\ y_{1}^{(n)}&\cdots&y_{k}^{(n)}\end{bmatrix}\in[0,1]^{2\times k},$ (15)
-obtained by first sampling $\mu_{n}\sim\mathsf{M}_{n}$ and then sampling $(x_{i},y_{i})\overset{iid}{\sim}\mu_{n}$ converge weakly by [kallenberg1973characterization, Thm.
+obtained by first sampling $\mu_{n}\sim\mathsf{M}_{n}$ and then sampling $(x_{i},y_{i})\overset{iid}{\sim}\mu_{n}$ converge weakly by [27, Thm.
 1.1], since the distributions of (15) are precisely the finite-dimensional distributions of $\mathsf{M}_{n}$.
@@
 Remark 4.2 (Edge-exchangeable random graphs).
-The random graphs we obtain by sampling edges from a grapheur appeared previously in the literature as edge-exchangeable random multigraph models [janson2018edge].
-As explained in [janson2018edge, Rmk.
+The random graphs we obtain by sampling edges from a grapheur appeared previously in the literature as edge-exchangeable random multigraph models [24].
+As explained in [24, Rmk.
 4.4], every edge-exchangeable random multigraph model is obtained from a random exchangeable measure on $0,1]^{2}$ as above after labelling the vertices by iid labels from $\mathrm{Unif}([0,1])$.
 Nevertheless, to our knowledge graph limits based on these edge-exchangeable random graphs as in Proposition 4.1 have not been previously studied in the literature, see also Section 1.2.
-As noted in [janson2018edge], if we repeat the above edge sampling process infinitely-many times, each edge $(T_{i},T_{j})$ corresponding to a positive edge weight $E_{i,j}>0$ will be repeated infinitely-many times.
+As noted in [24], if we repeat the above edge sampling process infinitely-many times, each edge $(T_{i},T_{j})$ corresponding to a positive edge weight $E_{i,j}>0$ will be repeated infinitely-many times.
 In contrast, all the other edges sampled above will only appear once.
 In particular, the vertices $T^{\prime},T_{1}^{\prime},T_{2}^{\prime}$ sampled in this process will only appear in one edge.
-Such vertices are called “blips” in [borgs2019sampling, janson2018edge], and the isolated edges and self-loops are called “dust”.
+Such vertices are called “blips” in [7, 24], and the isolated edges and self-loops are called “dust”.
 4.2 Edge-Based Szemerédi Regularity
@@
 Second, note that $f$ is $1/n$-Lipschitz in each $(x_{i},y_{i})$.
-Indeed, if $\mathsf{G}^{(n,i)}[\mathsf{M}]$ differs from $\mathsf{G}^{(n)}[\mathsf{M}]$ only in the $i$th location $(x_{i},y_{i})$, and we glue the two couplings $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})$ attaining the corresponding $\overline{W_{\square}}$ distances as in [villani2008optimal, §1], then
+Indeed, if $\mathsf{G}^{(n,i)}[\mathsf{M}]$ differs from $\mathsf{G}^{(n)}[\mathsf{M}]$ only in the $i$th location $(x_{i},y_{i})$, and we glue the two couplings $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})$ attaining the corresponding $\overline{W_{\square}}$ distances as in [46, §1], then
 $|f(\mathsf{G}^{(n,i)}[\mathsf{M}])-f(\mathsf{G}^{(n)}[\mathsf{M}])|\leq\mathbb{E}_{(\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})}\sup_{R\in\mathcal{AAR}}|\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}(R)-\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}(R)|\leq\frac{1}{n},$
 since the only difference between $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}$ and $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}$ is a weight of $1/n$ placed at one of their atoms.
-Since $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are sampled iid from a fixed measure $\mu$ (after it was sampled once from $\mathsf{M}$), the bounded-difference concentration inequality [wainwright2019high, Cor.
+Since $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are sampled iid from a fixed measure $\mu$ (after it was sampled once from $\mathsf{M}$), the bounded-difference concentration inequality [47, Cor.
 2.21] implies that $f(\mathsf{G}^{(n)}[\mathsf{M}])\leq\mathbb{E}f(\mathsf{G}^{(n)}[\mathsf{M}])+\epsilon/\sqrt{n}$ with probability at least $1-\exp(-2\epsilon^{2})$, proving the theorem.
@@
 ∎
+We remark that $\overline{W_{\square}}$ also metrizes weak converge in $\mathcal{M}$, as can be shown by combining Theorem 4.3 with Proposition 4.1.
+We opted to use the $W_{\square}$ metric in the rest of the paper since it is easier to relate to random quotients as in Proposition 3.2.
 We conclude this section by remarking that the above discussion applies to undirected random graphs sampled from symmetric grapheurs.
@@
 Define $\mathsf{M}_{<\infty}=\{\mathsf{M}_{G}:G\in\Delta^{n\times n}\textrm{ for some }n\}\to\mathbb{R}$ to be the grapheurs in $\mathsf{M}$ arising from finite graphs and define $\bar{f}\colon\mathsf{M}_{<\infty}\to\mathbb{R}$ by $\bar{f}(\mathsf{M}_{G})=f(G)$.
-This is well-defined because $f$ is permutation-invariant and isolate-indifferent, as $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ if and only if $G$ and $G^{\prime}$ differ by isolated vertices and node relabeling [kallenberg1990exchangeable, Prop.
+This is well-defined because $f$ is permutation-invariant and isolate-indifferent, as $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ if and only if $G$ and $G^{\prime}$ differ by isolated vertices and node relabeling [28, Prop.
 3].
@@
 Proof.
-For the first claim, it was shown in [levin2025deFin, Thm.
+For the first claim, it was shown in [31, Thm.
 5.2] that there exists a polynomial $q_{k}$ over $\mathbb{R}^{k\times k}$ satisfying $f(G_{n})=\mathbb{E}q_{k}(\rho(F_{k,n})G_{n})$ for all $G_{n}\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$.
@@
 for all large $n,m$, showing that $(f(G_{n}))$ is convergent. ∎
-We remark that the graphon analog of the characterization of quotient-testable parameters in Theorem 4.10 was used as the definition of testable parameters in [convergent_seqs1, Def. 2.11].
+We remark that the graphon analog of the characterization of quotient-testable parameters in Theorem 4.10 was used as the definition of testable parameters in [4, Def. 2.11].
 The upshot of this result is that a continuous graph parameter being quotient-testable is equivalent to the parameter being testable using edge sampling.
@@
 where $S_{2}^{(\mathrm{out})},S_{2}^{(\mathrm{in})}\in\mathbb{N}^{3\times 3}$ are outward- and inward-pointing stars with two leaves, respectively.
+This is an extension to weighted and directed graphs of the so-called first Zagreb index, originally introduced for chemical networks in [19].
 Define the polynomial $q_{2}$ on $2\times 2$ matrices by
@@
 For comparison, the value of this parameter on an Erdős–Rényi random graph with edge probability $1/2$ on 1,500 vertices is $0.001$.
+We end this section by giving two further examples of quotient-testable parameters, each quantifying different structural properties.
+Example 4.12 (Global Clustering Coefficient).
+One extension of the global clustering coefficient to weighted undirected graphs, a global analog of the coefficient proposed in [48, §7] for gene co-expression networks, is
+$f(G)=\frac{\mathrm{inj}(K_{3};G)}{\mathrm{inf}(P_{2};G)\max_{i,j}G_{i,j}},$ (16)
+where $\mathrm{inj}(H;G)$ denotes the injective homomorphism number and is defined as in (2) but with the sum restricted to injective maps, $K_{3}$ is an arbitrarily-directed triangle graph, and $P_{2}$ is the directed path on three vertices.
+We proceed to show that (16) is quotient-testable.
+Proposition 4.9 shows that $\mathrm{inj}(K_{3};G)$ and $\mathrm{inj}(P_{2};G)$ extend to continuous functions on all of $\mathcal{M}$.
+Similarly, we note that $G\mapsto\max_{i,j}G_{i,j}$ extends to such a continuous function, since if $(G_{n})$ quotient-converges to a grapheur described by $(E,\sigma,\varsigma,\theta,\vartheta)$, then $\max_{i,j}(G_{n})_{i,j}\to\max_{i,j}E_{i,j}$ by Proposition 3.6.
+Finally, note that $f(G)\leq 1$ for all graphs $G$ of any size, proving that $f$ extends to a continuous function on all of $\mathcal{M}$ and hence is quotient-testable by Proposition 4.8.
+Example 4.13 (Katz Centrality).
+Another measure of importance, or influence, of a node, originally introduced in [30] for social networks, is its Katz centrality.
+Specifically, for a parameter $\alpha\in(0,1)$, we consider the sum of the Katz centralities of all the nodes
+$f(G)=\mathbbm{1}^{\top}\Big[(I-\alpha G)^{-1}-I\Big]\mathbbm{1}=\sum_{k\geq 1}\alpha^{k}(\mathbbm{1}^{\top}G^{k}\mathbbm{1}),$
+which is another parameter quantifying the presence of hubs in a network.
+Once again, this parameter is quotient-testable.
+Indeed, if $(G_{n})$ quotient-converges, then $\mathbbm{1}^{\top}G_{n}^{k}\mathbbm{1}$ converges for each $k$ by Proposition 4.9, hence $f(G_{n})$ converges as well for any fixed $\alpha\in(0,1)$.
 4.4 Missing Proofs from Section 4
@@
 Therefore, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}\overset{d}{=}\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}=\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$, the latter being an empirical measure drawn from $\mu\sim\mathsf{M}$.
-Finally, we have $\mathsf{M}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}_{\mathsf{G}^{(n)}}$ as can be seen from (16), as $\mathsf{G}^{(n)}$ only depends on the multiplicities of $\{(t_{i},t_{j})\}$ rather than their locations, and the multiplicities and locations of a random exchangeable measure are independent of each other [kallenberg1990exchangeable, Thm.
+Finally, we have $\mathsf{M}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}_{\mathsf{G}^{(n)}}$ as can be seen from (17), as $\mathsf{G}^{(n)}$ only depends on the multiplicities of $\{(t_{i},t_{j})\}$ rather than their locations, and the multiplicities and locations of a random exchangeable measure are independent of each other [28, Thm.
 2].
@@
 Proof (Lemma 4.5)..
-Fix any $\mu\in\mathcal{P}([0,1]^{2})$. By the proof of [wainwright2019high, Thm. 4.10], we have
+Fix any $\mu\in\mathcal{P}([0,1]^{2})$. By the proof of [47, Thm. 4.10], we have
 $\underset{\begin{subarray}{c}\mu^{(n)}\\ \textrm{emp. meas.}\sim\mu\end{subarray}}{\mathbb{E}}\sup_{R\in\mathcal{AAR}}|\mu_{n}(R)-\mu(R)|\leq 2R_{n}(\mathcal{F}_{\mathcal{AAR}}),$
 where $R_{n}(\mathcal{F})=\underset{\begin{subarray}{c}X_{i}\overset{iid}{\sim}\mu\\ \varepsilon_{i}\overset{iid}{\sim}\mathrm{Unif}(\{\pm 1\})\end{subarray}}{\mathbb{E}}\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|$ is the Rademacher complexity of a class of functions $\mathcal{F}$ and we set $\mathcal{F}_{\mathcal{AAR}}=\{\mathbbm{1}_{R}:R\in\mathcal{AAR}\}$.
-We further have by [wainwright2019high, Ex.
+We further have by [47, Ex.
 5.24] that
@@
 If $\mathcal{F}$ is class of $\{0,1\}$-valued functions, then $\mathrm{diam}(\mathcal{F})\leq 1$.
-Finally, we have by [HAUSSLER1995217, Cor.
+Finally, we have by [20, Cor.
 1] that
@@
 We begin by arguing that $(\mathsf{G}_{k}[\mathsf{M}])$ is equipartition-consistent.
-Note that $d_{k,nk}\circ F_{nk,m}\overset{d}{=}F_{k,m}$ for any $k,n,m\in\mathbb{N}$ by [levin2025deFin, Lemma 4.13], since each $i\in[m]$ is mapped independently by $F_{nk,m}$ to one of the fibers of $d_{k,nk}$ with the same probability.
+Note that $d_{k,nk}\circ F_{nk,m}\overset{d}{=}F_{k,m}$ for any $k,n,m\in\mathbb{N}$ by [31, Lemma 4.13], since each $i\in[m]$ is mapped independently by $F_{nk,m}$ to one of the fibers of $d_{k,nk}$ with the same probability.
 Therefore, for any $G\in\Delta^{n\times n}$ the sequence of distributions $(\mathrm{Law}(\rho(F_{k,n})G_{n}))_{k}$ is indeed an equipartition-consistent model.
@@
 Proposition 5.1.
-Fix a grapheur $\mathsf{M}\in\mathcal{M}$ and consider the associated equipartition-consistent random graph model $(\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$.
+For any grapheur $\mathsf{M}\in\mathcal{M}$, denote by $(\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$ its associated equipartition-consistent random graph model.
   1. 1.
-    Draw $\mu\sim\mathsf{M}$ and let $\mathsf{G}_{k}=(\mu(I_{i}^{(n)}\times I_{j}^{(n)}))_{i,j\in[k]}$ for each $k\in\mathbb{N}$.
-Then $G_{k}\sim\mu_{k}$ for all $k\in\mathbb{N}$, and $(G_{k})$ is quotient-convergent to $\mathsf{M}$ almost surely.
+    Fix a grapheur $\mathsf{M}\in\mathcal{M}$, draw $\mu\sim\mathsf{M}$, and let $\mathsf{Q}_{k}=(\mu(I_{i}^{(n)}\times I_{j}^{(n)}))_{i,j\in[k]}$ for each $k\in\mathbb{N}$.
+Then $\mathrm{Law}(\mathsf{Q}_{k})=\mu_{k}$ for all $k\in\mathbb{N}$, and $(\mathsf{Q}_{k})$ is quotient-convergent to $\mathsf{M}$ almost surely.
   2. 2.
     Suppose $\mathsf{M}=\delta_{(T_{1},T_{2})}$ is the grapheur associated to a single edge.
-Draw $G_{k}\sim\mu_{k}$ independently of each other.
-Then $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely, so $(G_{k})$ does not converge to $\mathsf{M}$.
+Draw $\mathsf{R}_{k}\sim\mu_{k}$ independently for different $k$.
+Then $\limsup_{k}W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ almost surely, so $(\mathsf{R}_{k})$ does not converge to $\mathsf{M}$.
 Proof.
@@
 Therefore, there exists $K\geq N\in\mathbb{N}$ such that $T_{1},\ldots,T_{N}$ lie in distinct intervals $(I_{i}^{(k)})_{i\in[k]}$ for all $k\geq K$.
-Therefore, for all $k\geq K$ we can find a permutation $\pi_{k}\in\mathsf{S}_{k}$ such that $\pi_{k}G_{k}=\rho(\iota_{k,N})(E_{i,j}+\sigma_{i}/k+\varsigma_{j}/k)_{i,j\in[N]}+N_{k}+\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+\vartheta\frac{1}{k}I_{k}$ where $\|N_{k}\|_{1}\leq\epsilon$.
-Using the coupling $(\mathsf{M}_{G_{k}}^{\prime},\mathsf{M}^{\prime})$ obtained by drawing random locations $(T_{i})$, setting $\mathsf{M}^{\prime}$ to (5), and setting $\mathsf{M}^{\prime}_{G_{k}}=\sum_{i,j=1}^{N}(\pi_{k}G_{k})_{i,j}\delta_{(T_{i},T_{j})}$, we conclude that $\limsup_{k\to\infty}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\leq\epsilon$.
+Therefore, for all $k\geq K$ we can find a permutation $\pi_{k}\in\mathsf{S}_{k}$ such that $\pi_{k}\mathsf{Q}_{k}=\rho(\iota_{k,N})(E_{i,j}+\sigma_{i}/k+\varsigma_{j}/k)_{i,j\in[N]}+N_{k}+\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+\vartheta\frac{1}{k}I_{k}$ where $\|N_{k}\|_{1}\leq\epsilon$.
+Using the coupling $(\mathsf{M}_{\mathsf{Q}_{k}}^{\prime},\mathsf{M}^{\prime})$ obtained by drawing random locations $(T_{i})$, setting $\mathsf{M}^{\prime}$ to (5), and setting $\mathsf{M}^{\prime}_{\mathsf{Q}_{k}}=\sum_{i,j=1}^{N}(\pi_{k}\mathsf{Q}_{k})_{i,j}\delta_{(T_{i},T_{j})}$, we conclude that $\limsup_{k\to\infty}W_{\square}(\mathsf{M}_{\mathsf{Q}_{k}},\mathsf{M})\leq\epsilon$.
 Since $\epsilon>0$ was arbitrary, we obtain the claimed convergence.
-For the second claim, note that with probability $1/k$ we have $T_{1},T_{2}\in I_{i}^{(k)}$ for the same $i\in[k]$, in which case $G_{k}=\mathrm{diag}(0,\ldots,1,\ldots,0)$ and $W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ as can be seen by setting $S=[0,1/2]$ and $T=[1/2,1]$ in (7).
-Since these events are independent and the sum of their probabilities $\sum_{k}k^{-1}$ diverges, Borel–Cantelli implies that $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely.
+For the second claim, note that with probability $1/k$ we have $T_{1},T_{2}\in I_{i}^{(k)}$ for the same $i\in[k]$, in which case $\mathsf{R}_{k}=\mathrm{diag}(0,\ldots,1,\ldots,0)$ and $W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ as can be seen by setting $S=[0,1/2]$ and $T=[1/2,1]$ in (7).
+Since these events are independent and the sum of their probabilities $\sum_{k}k^{-1}$ diverges, Borel–Cantelli implies that $\limsup_{k}W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ almost surely.
 ∎
 In words, the quotients must be sampled in a manner that is related across different sizes for the samples to converge to the underlying grapheur; if we instead sample random quotients independently for each size, the resulting sequence almost surely does not converge to the underlying grapheur.
-In contrast, if we sample growing-sized random graphs from a graphon independently for different sizes, then the resulting sequence does converge back to the graphon almost surely as a consequence of Szemerédi’s regularity lemma [lovasz2012large, Lemma 11.8].
+In contrast, if we sample growing-sized random graphs from a graphon independently for different sizes, then the resulting sequence does converge back to the graphon almost surely as a consequence of Szemerédi’s regularity lemma [34, Lemma 11.8].
 Proposition 5.1(2) thus implies that the analog of Theorem 1.7 fails for random quotients.
@@
 Our goal in this section is to prove Theorem 1.9, showing that every equipartition-consistent random graph model arises from a unique random exchangeable measure.
-We deduce the existence of the desired random measure from results on inverse limits of random histograms [orbanz2011projective], and proving that this random measure is exchangeable proceeds via standard approximation arguments.
-To apply the results of [orbanz2011projective], we require the following lemma on the sequence of means of an equipartition-consistent model.
+We deduce the existence of the desired random measure from results on inverse limits of random histograms [38], and proving that this random measure is exchangeable proceeds via standard approximation arguments.
+To apply the results of [38], we require the following lemma on the sequence of means of an equipartition-consistent model.
 Lemma 5.2.
@@
 ∎
-Combining the above lemma with [orbanz2011projective], we are ready to prove Theorem 1.9.
+Combining the above lemma with [38], we are ready to prove Theorem 1.9.
 Proof (Theorem 1.9)..
 Uniqueness is clear since squares of the form $I_{i}^{(k)}\times I_{j}^{(k)}$ generate the Borel $\sigma$-algebra on $0,1]^{2}$.
-To prove existence we appeal to [orbanz2011projective, Thm.
+To prove existence we appeal to [38, Thm.
 1.1], which states that there exists a random probability measure $\mathsf{M}$ on $0,1]^{2}$ (not guaranteed to be exchangeable) satisfying $\mathsf{M}(I_{i}^{(k)}\times I_{j}^{(k)})\sim\mu_{k}$ for all $k$ if there is a determinisitic measure $\mu\in\mathcal{P}([0,1]^{2})$ satisfying $(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}=\mathbb{E}_{X\sim\mu_{k}}X$ for all $k\in\mathbb{N}$.
@@
 To do so, it suffices to prove that for any continuous function $f\colon[0,1)^{2}\to\mathbb{R}$ and any measure-preserving bijection $\sigma$, we have $\mathbb{E}_{\mathsf{M}}f\overset{d}{=}\mathbb{E}_{\mathsf{M}}f\circ(\sigma,\sigma)$.
-By [kallenberg2005probabilistic, Prop.
+By [26, Prop.
 9.1], it further suffices to prove this equality for $\sigma\in\mathsf{S}_{m}$, a permutation of the $m$ intervals $I_{1}^{(m)},\ldots,I_{m}^{(m)}$ acting by $\sigma((i-1)/m+x)=(\sigma(i)-1)/m+x$ for $x\in[0,1/m)$ for $i\in[m]$.
@@
 The limiting grapheur of a quotient-convergent sequence of graphs describes the asymptotic distribution of edge weight in the graphs.
-We then presented another equivalent view of quotient convergence via edge sampling, allowing us to prove a dual analog of Szemerédi’s regularity lemma and to test properties of arbitrarily-large graphs by sampling edges from them.
+We then presented another equivalent view of quotient convergence via edge sampling, allowing us to prove an edge-based analog of Szemerédi’s regularity lemma and to test properties of arbitrarily-large graphs by sampling edges from them.
 The number of edges needed is indpendent of the size of the graph, only depending on the desired accuracy and the Lipschitz constant of the graph parameter in question.
@@
     Is the rate of $O(1/\sqrt{n})$ for approximating a grapheur by a graph on $n$ edges optimal?
-  2. item(Rates for random quotients)
-    Is there a Szemerédi-type regularity lemma for random quotients, i.e., does $\mathbb{E}_{\mathsf{G}_{k}[\mathsf{M}]}W_{\square}(\mathsf{M}_{\mathsf{G}_{k}[\mathsf{M}]},\mathsf{M})$ converge to zero at a universal rate independent of $\mathsf{M}$?
-If not, how does this rate depend on $\mathsf{M}$?
   2. item(Extending graphon duality)